A Complete Bibliography of the Journal of Econometrics (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.01

Title word cross-reference

AR(1) [Kru08, Kru14, PSS03a, PSS03b]. ARCH(∞) [HZ07b, HZ07c]. F [BM00b]. GARCH(1, 1) [Cor00]. hv [Rac00]. k [GL02]. m [Deo00a, GH06a]. N [GV07, HR07]. Q [HLNS06]. t [BM00b, Des06, GGHW08]. τ [Fit04].

-block [Rac00]. -class [GL02]. -consistent [HR07]. -dependent [Deo00a]. -distributions [GGHW08]. -estimators [Fit04]. -tests [BM00b].

/unit [Kru14, Kru08].

106 [CK07, Dav02a]. 108 [BT03a]. 109 [BZ04]. 112 [GKLT05]. 114 [HI07]. 117 [PSS03a]. 119 [Cha08]. 134 [HLT12]. 138 [Zel07a]. 141 [HLT08]. 144 [Kru14]. 1980s [CL00]. 1990s [CvdK08].

2000 [Ano00s, Ano00r, Ano00p, Ano00n, Ano00q, Ano00l, Ano00m, Ano00k, Ano00j, Ano00o, Ano00i]. 2001 [Ano01o, Ano01m, Ano01p, Ano01r, Ano01q, Ano01n, Ano01l, Ano01k]. 2002 [Ano02s, Ano02q, Ano02r, Ano02p, Ano02m, Ano02o, Ano02n, Ano02l]. 2003 [Ano03s, Ano03t, Ano03w, Ano03v, Ano03y, Ano03u, Ano03r]. 2004 [Ano04s, Ano04u, Ano04t, Ano04v, Ano04r, Ano04q, Gra06]. 2005 [Ano05z, Ano05-28, Ano05x, Ano05v, Ano05y, Ano05-27, Ano05w]. 2006 [Ano06x, Ano06u, Ano06p, Ano06v, Ano06r, Ano06t, Ano06s, Ano06w, Ano06q]. 2007 [Ano07p, Ano07q, Ano07o, Ano07n, Ano07r]. 2008 [Ano08r, Ano08q, Ano08s, Ano08p, Ano08o]. 2009 [Ano09n, Ano09t, Ano09u, Ano09s, Ano09o, Ano09r, Ano09q, Ano09p]. 201-400 [Ano03x]. 2SLS [DLM06, fL07a].

30 [SCF08].

50th [Gra07].

75 [RU05].

’87 [Bat00].

95 [KA02].

Ability [Arc04, CM05, CSO01, HHM04]. absence [Pes04]. academically [Reb04]. accomplishments [SEGG06]. accuracy [BK06, CM01, CW07b, CS02, CS06c]. accurate [FRW05]. ACD [ACMP08]. achievement [HHM04, Mat08, vdK08]. Achievements [Hen01]. across [AT07, Lav01]. action [Lon04b]. actively [AAD05]. activity [BMS07, IV06, Phi01]. Adaptive [BBvDvO04, BGG08, XP08, CG07, SS06, VKG09]. addendum [PJH07]. additional [Hu06]. additive [PS08]. additivity [GL01]. Adjusted [Per00]. adjustment [DHL06, SS01]. adjustments [SS00]. Administrative [LSW08]. admissible [MS04]. admissions [MZ04]. advances [BLvD04]. adverts [Ano02t]. affect [Lal08]. affine [DMN05, EHL06, Sin01]. affinity [HD08]. affirmative [Lon04b]. against [BT03b, BT04b, HL09, Kle04]. age [Cos03, HSWW08]. agents [KM09]. aggregate [JMR09]. aggregated [CM07a, vDP08]. Aggregation [GG04, Zaf07, Bar07, DS05, Haf08, KLV04, MR04, Zaf04, vGLP00]. aggregation-theoretic [Bar07]. agreement [BEP08]. agricultural [PN05]. AHEAD [Gew03]. AIDS [WKV03]. Aigner [Ano05h, Ano08a]. AIM [FS08a]. air [AD05, MMS^{+}08b]. Aitken [GL00a]. al [Bas03b]. algorithm [JST06]. algorithms [BW02, BD04, Nie00]. allocation [MS09a, WW09]. allocative [KT05, KW06a, KW06b]. allowing [KP09b, MR03]. Alternative [CS07a, CGGT03, Dav06a, FKS05, RT04, SA04, DGR08, GO05, GT06, Hid03, HL09, KP09b, MMS^{+}08b, Phi00, SV07b]. alternatives [AD07, Nev02]. always [BN06b]. ambiguity [Man00]. Amemiya [Hsi09]. American [BDGT00]. Americans [LP03]. among [FTC03, GK00]. Analysis [BKP07, Bha03a, CNS06, Chi07, CJ08, ASM08b, And04, BNS06, Bas03b, BV04, BH07, Bie07, BN06b, BR09, CCL03, CP02, CMvS01, CH00, CH02, CKP01, DS07, DFG04, FKS00, FRW05, GK00, Hec03, HPU06, Hou03, IvD05, JJ00, JPR04, JLWY08, JMR09, Joh06, JDGZ03, Kim02, KM09, KP02, KW06b, LG02, Man00, MR01, MV02, MCM07, McC08a, MPR00, MT03, MT08b, Nak00, PSS00, PP04b, RT06, RR02, SMB07, SBES05, SI04, VI02, WW08, YS07, Zel02]. Analytical [Pes04, WZ03b]. Angrist [HKvD07a]. anniversary [Gra07]. Announcement [Ano09a, Hsi09]. ante [Mos01]. Antitrust [HMSA03]. apartments [EL03]. applicant [CH09]. Application [ACMP08, AHV05, ASBS01, AL01, BMM09, BIKY07, CL09, CS07a, CP04b, DS06, DJ06, EL03, FS07, GJ06, Gra06, HZ07a, Hau03, Hid02, Hon01, HKvD07b, Kar02, LW07, Lon04a, Mac04, MCM07, MT03, PW08, PSZ08, RT06, RJ02, SCF08, SV07a, SEGG06, WvDW03, WW08, Wu03, Yu07, ZBK09, vDFP02]. applications [BK08, BHPW04, CMP09, DEG01, GLS02, IL08a, KP01a, Kri08, Lon04b, LT06b, SR02, ZRT01]. applied [HKvD07a, Jin09]. approach [BR05a, BP07a, BM05, BC06, CFT06, CFS02, CH03b, CH08, CHMR08, DS07, DIW00, DP08b, DRA06, DLY08, Duf06, FVRR04, GRZ07, GKvD07, Gir00, GT09, HY05, HTV07, HK03, HS00, IL09, JO08, KV00, KT05, KW06a, KPST07, Lal08, Lee07, LM08a, MMS^{+}08b, MNRS03, Moe08, Nak00, NS07b, OC05, Pes07, Phi00, PY09b, ST01a, SL03, SK06a, Sko07, Stu03, ZBK09, vGK02]. approaches [BB02, KZ03, SP01]. approximate [KP09a, SK06b]. Approximately [CW07b]. Approximating [PS07b]. approximation [Bon03, rMCGJL08, HW08a, Yu07]. approximations [ASY06, CS07a, FS05, Lin02, Phi00, Ull02, XP02b]. April [Ano03x, Ano06t, Ano00s, Ano01o, Ano07p, Ano08r, Ano09n, Ano09t]. arbitrage [ABD07, Moe08]. arbitrarily [Mor09]. ARCH [BK07, DKBG04, HP08, KLV04, OP04]. ARCH-type [OP04]. ARFIMA [BS06]. arguments [ABW07]. arising [Gas03]. ARMA [FZ01, Kue01, Lin07, Nak00]. ARMA-GARCH [Lin07, Nak00]. ARMA-GARCH/IGARCH [Lin07]. Army [Cos03]. Arnold [Ano05a, Ano07t]. arrays [FPW04, JP09a]. arrivals [MMS^{+}08b]. Arrive [Coo08]. ascending [HS03]. aspects [Gra03, Zel07d]. Assessing [KP09a, KYH09]. assessment [GLR03, GRZ07, LHMC09]. asset [ASWY01, BL02a, Ber01, JLWY08, JO08, KS02, Sen04, Sin01, WW09]. assets [Dub07, PS00]. assistance [LM08a]. Assumptions [MR03, GO05, KSC09]. asymmetric [BP06, HS07, HS03, MCM07, PT07a, SL03, SS06, ZZW09]. asymmetries [MM08, PQT01]. Asymmetry [SL07b, EZ01]. Asymptotic [Bha05, CN06, DF07, DGR06, Gos08, Han07a, KH09b, McK04, OK00, Rei07, Seo07, AG09, CP04a, DM06, DV09, KP01a, LLC02, LP08d, Maa01, SB02, dJ02b]. asymptotically [Plo04]. Asymptotics [HHS07, McC07b, Duf06, Phi02]. attributable [JDGZ03]. attributes [KAR07]. attribution [GM06b]. attrition [Bha08, Das04]. auctions [BP06, Bre07, BP07c, EL03, HKL05, HS03, LPV00, Li05, McA08]. augmentation [IvD05]. augmented [IS08, Moe08]. August [Ano00r, Ano01m, Ano02r, Ano03s, Ano05z, Ano06x]. Australian [ACMP08]. Author [Ano01e, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano03c, Ano03a, Ano03d, Ano03e, Ano03f, Ano03b, Ano04b, Ano04c, Ano04a, Ano04d, Ano04e, Ano04f, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e]. Authors [Ano02t]. auto [CR08, XL04]. auto-regressions [XL04]. autocorrelated [Mül05, RT06]. autocorrelation [BSJK07, BLS06, Cho00]. Autocovariance [AT05, LP08d]. automatic [EL09]. autoregression [Gal09, PPC04, Rot02, Shi01]. autoregressions [Des06, DT00, GK04]. Autoregressive [GJS09, KYH09, MM08, BGG08, Cho05, DS04, Dis08, DSS07b, DT00, ES01, FG06, GW05, Gos02, Han03, HW06, fL07a, fL07c, MN00a, NS03, PB09, PT05, RJ02, RT04, XP08, ZRT01]. average [Ban07, For05, Frö07, IL09]. Averaging [Wri08, FLS01, Han08]. aversion [ASL00]. Award [Ano08a, Ano05a, Ano05h].

B [WW02]. B-robust [WW02]. back [Gra07]. Bagging [LY06]. Bahadur [KP09a]. balanced [DS04]. balances [HKL05]. band [CN06]. bandwidth [ZBK09]. bargaining [Fré09]. based [ABW07, AG09, BBvDvO04, BGG08, BL08, Ber01, Bow07, BT01, CDPX07, CF06b, CF06a, CP07a, CS07b, CDS09, DGR03, Dah09, DFG04, DG07, DK02b, DKBG04, DJ06, DT07, FFP07, GP02b, HIJS07, HH00, Hor02, Kar02, KP09a, Koo01, Kra06, LL09, LHZ03, Li09, MT03, PS09b, Ram02, RS00, Sak07, SS00, Sko07, ST01b, SS01, SW07b, Xu09, dJAS07, vAJM02]. baseline [CJ07]. basis [DG07]. Bayes [Zel07a, CH02, KP02, Rei07, Zel07b]. Bayesian [AD05, BLvD04, BHH08, CC04, CH00, CP04b, Cla07, CHMR08, DIW00, DGR08, FVRR04, FKS00, FLS01, GSN08, Gew01, GS04, HTV07, HK03, HM03, Hou03, IvD05, JJ00, JPR04, JMR09, Kim02, KP02, KZ03, Kle04, KY09, KV00, Koo01, KP04b, KT06a, KT05, LG02, LS08a, LW01, MN00a, MZ04, McC08a, MPR00, MS04, MT03, MT08b, Nak00, NS03, Nob00, OC05, OP04, PS08, RT06, Sar03, SP01, SI04, Vil06, Wri08, YS07, Zel02, Zel07d, ZBK09]. be [GM06a, HLZ07, MPS09, Mor09]. beat [EHL06]. beaten [HLZ07]. behavior [CDSV08, vdKW08]. behaviour [HI03, HI07, Rei07]. behind [KSC09]. beliefs [Leh09]. Benchmark [FLS01]. benefits [EL05, Lal08]. bereavement [WKV03]. Bertrand [FTC03]. better [BN06b]. between [AHM^{+}03b, ABCD07, CHS07, CSW00, HLW09, KS02, OZC08, Rob05, vdK08]. Beveridge [Kim08, Nel08, OZC08]. Beyond [ADG09]. Bias [PS07a, BU07b, CS07a, HS07, HD08, IL09, JL08, KMR07, LS09a, Mal08, RSU96, RU05, TC09]. bias-corrected [IL09]. biased [BR05a]. biases [CES08]. bilinear [FMZ08]. Binary [AC03b, BMM09, BD07, Che00a, Che00b, Cop01, FL03, KY09, LY06, LS08b, MM07, Moo04, MS06, Rot09]. binding [Li05, MT08a]. binomial [Mac04]. Birth [BvS08, LP03]. Birth-spacing [BvS08]. bivariate [CZ07, GTP06, Lie01, MS06, OP04]. black [CL00]. black-white [CL00]. block [Rac00]. blocks [And06, Hid03]. Board [Ano04p, Ano02i, Ano02j, Ano02k, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m]. bond [DL06, EHL06]. bonus [BR05b]. Bootstrap [Bie02, Cha04b, CS05, CS06a, CI08, DM02, HN05, HK06, Hor02, MPS09, Seo06, AB01, ALM06, DH05, Dav02c, Dav06a, DM06, DF07, DF08a, DF02, GW04, Gos08, HM03, Hid03, Hid05, HS00, Hor01b, KS08a, LT06a, Par06a, PPP06, Wha00, YH06, vGK02]. Bootstrapping [BT04a, CPS06, FNS04, GK04, HLNS06, IS06]. both [KP09b, RL04, YdJfL08]. bound [JST06]. Boundedly [Can07]. Bounding [Hu06]. bounds [DFH06, San07, ST01a]. Box [HLNS06]. branch [JST06]. branch-and-bound [JST06]. brand [CKP01]. break [EM07, KLN02, KP09a, KP09b, Kur02, MOC05]. Breaking [LP08a, vdK08]. Breaks [BM06, AC03a, Ano05u, BU05, BT03b, CM05, DH00, GTU05, GGV08, MN00a, PZ05, PT05, PT07b]. Britain [GO07]. buffer [HS02b]. Burg [BD04]. bus [EL05]. Business [PQT01, GRS06].

C. [Dei01]. Cactus [Bas03b]. Calculating [Lew03, CJ07]. Calculation [Wu03]. calendar [STW01]. calibration [DGR07]. Can [EHL06, HLZ07, DM04, Mor09]. canonical [DFG04]. cap [JKS03]. capacity [FTC03]. cards [CGS08]. Care [Vin01, EL05, SW01b, KYH09]. Carlo [BM00a, BBvDvO04, CNS02, CS06b, DGR06, Duf06, Dur06, GRZ07, KW06b, LG02, MN00b, VI02, Vil01]. case [Bas03b, CP02, CR08, Cop03, DMN05, JM07, San07, STW01]. cases [Can07, Gas03, JDGZ03, Lew03]. cash [HKL05]. categorical [LMR09, RL04]. Cauchy [SS00]. causal [AHM^{+}03b, Gra03, Hoo03, MR03]. Causality [BBU06, Vil01, ACM03, BC06, DJ06, DPR06, Eic07, Flo03, GRZ07, Hau03, Hec03, Hid05, HLW09, McC07b, MC06, MS06]. cause [DM04]. caused [Bha08]. causes [BM06, WvDW03]. CCA [DS04]. celebrate [Gra07]. Censored [BP07b, BCY00, CMvS00, CK00, DT09, HH04, KP01b, KT09, jLV06]. censoring [CGM08, HKP02, KT07]. Central [JP09a, Dav02a, Dav02b, PSS07b]. certain [Zd03]. ceteris [BS00]. Chain [Oul08, CNS02, LG02, Nak00]. chains [McC07a]. challenges [Bol01, FvD07, Hen01, Hsi01, Lüt01, Phi01]. Chamberlin [Bas08]. Change [Car02, BR05a, BMS07, BIKY07, BT04b, Can07, CT08, CCZ05, DP08a, Fit04, GG03, GKvD07, Han00, HLT06, HLT12, Kim00, KA02, Laz05, LS08b, PR03, PQ06, Qu08]. change-point [LS08b]. changed [Kim08, Lon04a]. Changes [CL00, HW08b, BC08b, CL09, CCZ08, Han03, Hil05, KP08]. changing [ABG05, JX05a, JX09, Zaf07]. chaos [SL04]. characteristic [Sin01, SW07b]. Characteristics [FGNW05, KS08b]. Charles [CDN08]. check [WW08]. chi [BSW04]. chi-squared [BSW04]. children [FGS03, JR03, LHMC09]. Chinese [Yu07]. choice [AC03b, BD07, BHH08, Car07, Che00a, Che00b, CW07a, HN07, HP04, HN00, Jin09, KY09, fL00, Lon04a, Man00, Man07, Moo04, MME08, Oku09, PJH07, Ram02, Sto09, Stu03, Tab00]. choice-based [Ram02]. choices [CKP01]. Choosing [DIN09, AB01]. CIR [JKS03]. citation [SV07a]. claim [JJ00]. class [FMZ08, GL02, Plo04]. classical [CH03b, Das03b, KZ03, KP04b]. classical-type [Das03b]. Clive [SEGG06]. Closed [Yu07]. Closed-form [Yu07]. clustered [CH00]. clustering [GM06b]. co [CHP09]. co-movements [CHP09]. coal [SBES05]. coefficient [BH07, CDXW06, Cai07, CLP09, CZ07, Dis08, JMR09, KLV04, KT06a, MW08, Xia09a]. coefficients [CK03, DFH06, ES01, Li08]. coincident [IV06]. cointegrated [And02, BW02, CM07a, CSO01, Han03, KP08]. cointegrating [ABG05, CPS06, Joh02, Joh06, KH09b, LS00, LST03, NS07b, RY02, Seo07, Shi01, Xia09b, dJ01, dJ02a]. Cointegration [RI05, ABG05, AV09, BKN09, BL02b, BH02, Bre02, BT03a, CH03a, CHO08, CN06, CS06b, Dav02c, Dav06a, HS02a, HML02, HPU06, HR07, HS02b, Jan05, JX05b, KP02, MR01, Pes04, Rob08b, Seo06, Vil06, XP02a, Xia09a]. collective [LP08c]. collectively [CDSV08]. college [Arc04, BS04, BGKT04, Cap04, CL09, Gro04, HD08, JLS05, Lon04a, Reb04, SS04, Lon04b, Rot04]. college-educated [BGKT04, Gro04]. collinear [DT07]. colloquy [KY09]. combination [AT06, GT09]. combinations [ET04, San07]. Combining [KPS07a, MJ05, vdBvdK01, DT09]. Comment [Lüt01, Nob00, WWM01, Kri01]. Comments [Dei01, Zel01, Gra05, Gra06, Hor01a]. Common [AIV06, EM06b, GTP06, HPU06, Par06b, Bai04, CMP09, CKL06, CS06b, GRS06, VI02, VF05]. community [JLS05]. comovement [CFT06]. comovements [CSW00]. company [CR08]. comparable [Das03b]. Comparing [FVRR04]. Comparison [GP02a, LST03, AL04, BR08, CS07b, FS07, Ink00, MSW06, MN00b, OP04]. comparisons [HMSA03]. compatible [CLP03]. compensation [Jur02, Lew03]. competing [JM07, Li09]. competition [FTC03, MC08]. complete [LP08d, Plo04]. completely [Can07]. complex [BEP07, Gre06, PSS07b]. compliance [BR05b]. component [BSJ01, BBP06, BSEY01, BR09, Cho02, fL00, PY09a]. components [CKL06, Dav02d, DGR08, DS06, KKP07]. composite [SK06b]. computation [rMCGJL08, FS08b, McC08a]. computing [HH00, ST01a]. concentration [PS07b]. concentrations [OK00]. condition [BS00, CHS07, EM06a, NS07a]. Conditional [KYH09, Ots08, AC07, AT06, AMS07, BN01, BV04, BK07, BZ02, BZ04, CW08, CGM08, CMvS00, CCZ08, CP09, CHJ06, CS06a, CS06c, DS07, DE07, DS08a, Deo00b, DIN03, DIN09, Esc08, FW08a, FG05, FG06, GK04, GTP06, Han00, JP09b, KT09, Kom05, LT06a, Lon04a, Lug03, MFY08, MMS^{+}08b, Muk07, Pal09, PB09, PQT01, RJ02, Sak07, Seo07, SSVY09, Smi07, SW07b, Vil01, ZRT01]. Conditionally [LPV00, SF01, SCF08]. Conditioning [Hec03, AC07, MS04]. conditions [CCFG07, Dav02a, Dav02b, MS09b]. Conference [Ano00a, Gra06]. Confidence [ES01, EM07, IS02, Ros08, CCL03, CS06c, GH09, vGK02]. conjugate [Gre06, HKvD07a]. Connections [KS02]. consequences [BS00, GL00a]. considerations [Rob03]. consistency [FHLR04]. Consistent [AL01, BR09, DG07, HL06, Hid02, LHZ03, LS09b, Rac00, Wha00, AG09, BN01, BGG08, CS02, DEG01, DIN03, GV07, HLR07, HR07, Li02, LT06a, Mac04, SW07b]. consistently [KL08]. constancy [ET07]. constant [Per00, Tse00]. constrained [YH06]. constraint [Nob00]. constraints [BMM09, Cho05, FTC03, MT08a]. construct [IV06]. Consumer [BS08, FW08b, Slo08]. Consumption [JS08, Ber01, BC08b, CDSV08, HZ07a]. consumption-based [Ber01]. contaminated [DF09, Mul06, RS00]. Contemporaneous [DSS07b, EZ01, Zaf04, DK02a]. content [ABR07, BPT01, CP04b]. contents [Ano04w, Ano07m]. contexts [HY05]. contextual [fL07b]. contingent [JJ00]. continuous [ASY06, ABD07, Bow07, Can07, CV03, Cor00, DP08b, HLR07, LMR09, RL04, Tod09]. continuous-time [ASY06, ABD07, CV03, Tod09]. continuum [CCFG07]. contracts [AT07]. Contributions [PP04b, CDN08, Dei01]. control [AG08, Jun09, Lee07]. Convergence [dJ02b, CK03, MRS07, Pes07]. convolution [Bon03]. coordinate [FS07]. coordinate-free [FS07]. Copula [LL09, CF06b, CF06a]. Copula-based [LL09, CF06b, CF06a]. copulas [PS09b]. copyrights [SMB07]. correct [Mor09]. corrected [IL09]. Correcting [BR05b, KMR07]. correction [BLS06, CS07a, HS07, HS02a, Hen06, HN05, Joh02, PSS00, Sai05, Seo06, Seo07, SI04, TC09, Win05]. correctly [ASWY01]. correlated [BH07, Cai07, HPT08, JPR04, KKP07, MW08, Tur00]. Correlation [Rob08a, BSK03, BSJK07, CZ07, CM07b, DK02a, DFH06, EL09, Han07b, Yam08]. correlations [AM09, Pal09, Pel06, Tse00]. Corrigendum [BZ04, BT03a, Cha08, CK07, Dav02a, GKLT05, HI07, HLT12, KA02, Kru14, PSS03a, RU05, PJH07]. cost [AP02, BC08b, FS08a, KT05, KW06b, Mos01, Oul08]. cost-of-living [Oul08]. costs [DMN05, JDGZ03]. count [BGW02]. countries [Bar07]. country [BEP07]. counts [HPT08]. course [Reb04]. covariance [BM00b, ET07, FFL08, For05, Hid05, HK06, Kle07, Kru08, Kru14]. covariate [LS08b, Mul06]. covariates [EJ03, Frö07, Gal09, HP08]. Cover [Ano02j, Ano02k, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04p, Ano04l, Ano04m, Ano04n, Ano04o, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano06o]. coverage [DF08b]. covering [HPT02]. crash [Bat00]. credible [HKvD07b]. credit [AT07, AG08, DF09, KLM08]. crediting [AD05]. crime [SW08b]. critical [HS00, Hor02, WZ03b]. critique [ST05b]. Cross [CG02, vGLP00, Bai04, BGKT04, Cha02, Cha04b, CH00, Gir00, GH06b, HM08, HH04, HS02b, KP04a, Mor04, PS07a, Rac00, Rob08a, SYR09, SK06a, VV05]. cross-equation [HS02b]. cross-section [Bai04, CH00, GH06b]. Cross-sectional [vGLP00, Cha02, Cha04b, Mor04, Rob08a, SK06a]. cross-sections [Gir00, VV05]. cross-state [BGKT04]. Cross-validated [CG02]. cross-validation [Rac00]. cross-validatory [Rac00]. currency [BGW00]. Current [DK05, Hsi01, Lüt01, Lüt07]. curriculum [HD08]. curse [LP08a]. curvature [Bar02, McC08a, OC05]. curve [APW06, CP04b, DRA06, DLY08, LMNT01, Moe08]. curves [CLP03, HK03, LP08c, ST02]. CUSUM [DP08a, XP02a]. cuts [Tho00]. cycle [CS06b, GRS06, MP08, PQT01]. cycles [DH05, HP06, HTV07, Par06b]. cyclical [HPU06, VI02].

daily [BKPT09, CDS09, EG06]. Dangers [STW01]. Danish [Kar02]. data [ALS01, AL01, Ant01, AC03b, Bai04, BSK03, BKP07, BSJK07, BMM09, Bha03a, Bha07, Bha08, Biø04, BGW02, BN06b, BW00, BK06, CLP09, Car07, Cha04a, Cha08, CMvS00, CMvS01, CH00, CH02, Chi07, CJ08, CKP01, Cop03, CS06b, CS07b, CS01, CHP09, Das03a, Das04, Dav02d, DLM06, Dom01, DS06, EG06, FVRR04, FFP07, GSCV06, Gre05, GH06b, HOS05, Han07a, Hay09, HCL08, Hid02, Hid08, HH04, HKvD07a, HL04, HM06, HPT02, HLR07, IvD05, Ino08, IL09, JMR09, KKP07, LSW08, LLC02, LS07, LT06a, LMR09, Mac04, Man07, Moe08, Mol08, Mor04, Mul06, MW08, Nic06, Oku09, PW08, PSS03a, PSS03b, PSS07a, PY09b, PS09a, Rac00, RL04, RR02, Rob08a, SC07, SBES05, STW01, Tod09, WW02, Wan01, Woo07, WP07, Yam08, YH00, YdJfL08, ZRT01]. data [dMF07, vDP08, vdBvdK01]. data-rich [Moe08]. datasets [DT09]. date [EM07]. de-trending [Pol00]. death [Lew03]. December [Ano06s, Ano00p, Ano01p, Ano03t, Ano07q, Ano09u]. decision [Cha00, Deh05b, KS08a, Sko07]. decision-based [Sko07]. Decisionmetrics [Sko07]. decisions [Lon04a, Man00, SV07b]. declines [Cos03]. decomposing [WvDW03]. decomposition [Kim08, KP06, MP08, Nel08, OZC08]. Decompositions [BSW04, OZC08]. defining [Flo03]. deformed [BNS06]. Deistler [PP04b]. Delay [AHR09]. delivery [HW08b]. Demand [CES08, BS08, Bas08, BIKY07, CLP03, Des00, FGHM08, HHP09, HL07, HPT08, HW08b, HS02b, LBPA02, LaF08, MMS08a, MMS^{+}08b, MT08a, MT03, PS00, SW01b, Slo08, Tho00]. Dennis [Ano05h]. densities [Bon03, HH09, HKvD07b, PQT01, RJ02, Wu03]. Density [LP04, CG02, CS06c, CDS09, EHL06, HS07, HLZ07, LS07, LT06a, McC08b, RS00, VKG09, YH06, ZBK09]. depend [IS08]. dependence [BM00a, DG08, GG07, MV02, Oul08, PS07a, SYR09, SK06a]. dependencies [BW00]. dependency [Cha02, Cha04b]. dependent [BK08, BKP07, CG08, Deo00a, Hid02, Hid08, HHS07, JST06, KH09b, LL09, RHSP05, Rac00, SW01a, Zd03, ZC01]. depends [KS08b]. derivative [Ban07, GG00]. derivatives [PSZ08]. design [Can08, Coo08, IL08a, McC08b]. designs [BR08, CJ08, IL08b]. detecting [AC03a]. Detection [Kim00, KA02, HLW09]. Detections [CCZ08]. Determination [RY02, CP04b]. Determining [NS07b]. Deterministic [Wil04, PZ05, PY09a, Phi02, RI05, VF05]. detrending [PR03]. developments [Bol01, DK05]. deviations [PM07, Stu03]. Diagnostic [Rob08b]. diagrams [Eic07]. diarrhea [JR03]. Diego [Gra06]. differ [SH07]. Difference [HK08a, CW06, EV06, HHK07]. differences [Bha03b]. differencing [Cho00, Gir00]. different [AC07, GSCV06]. differential [Kri08]. differentials [CL00]. diffusion [BN03, BZ02, BZ04, CS05, DG01, DGR06, DP08b, FF07, GC08, PY09b, Pos09, TC09]. diffusions [ASM08b, BP07a, ELX03, Xu09, Yu07]. dimensional [CNS06, FFL08, MT08a, PS08, RZ07]. dimensionality [LP08a]. dimensions [LS09a]. Direct [Ano02t, BS03, AHM^{+}03b, MSW06, MR03, SL04]. direction [BBvDvO04]. directional [LW07]. directions [ABG05]. Disability [LSW08, CvdK08]. disagreement [LS08a]. discipline [Eng01]. discontinuity [BR08, Coo08, IL08b, IL08a, Lal08, LC08, LM08a, McC08b]. discount [GM07]. discourse [Sim00]. Discrete [FHH07, Jin09, AHV05, BKPT09, Car07, CK03, CW07a, CSW00, Das05, HN07, HLR07, HP04, KS08a, fL00, MME08, PJH07, PY09b, RHSP05, Tab00]. discrete-time [BKPT09]. discretely [ASM08b, ZW09]. discreteness [KAR07]. discretized [Cop03, GC08]. Disease [JDGZ03]. disincentive [CvdK08]. dispersion [SH07]. displaced [JLS05, Jur02]. disruptions [AG06]. distance [AP02, CM07b, DG07, FGHM08, OC05, Rob05]. Distribution [DS08a, DFH06, AM08, BHPW04, Chi07, CS06a, CF07, Des06, HK08b, KP08, KP01a, LBPA02, MRS07, Mac04, PS07b, Rot02, Seo07, SSVY09, SV07a, WS09, Wu03, WP07, ZZW09]. Distribution-free [DS08a, DFH06, AM08, HK08b]. Distributional [OS06, ABD07, CS07b]. distributions [BC08a, CL09, Cop01, ET04, FS07, GGHW08, GTP06, Hor05, LMR09, MhL02, Mol08, San07]. disturbances [Bek02, DK02a, Tur00]. diurnal [KL08]. divergence [SNT06, Ull02]. diverging [BHPW04]. diversification [BLT08, PN05]. dividends [BDGT00]. Do [ASWY01, AAD05, Cap04, IS08, Lal08, Mul06]. Does [JR03, MC08, ST05b, APW06, GM06a]. domain [BC06, CM07a]. dominance [HKZ06]. double [BR08, GL02]. Dow [EM06b, SCF08]. drift [Lug03]. driven [HHS07]. drug [CS04]. DSGE [CFP09]. dual [FGHM08]. due [CES08, LS08b]. durability [CES08]. Duration [BM00a, CDS02, BV04, FG05, FG06, FHH07, GGJ04, GW02, HHS07, KT07, Lal08, RS00, ZRT01, vdBvdK01]. duration-driven [HHS07]. durations [BV04]. Durbin [BD04]. Dynamic [CRW04, GJ08, HN07, AL01, AM09, Bie07, BK06, CH09, CCFG07, Car07, CF06a, CT07b, CKL06, CS06a, CS07b, DIW00, DE07, Des00, DRA06, DLY08, DS06, DGR07, FVRR04, FHLR04, GV07, Gir00, GW04, HHK07, Hay09, HS06, Hou03, HPT02, McK04, MP04, Nic06, Oku09, PSS07a, Pel06, PP04a, PS07a, SYR09, SW08b, Tab00, Tho00, VV05, YdJfL08, Zaf04, vGK02]. Dynamics [BvS08, HH09, ADGG03, BS03, BGW02, CGGT03, DG01, DLY08, GK00, HTV06, Jon03, MRS07, RR02].

earnings [CH09, GK00, HD08]. EC2 [Ano02h]. Econom [BZ04, BT03a, Cha08, GKLT05, HLT12, KA02, Kru14, PSS03a]. Econometric [Bie07, CG08, CGM00, Ehr04, GMA08, Gew03, GM07, HS03, SMB07, BNS06, BS00, DP02, GT06, HH00, Hen01, HTV06, KMR07, fL07b, LP03, Li09, Lüt07, MCM07, MNRS03, Sim00, Sko07, Ull02]. Econometrics [Cha00, CK07, Dav02a, Die01, Dir01, Gol02, HI07, HLT08, Hec01, Li05, RU05, Zel07a, ABCD07, Ame09, Bas03a, BLvD04, BBU06, Bol01, CR08, DEF^{+}06, DP06, Dur01, Eng01, FLS07, FvD07, Gew01, GT03, Gol07, Hor01b, Hsi01, Koe00, LW01, LMPX09, Lüt01, Mar04, McA07, PP04b, SW08a, SB02, Sto01, Tau01, Vin01, Zel07c, Gra07, Ano05j, CDN08]. Economic [Lam01, ACM03, ABCD07, BN08, BS00, HL05, HTV07, IV06, IL09, MC06, PN05, Pot03]. Economics [Ano02t, Coo08, Gra06, Hec01, LP04, Maa01, OZC08, SEGG06]. economies [Zaf04]. Edgeworth [Koe00, Lin02, SvG09]. editor [ABCD07, IL08a, KH09a, Mar04, SW08a, Del09a, Del09b, Gol02]. Editorial [Ano00b, Ano00c, Ano04p, Ano02i, Ano02j, Ano02k, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m]. educated [BGKT04, Gro04]. education [Ehr04, Mar04, Mor04]. EF [BB02]. effect [CH06, CJ07, CS06b, GZ06, Gro04, HD08, HHM04, Hay09, Jur02, Lon04b, SS09, YH00]. Effective [Cop03]. effectiveness [HSWW08]. effects [AHV05, Abr00, ALS01, ABD07, Ant01, BEP07, BSJK07, BEP08, BS04, BGW02, BKPT09, BDF09, BK06, CCL03, CR05, CH09, Car07, CvdK08, DKBG04, FV09, FGS03, FFP07, Frö07, HHK07, HOS05, Han07b, HN07, HCL08, HL04, HPT02, Jun09, Kra06, LS09a, fL07b, LM08a, MR03, MW08, PS07a, RHSP05, STW01, Whi06, YdJfL08]. efficacy [HIM05]. Efficiency [But00, MC08, AD05, BL02a, DS07, DK05, FKS05, KM09, KT06b, KPS07b, ST01a, Sic05, WS09, WW08]. Efficient [AMS08, CCFG07, Che00a, CP09, FW08a, Gre06, Ino08, RZ07, RT07, Smi07, ABR07, Cop07, GG07, Hid02, IS08, KH09b, OCSN07, OT05, PSS03a, PSS03b, PSS07a, PR03, SS06, Win05]. efficiently [CMvS00]. EGARCH [Zaf09]. elasticity [HLS08]. elections [Lee08]. electric [AD05]. electricity [HN06]. electronic [HW08b]. element [Das02]. elementary [Rot02, SH07]. eligibility [CJ08]. eligible [BR08]. elimination [fL07c]. Empirical [Bat03, Che03, DIN03, EvdB07, GLR03, Man00, Xu09, And04, ABR07, BS06, Can08, CG07, CC07, CS04, Dah09, DW01, EL03, GK00, GRS06, Gra05, GS08, Hec01, Ots08, Oul08, RS02, Sin01]. Empirically [HS00]. empirics [BDW07]. employers [Cap04]. employment [BR05b, HSWW08, LHMC09]. Encompassing [Wes01, CK02, CK07, CM01]. Endogeneity [CDvD07, Lee07, Che09, HKL05, HH04, jLV06]. Endogenous [Lew07, BP07b, CP02, Das05, FL03, HL09, KL08, Kim04, KPS08, Kim09, MT08b, OK00, PSS07b, Rot09, Stu03]. endogenously [KT09]. energy [BIKY07]. Engel [CLP03, LP08c]. engineering [Che03, GT03]. Engle [Zel01]. Engler [Bas03b]. English [BP06, Bre07, BP07c]. enrollment [CL09, Mal08]. entropic [KS02]. Entropy [Gol02, MR02, RJ02, GP02a, Gol07, HIJS07, MGM08, PB09, RM07, SP01, Ull02, Wu03, WP07]. entropy-based [HIJS07]. entry [CR05, CS04, Li05]. environment [Moe08]. epidemiological [FGS03]. equal [CM01, CW07b]. equality [Das03b, Lav01, LMR09]. equation [HN05, HS02b, Li08, MK06, Phi00, RT06, SWZ08, WZ03a, dLJ06]. equations [And06, ABR07, CP02, DG01, GL02, JM07, KP04a, Kri08, OK00, SU08, Tur00]. equilibrium [Bie07, CS07b, DGR07, FVRR04, GRS06, Hen06, Koo01]. equilibrium-correction [Hen06]. equity [BL02a, DF09]. Erasmus [Gra07]. ergodicity [CSW00, DEG01]. erratic [SM02]. Erratum [Ano05i, HLT08, Zel07a]. Error [Ano07m, BSJ01, BSK03, BBP06, BU07b, BSEY01, BLS06, CDS02, Cho02, Dav02d, ET07, ET04, HS02a, HN05, HT03, Hu08, JJ00, KL08, KKP07, LC08, PSS00, RT06, RSU96, RU05, Sai05, Seo06, Seo07, SI04, WZ03b, Wan01]. error-correction [HS02a]. errors [Cai07, CW06, CM07b, Des06, HMS05, Has01, JPR04, KPS07b, KH09b, LL09, Li02, LH04, MMS08a]. errors-in-variables [KPS07b, Li02]. Esfandiar [Hor01a]. essays [Kri01, Lüt01]. Establishing [Dav02a, Dav02b]. establishment [Hsi09]. estimated [GL01, WS09, XL04, dJ01]. estimates [CP04a, CG02, HLS08, KP08, Per00]. Estimating [AHV05, Bai04, BEP07, BEP08, BW02, BZ02, BZ04, CL09, Car07, CK00, CH03a, CZ07, Cho00, CT07b, Dav02d, DF09, FGHM08, FL03, GO05, GH06b, HPT08, JLS05, Jur02, KL08, MN00b, MNRS03, MT08a, Mor04, PQ06, PY09a, RS00, Slo08, VV05, ABR07, Ban07, CK03, Dur06, JM07, MMS^{+}08b, SA04, WZ03b]. Estimation [AC07, Ano01f, Bon03, CMvS00, CF06b, CD06, CF06a, Cop01, CLP03, Dav06b, DLM06, Dom01, DSS07b, GP02b, HOS05, HK03, HL07, HH04, HP01, JM07, KP04a, KPS08, Kri08, KT06b, KW06a, LP08c, Li08, MS09b, PWT08, SW07a, Sin01, Tod09, Aba03, ADG07, Abr00, ALS01, AL01, Ant01, Art04, AP02, BS03, BN03, BP07a, BB02, Ber01, BW00, BP06, Bre07, BP07c, But00, CW08, CCFG07, CG08, CFS02, CV03, CM07a, CDPX07, Che00a, Che00b, CCZ05, CES08, CP09, CH03b, CIN07, Cho02, CHO08, CP07a, Cop03, Cop07, DG01, Das03a, Deo00a, DHL06, DF08c, DP02, DIN03, ELX03, FFL08, FV09, Frö07, Gol01, GLS02, HHK07, HIJS07, HML02, Hau03, HCL08, Hid02, HL04, HM06, HPT02, HW06, Hu08, HR07, IS08, Ino08, KS08a, KP07, KP01b]. estimation [Kha01, KT07, KS02, Kom05, Kra06, Kru08, Kru14, Kue01, KW06b, fL07a, fL07b, fL07c, LS08b, Lew00, Lew07, Li02, LHZ03, LH04, LMNT01, LS09b, MFY07, MR03, MJ05, Moo04, Mos01, Mül07, Mul06, MT03, MW08, OFRP05, Ots08, PS08, PSS03a, PSS03b, Par03, PSS07a, PSZ08, PT07b, PPC04, PS07a, PY09b, Pol00, Pos09, PS09b, RHSP05, RL04, RH03, Rot09, SW01a, Sak07, SS09, SF01, SCF08, SP06, Shi07, SS00, SO04, SL04, SU08, SS06, TC09, Vil06, VKG09, WW02, Wan01, WW08, Whi06, Woo07, WP07, XP08, YH06, Yu07, Zaf09, ZBK09, ZZW09, dJ01, vAJM02]. estimator [BU07a, BK06, CS07a, Cop01, DG07, GG07, GL00a, GL02, HI03, HI07, Han07a, HT03, Hor02, jLV06, LP08b, LS07, Mac04, MFY08, PS07b, Seo07, SL07a]. Estimators [FMRT09, ADG09, ASM08b, AHT07, And05, AG09, BU07b, Can08, CI08, CDS09, Das04, Das05, DGR06, DS06, Fit04, FS08b, FNS04, GV07, GP02a, Gre05, HS07, HW06, Ink00, IS06, JL08, Kar02, KH09b, LS09a, LP08b, Lin07, Lin02, MN00b, MJ05, Muk07, NS03, OK00, RV02, Rei07, RSU96, RU05, RT04, RT01, Sic05, ST05b, Win05, YdJfL08, dJ02b, dJ02a]. Euclidean [ABR07]. Euler [Li08]. European [LP04]. Evaluating [BN06a, HSWW08, LT02, LW07, Ull02]. Evaluation [AB01, CS07b, BDW07, Deh05b, DG08, HL05, JKS03, MS04, Pal09, Pes04, vdK08]. events [Bow07, GRZ07]. Evidence [BGW00, CH09, CGS08, FS08a, PQT01, SH07, Bha03b, BS04, CR05, Cos03, EL05, FGS03, Gas03, GLS02, Jon03, KPS07a, MN00a, Mor04, Vil01, YH00, AAD05, BS04, HLZ07]. Evolution [LS08a]. evolving [KV00]. ex [Mos01]. ex-ante [Mos01]. EX1 [Ano01r, Ano02s, Ano02q, Ano02r, Ano02p, Ano03x, Ano03w, Ano03v, Ano03y, Ano04s, Ano04u, Ano04t, Ano04v, Ano05-28]. EX2 [Ano01r, Ano02s, Ano02q, Ano02r, Ano02p, Ano03x, Ano03w, Ano03v, Ano04s, Ano04u, Ano04t, Ano04v, Ano05-28]. EX6 [Ano03y]. Exact [Bek02, Des00, DK02a, DV09, FS08b, Lug03, Lug06, NS07b]. Exactly [AM08]. Examining [JL08]. examples [Sim00]. Excess [Che09]. exchange [BGW00, Hon01, HLZ07, Jin09, LT06b, OP04, Wri08, Yan06]. exclusion [Das03a]. executive [EL03]. existence [LM02]. Exogeneity [dLJ06, BBU06]. exogenous [PSS00]. expansions [SvG09]. expectancy [MNRS03]. Expectations [CFK06, Dom01, JL08, Leh09, LS07, MN00b]. expected [CW08]. Expectile [KYH09]. expectiles [DH09]. expenditure [CMvS01, JDGZ03]. experiences [DG08, HIM05]. experiment [BR05b, CH09, Fré09, Mal08]. Experimental [ABW07, HL05, CR05]. experiments [HKL05, Lee08, Whi06]. Expert [FGS03]. Explaining [vDP08]. explanatory [FL03, HL09, Kim04, Kim09]. exponential [Kaw06, LP07, Zaf09, ZZW09]. exponentiality [HK08b]. exponents [SL04]. exposure [FGS03]. expressions [XL04]. extended [ADG07, Das03b, Lal08]. extension [DSK07]. External [DF02]. Extracting [CMP09]. extreme [CF07, DP02, HLW09, SV07a].

F [Bas03b]. Factor [Sen04, BS03, BN06b, CRW04, DRA06, Dur06, FFL08, FHLR04, GM07, JKS03, Kle09, KLM08, fL00, Moe08, PP04a, SF01, SCF08, Tho00]. factor-augmented [Moe08]. factors [BN06a, CFK06, GRS06, GTP06, KP02, fL07b, MP04, OP05]. facts [Dir01]. failure [GS08]. families [Das03b]. Family [Bha03b, DSS07a, FG06, LM02]. Fast [OCSN07, HS06]. fat [JPR04]. fat-tails [JPR04]. FDI [BEP07, BEP08]. fears [Bat00]. features [AIV06, HPU06, VI02, EM06b]. February [Ano01q, Ano02s, Ano03u, Ano05-28, Ano06u, Ano09s]. fecundity [BvS08]. Federal [WW08]. feedbacks [BK06]. Feeders [Bas03b]. feeding [Vin01]. Fellows [Ano08m, Ano08n, Ano05j]. female [HSWW08]. fertility [BvS08]. Festschrift [PP04b]. field [AT07, LP04]. fields [DGR03, JP09a]. figures [Dir01]. filtering [Dur06, PFC08, Wil04]. filters [Pol00]. finance [BN06a, BHPW04, DW01, DEF^{+}06, GMA08, GGHW08, Gra05, Gra06, Kri08, SEGG06]. Financial [Bol01, CFK06, Eng01, AG06, BK08, BV04, BS06, CR05, GT03, HLW09, SV07b, Tau01, ZRT01]. findings [WZ03b]. Finite [ACMP08, BU07a, BIKY07, CHJ09, DJ06, SB02, DS06, DT00, DK02b, DKBG04, Duf06, LS09a, LLC02, McC07a, OK00, Ros08, Sto09, Win05]. Finite-sample [DJ06, DT00, DKBG04, Duf06, LS09a, LLC02]. finite-state [McC07a]. firm [AP02, KS08b, Sic05]. firm-specific [Sic05]. firms [Dav06b]. first [Li05, Maa01]. first-order [Maa01]. first-price [Li05]. Fiscal [JLWY08]. Fisherian [KY09]. Fisherian-Bayesian [KY09]. fit [ASBS01, CGM08, Deo00a, HK08b, HZ07b, HZ07c, Hid09]. fitted [Rot02]. five [Ame09]. Fixed [FV09, MW08, Abr00, Car07, HHK07, Han07b, HCL08, HL04, HPT02, fL07b, PS07a, YdJfL08]. fixed-effects [Abr00]. flexible [BBvDvO04, Des06, GT09, HL07, HKvD07b, vSDG02]. flow [Cha04a, Cha08]. fluctuations [GZ06]. food [LBPA02, LaF08]. forecast [CM01, ET04, FW08a, GT09, Han08, IL09, LS08a, San07]. Forecasting [BN08, BPT01, CF01, CC04, DGR08, DHL06, DW01, DL06, Gra07, Moe08, PP04a, AT06, BRY08, CN06, DSS07b, Gew01, GGHW08, IK06, IL09, MSW06, Sch05]. Forecasting-looking [Gra07]. Forecasts [GT09, ASM08a, EHL06, FW08a, Gos02, Han06, Hen06, HLZ07, Mar08, PT07a, PT05, Wri08]. foreign [BGW00, HLZ07, Yan06]. form [BK07, DP08b, GK04, Li08, Yu07]. forms [KT07]. formulas [SH07]. Forum [Ano01j, Hsi01, Kri01]. forward [Gra07, HS02b]. forward-looking [HS02b]. four [AAD05]. Fourier [FMRT09]. fractional [AV09, CH03a, CN06, Dav02c, HR07, MR01, NS07b, PW08, RY02, RI05, Rob05, SP06, SP03]. fractionally [BH02, Dav06a, DG02, Shi07, TC00]. frailty [BvS08]. framework [BZ06, EE00, HN00, KSS03, KP07, KP09a]. free [AM08, DS08a, DFH06, FS07, HK08b]. frequencies [GSCV06]. Frequency [CM07a, Bas08, BPT01, BW00, BC06, Cha04a, Cha08, MMS^{+}08b, Tod09, dMF07]. frequency-domain [BC06]. frequentist [NS03]. Frisch [Koe00]. Front [Ano02j, Ano02k, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04p, Ano04l, Ano04m, Ano04n, Ano04o, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano06o]. frontier [CFS02, Gre05, GS04, KT06b, MFY07, MFY08, WS09]. Frontiers [GT03, FKS00, FS07, FS05, GO05, KPST07]. full [LHMC09]. full-time [LHMC09]. Fully [SO04, Des00, MPR00]. function [Cho00, FNS04, KP09b, KW06b, Lee07, LP08d, LS09b, Sin01, SW07b]. function-based [SW07b]. Functional [CDXW06, CLP09, RR02, Xia09a, BN03, Dav02a, Dav02b, HY05]. Functional-coefficient [CLP09, Xia09a]. functionals [HM06]. functions [AT05, AP02, Cop04, ET04, FGHM08, FS08a, KT06b, LT06a, McC08a, Mos01, OC05, PS08, PY08b, RL04, Wha00]. Further [DT07, GL00a]. future [Bol01, Gra01, Gra05, HIM05, Hsi01, Lüt01, Phi01]. futures [Bat00].

Galton [Koe00]. game [dP09]. games [Dav06b]. gamma [DSS07a]. GARCH [BRT03, BC08a, CDPX07, CI08, EZ01, Gos08, Haf08, Hil05, Kaw06, LL09, LM02, LT02, MPR02, Nak00, PWT08, SS06, Tse00, Yan06]. GARCH/IGARCH [Lin07]. gasoline [RT06]. gauging [BZ06]. Gaussian [AL04, ALM06, Art04, BS03, CM07a, rMCGJL08, GS06, Shi07, SS00, VKG09, Zd03]. GDP [APW06]. gender [Bha03b]. General [Lüt07, Rob03, BR05a, Bie07, CCFG07, CS07b, DGR07, Duf06, ET04, Hu08, KT07, LS09b, PS09a, Sch04, WZ03b, Woo07]. General-to-specific [Lüt07]. Generalised [Tur00]. Generalized [Ber01, BD04, EV06, GS08, Han07b, KP06, Muk07, RS02, vAJM02, Abr00, Can08, CZ07, DSS07a, FHLR04, GL01, HI03, HI07, HIJS07, HK08a, IS02, LH04, Lie01]. Generalizing [Kle07, Zel07a, Zel07b]. Generating [DS05]. getting [GSCV06]. Geweke [LW01, Zel01]. Gini [Dav09]. Glejser [Im00, MS00]. Global [DLY08, BKN09, BL02a]. GLS [PR03, SO04]. GMM [ALS01, AL01, BB02, BM05, But00, Can07, Can08, CV03, DS06, GTU05, Ink00, IS06, fL07a, fL07c, Nev02, PS09a, RT01, Wan01, Win05, WP07]. Goodness [ASBS01, CGM08, Hid09, Deo00a, HK08b, HZ07b, HZ07c]. Goodness-of-fit [ASBS01, CGM08, HK08b, HZ07b, HZ07c]. goods [RR02]. government [DL06]. graduate [MZ04]. Granger [Dei01, DJ06, Eic07, HLW09, McC07b, MC06, SEGG06, Zel01]. grids [HW08a]. group [BR08, FHH07, fL07b]. group-level [FHH07]. grouped [DLM06]. groupwise [Bek02]. Growth [MRS07, APW06, AP02, Dur01, Pes07]. Guest [KH09a]. guide [IL08b].

HAC [KP07]. Hansen [ASM08b]. Hansen-Scheinkman [ASM08b]. harm [DM04]. Hausman [CHO08, Das03b, WWM01]. Hausman-type [CHO08, Das03b]. hazards [HL04]. health [AHM^{+}03b, ACM03, Bha03a, CGS08, DF08b, Pot03, SW01b, WvDW03]. healthcare [MT03]. Healthy [AHM^{+}03b]. heavy [CDPX07]. heavy-tailed [CDPX07]. Heckman [Dei01]. hedging [BCC00, GG00]. HEGY [BT04a]. heterogeneity [BM00a, Che09, FHH07, Gre05, Tab00]. Heterogeneous [DS08b, MM08, Bar07, Cho02, Dav06b, IPS03, McK04, SW01a, WKV03]. heteroscedastic [Bek02, CCZ08]. heteroscedasticity [BT01, Deo00b, Lew00, Lug03, Muk07]. heteroskedastic [CK03, DFH06, MMS08a, SF01, SCF08]. heteroskedasticity [CK00, CP07b, DKBG04, GK04, Im00, Ink00, Par02, PB09, Seo07, Vil01]. hidden [CP04b]. High [FFL08, BPT01, BW00, CNS06, MMS^{+}08b, MT08a, PS08, RZ07, Tod09, dMF07]. high-dimensional [MT08a, PS08, RZ07]. high-frequency [BPT01, BW00, Tod09]. Higher [ALM06, HK08a, RH03, XP02b, Ehr04, GP02a, IS08, Mar04, Mor04, PQT01]. Higher-order [ALM06, RH03, GP02a]. Highly [FRW05, Gos02, Mül05]. hint [GG00]. historical [CS07b]. histories [Jur02]. history [Coo08, Zel07d]. Hodges [LP08b]. homogeneity [GG00, PY08a]. homoskedasticity [BBP06]. Honor [Ano07t, PP04b, SEGG06]. honoring [CDN08]. House [Lee08]. Household [DF08b, Bha07, CKP01, CLP03, HW08b, LP08c, PS00]. household-level [HW08b]. households [vdKW08]. housing [CMvS01]. hybrid [ADGG03]. hypotheses [AD07, Bas08, CK02, CK07, DV09, HY05, HK08a, Joh02, Joh06, KP09b, RM07]. hypothesis [CW06, Deo00b, EV06, HLT07, HLT08, Has01, HS00, Hor01b, Jan05, RT07, ST01b, SK06b, Tay05].

i.i.d [ABD07]. Identification [BDF09, BP06, BD07, Das03a, HM06, Hu08, fL07b, MM07, Man00, Mul06, SF01, BP07c, Can07, CS07a, CDvD07, CW07a, CFP09, GS08, Jun08, JP09b, Kle07, McA08, Mol08, PS08, PP04b, Sic05, Tab00, vS04]. identified [GH09, MPR00, NS07a, Ros08]. identifier [HD08]. identifying [GH09]. IFC [Ano02j, Ano02k, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04p, Ano04l, Ano04m, Ano04n, Ano04o, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano06o]. IGARCH [Lin07]. ignorable [HD08]. ignores [KW06b]. ignoring [DM04]. II [Kim09]. Impact [BNS06, OP05, EM06a, Reb04]. impacts [DS08b, WKV03]. implement [vGK02]. Implementation [JST06]. implications [ABD07]. Implied [ABR07, ASL00, ASBS01, BPT01]. importance [KSC09, RZ07, VI02]. important [CR05]. imposing [HHP09, OC05]. improve [MJ05]. improved [DS06]. improvements [ALM06]. Inc. [Bas03b]. Incentive [LM08a, CR05]. Incidental [MPP07, Lan00, Mac04, PS07a]. Income [CF07, Dom01, LBPA02, Wu03, vdKW08]. incomplete [HM06]. Incorrect [AG09]. Increasing [HM08]. incremental [BPT01]. independence [MGM08, SW07b]. independent [BR09, Gir00, Hor05, LPV00]. Index [Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano01g, Ano01h, Ano01i, CP03, Ano01e, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano03c, Ano03a, Ano03d, Ano03e, Ano03f, Ano03b, Ano04b, Ano04c, Ano04a, Ano04d, Ano04e, Ano04f, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, BR05a, BPT01, Che09, Dav09, Hil06, Kha01, RR02, SH07]. India [BvS08, Bha03b, JR03]. Indian [BMM09]. indicator [IV06]. indicators [EG06, dJAS07]. indices [IV06, Oul08, SR02, ZC01]. Indirect [DGR07, SCF08]. Individual [BGW02, ALS01, HOS05, SS09, vDP08]. indivisibilities [Dav06b]. Industrial [GO07]. industry [Kar02]. inefficiency [AP02, KS08b, KT05, KW06a, KW06b]. Ineligibles [BR08]. inequalities [KT09, Ros08, WvDW03]. inequality [Bha07, Bie02, CL09, CF07, DF07, MS09b, SvG09, ZC01]. infant [Bha03b]. Inference [Bha07, Bha08, BH02, GS06, HT09, KT09, Pal09, YH00, dP09, AM08, Bek02, BL08, Bha05, Bie02, CH06, CH08, CHJ09, CP04b, CM07b, Dah09, DF07, Dav09, DH00, Des00, DP02, DGR07, DT00, DJ06, Duf06, DPR06, DT07, FSZ04, GP02b, GW05, GS04, Han07b, HML02, HMV06, HN05, Ino08, JM07, KS08a, Kle04, Koo01, KT06a, Kru08, Kru14, LC08, Leh09, LP04, McC00, MJ05, Mul06, NS07a, OCSN07, Ots08, RT01, Sar03, SW07a, SWZ08, Smi07, Ull02, Xu09, Zd03, ZW09, Zhe01, ZC01, vAJM02]. Inferential [HLS08]. infinite [Has01]. inflated [HZ07a]. inflation [CP04b]. Information [Cla07, DSS07a, Gol02, Gol07, GLS02, HIJS07, Kar02, LBPA02, SR02, Zel02, And05, BPT01, CFK06, CGS08, CP04b, GL00b, Hu06, Kim02, LPV00, MM07, MhL02, NS07a, Nev02, RT06, Smi07, Zel07d]. Information-based [Kar02]. Information-theoretic [GLS02, Nev02]. informational [ABR07]. informing [SW08a]. Infrastructure [DSK07]. inhomogeneous [ELX03]. initial [EM06a]. innovation [KLN02]. innovations [BRT03, CDPX07, SV07a]. input [KT06b]. input-oriented [KT06b]. inputs [LP03]. Institute [Gra07]. instrument [Kle07]. Instrumental [CH06, CIN07, CH08, Che07, Cho02, Das05, Sak07, Aba03, AM08, BL08, CDXW06, CL09, CHMR08, DIN09, HHK07, HKvD07a, HKvD07b, Hu08, Jun08, KZ03, Kue01, Lew00, Lin02, MW08, Par03, PPC04, SL03, SK06a, WW02]. instruments [AMS07, AS07, AM08, AMS08, CDvD07, DT07, MPS09, Mor09]. insufficient [Das03a]. Insurance [LSW08, AAD05, CvdK08, DF08b, Jur02]. intangible [Dub07]. integer [CDSV08]. integrated [AL04, BZ02, BZ04, BH02, CP03, CP07a, Dav06a, DG02, HW06, Par06a, PY09a, Shi07, TC00, XP02b]. integration [BBvDvO04, HW08a, KV00, SP06]. intellectual [GO07, McA07]. Intensity [Bow07, KLM08]. inter [GW02]. inter-transaction [GW02]. interaction [KS07]. interactions [BD07, CT07b, DLY08, fL07b, PSS07b, dP09]. interdependence [GW02]. interdependent [BEP08]. interest [CFK06, Das02, GT09]. interface [ABCD07, DEF^{+}06]. international [HMSA03]. Internet [HH00]. Internet-based [HH00]. interrelated [KP04a]. intertemporal [GLR03]. Interval [CDPX07, Han06, CS06c, WP07]. intervals [CCL03, ES01, IS02]. interventions [DG08]. intra [EG06]. intra-daily [EG06]. intraday [HLZ07]. Introduction [Bas03a, CDN08, GH06a, ABCD07, Del09a, Del09b, IL08a, KH09a, SW08a]. introspective [ABW07]. Invariance [Kem01]. Invariant [Kle04, AMS08, SS01]. Inverse [Woo07, LS07, PY08b]. investigation [BS06, BDGT00]. investor [WW09]. issue [ABCD07, Ano07m, CDN08, HL05, IL08a, SW08a]. Issues [Gas03, CCL03, Flo03, Gew03, HMSA03]. iterated [GL00a, MSW06]. IV [AMS07, AMS08, Cha02, CS07a, DT07, Frö07, Kle07].

J [Ano05h, BZ04, BT03a, Cha08, DM02, Dei01, GKLT05, HLT12, KA02, Kru14, PSS03a, SEGG06, Zel07a]. J. [Dei01, WWM01]. Jacobi [GJ06]. January [Gra06, Ano00n, Ano02m, Ano05x, Ano06p, Ano07o, Ano08q, Ano09o]. Japanese [YH00]. JE [Kri01]. Jeffreys [CP02, Kle04]. job [BC08b, HSWW08]. Joint [BBP06, Esc08, AD07, Chi07, EHL06, HTV06, MhL02, Yam08]. Jones [EM06b]. Journal [CK07, Dav02a, Dir01, HI07, HLT08, RU05, Ame09, Ano05j, CDN08]. July [Ano00l, Ano02q, Ano03v, Ano05v, Ano06r, Ano09r]. jump [BN03, Pos09, Yu07]. jump-diffusion [BN03, Pos09]. jump-diffusions [Yu07]. Jumps [BKPT09, ABD07, BNS06, BLT08, Das02, JO08, Tod09]. June [Ano00q, Ano01n, Ano02o, Ano03w, Ano04v, Ano06v, Ano08s].

Kalman [PFC08]. Kernel [DFG04, ASBS01, Che03, HS07, LMNT01, RH03, ZBK09]. Kernel-based [DFG04]. Kolmogorov [Cop04]. KPSS [dJAS07]. Krueger [HKvD07a]. Kullback [DG07, SNT06]. Kurtosis [BRT03, RJ02].

labor [BGKT04, DG08, EvdB07, Gro04, Hou03, JS08, Jur02]. laboratory [HKL05]. labour [vSDG02]. LaLonde [ST05b]. Large [GM06a, CHP09, DGR08, DK02b, HI03, HI07, Han07a, JP09a, PY08a, Sen04, SCF08, SWZ08, Stu03, YdJfL08, Zaf04]. last [DF08a]. latent [BN06a, BV04, DRA06, GLR03, JJP07, KLM08]. lattice [Hid09, SC07]. laws [JP09a, LPS05]. lead [FGS03]. leading [IV06]. Learning [CGS08, Fré09, CS04, LS08a]. Least [Han08, Sch04, And05, CN06, GL00b, Han07b, HK08a, HW06, IS08, PS07b, SL07a, Wil04, YH06]. Least-squares [Han08, HW06]. leaving [GH06b]. lecture [Hsi09]. Leibler [DG07, SNT06]. lessons [Hoo03]. level [FHH07, HW08b]. levels [GKLT03, GKLT05, Sic05]. leverage [ABD07, BKPT09, OCSN07, Yu05]. Levinson [BD04]. Lévy [BNS06, ZW09]. LIBOR [JKS03]. licensing [HL07]. Life [Coo08, Cos03]. Likelihood [HW08a, PS09b, WZ03a, And05, ABR07, BU07a, Biø04, Can08, CG07, CHS07, CC07, DIN03, ELX03, FPW04, FdV07, FRW05, GW04, GS08, HPT02, JJP07, KS08a, Kim02, Kom05, Kru08, Kru14, KPST07, Lin07, OCSN07, Ots08, PSZ08, Plo04, RHSP05, RS02, Sim00, Xu09, Yu07, YdJfL08]. Likelihood-based [PS09b, Xu09]. Limit [PM07, Cor00, Dav02a, Dav02b, JP09a, KP08, LPS05, NS07a, PSS07b]. Limited [Kim02, RT06, And05, JST06, MhL02, RHSP05]. LIML [GL00a, OK00]. Lindley [Kle04]. linear [ALS01, Bek02, Cho05, DM02, EM07, FFP07, For05, FL03, HK06, Hu06, Ino08, JL08, JX05b, Kim00, KA02, KS02, Kle09, LH04, Lug06, MFY07, SYR09, WW02, Win05, Yam08, Zaf04, vGLP00]. linearity [DS05, Gos08]. linearized [Bie07]. lines [Zhe01]. link [IV06, vdK08]. links [KY09]. list [Ano08m]. listing [Ano07m]. litigation [Bas03a, Bas03b]. living [Oul08]. LM [BBP06]. Local [Dah09, For08, GL00b, HS07, Jun09, PSZ08, SP06, SU08, ADG07, BSEY01, CL09, CT07b, DP08a, Frö07, GZ06, KPST07, Lin07, MFY07, NS07b]. locally [Dah09, Plo04]. location [Che00b, CM07b, Fit04, Gro04, MS09a]. location-scale [MS09a]. locations [HIM05, PSS07b]. Log [ACMP08, HMV06, Phi07, SP03]. Log-ACD [ACMP08]. log-periodogram [HMV06, SP03]. logistic [Lon04a]. logit [BHH08, Dub07, JMR09, LG02]. logit-probit [BHH08]. loglinear [vG01]. Long [BP08, DT02, DI01, HHK07, Phi09, ADG09, ALM06, Art04, BCC00, BU05, BW00, BC06, BC08b, CH09, CG08, CD06, DS05, DV05, DHL06, DPR06, EM06b, FMRT09, GKLT03, GKLT05, HN06, HMV06, HHS07, Laz05, LP08d, Liu00, MV02, MM08, Mül07, RH03, Zd03, vDFP02]. long-memory [ALM06, BW00, DV05, DHL06, FMRT09]. long-range [CG08, MV02, Zd03]. Long-run [BP08, ADG09, BC06, BC08b, EM06b, HMV06, Mül07]. long-term [BCC00, CH09]. longitudinal [CH02, Mor04]. looking [Gra07, HS02b]. Lorenz [HK03, ST02]. loss [BC08b, ET04, MJ05, PT07a, WZ03b]. lots [FTC03]. low [vdK08, vdKW08]. low-income [vdKW08]. LS [BK06]. Lyapunov [SL04].

M [RH03]. M-estimation [RH03]. Maasoumi [Hor01a]. Macro [Sto01, vdBvdK01]. Macro-econometrics [Sto01]. Macroeconometrics [Gra01]. macroeconomic [BM06, BS06, CFK06, GLS02, HTV06, KPS07a, MSW06, Phi01, YH00]. macroeconomics [BN06a, DIW00, DW01, DEF^{+}06, Hoo03, Pos09]. macroeconomists [CS01]. macroeconomy [DRA06]. main [OP04]. major [Arc04]. makers [SW08a]. Making [KPS07a, CR08]. malnutrition [WvDW03]. management [ASL00, GMA08]. Mandatory [Mat08]. Manfred [PP04b]. Manifesto [Dur01]. Manipulation [McC08b]. manufacturing [BR05a, FS08a]. many [AS07, MT08a]. March [Ano00m, Ano01l, Ano04u, Ano06t, Ano07n]. Marginal [FdV07, Esc08, FV09, IvD05]. Market [BL02a, BU05, Bas08, Bat00, BM06, Bow07, DG08, Dub07, FFP07, GZ06, HMSA03, Jur02, Liu00, MR02, RT06, WW09]. market-response [FFP07]. markets [ASWY01, AG06, BL02a, HLW09, JLWY08, Jon03, Sen04]. Markov [ASY06, Car02, CNS02, CP04b, CSW00, Des06, FZ01, KS08a, Kim04, KPS08, Kim08, Kim09, LG02, McC07a, Nak00, SWZ08, SNT06, Tim00, dMF07].

Markov-switching [Car02, FZ01, Kim04, Kim08, Kim09, SWZ08, SNT06]. Markovian [DT00]. markups [DF08c]. martingale [ABD07, CW06, Deo00b, EV06]. Masking [CW07a]. match [KPS07a]. matched [LSW08]. matching [BS04, Deh05a, ST05b]. Maternal [LHMC09]. matrices [FSZ04, Kle07]. Matrix [Kaw06, DS06, ET07, FFL08, Han07a, Kar02, LP07]. matter [GM06a, MC08]. max [FS08b]. Maximum [ELX03, GW04, HPT02, Kru08, Kru14, Moo04, PB09, RHSP05, SP01, AL04, And05, BU07a, Biø04, GP02a, Hor02, JJP07, Kom05, KPST07, Lin07, Wu03, WP07, YdJfL08]. may [MPS09, Ano00k, Ano01k, Ano02n, Ano03r, Ano04t, Ano05y, Ano08p, Ano09q, GM06a]. McFadden [HKZ06]. McFadden-type [HKZ06]. MCMC [CH03b, JJP07]. ME [BB02]. mean [BU07b, BK07, CW06, Esc08, Hay09, JX09, MS09a, RSU96, RU05, SS00]. mean-nonstationarity [Hay09]. mean-variance-skewness [MS09a]. means [DLM06]. measure [Dah09]. measured [BC08b]. Measurement [HMS05, AD05, Bie02, CDS02, CF07, DK05, FSZ04, KL08, LH04, Wan01]. measures [CDS09, DSS07a, DF07, FKS05, LBPA02, SvG09, SV07a, Ull02, Zhe01]. Measuring [KT05, MvD07, Slo08]. Median [Gos02, Sak07, SS01]. medical [JDGZ03, Mal08]. meeting [Ano02h]. meets [HK08a]. memory [ADG09, ALM06, Art04, BU05, BW00, CD06, DT02, DS05, DV05, DHL06, DI01, FMRT09, GKLT03, GKLT05, HN06, Laz05, LP08d, Liu00, MM08, Phi09, Rob01, RH03, Zaf07, vDFP02]. men [GK00]. mergers [FTC03]. method [AB01, Ban07, DS04, GG03, HI03, HI07, HIJS07, HS06, HN00, IS02, KMR07, fL07c, OT05]. methodological [Rob03]. methodology [Gra06, Hen01, SEGG06]. Methods [SWZ08, BBvDvO04, BCY00, CNS02, CW07a, DW01, DP06, Dur06, Eng01, FLS07, FPW04, HLS08, HKvD07b, Kru08, Kru14, LHZ03, LMPX09, LMNT01, MK06, MSW06, PW08, SA04, SB02]. metric [RM07]. micro [vdBvdK01]. microaggregation [SS09]. Microeconometrics [Hau01a, WWM01]. mid [Cos03]. mid-life [Cos03]. migration [Gro04]. minimal [Rei07]. Minimax [Man07, Sto09]. Minimax-regret [Man07]. minimization [dJ02a]. Minimizing [EM06a]. minimum [AL04, LS09a, PY08b]. mining [STW01]. mis [CD06]. mis-specified [CD06]. misclassification [Hu08, Ram02]. misclassified [Mol08]. mismeasured [FL03, Hu06]. missing [DT07, IV06, Man07, PS09a, Woo07]. misspecification [BM00b, BSEY01, CF06a, CS06a, JL08, Sch05]. Misspecified [Car02, Ink00, AC07, HI03, HI07]. mixed [BHH08, HLR07, fL07a, fL07c, LMR09]. Mixing [FMZ08, GK07]. mixture [Cop01, Lie01]. Mixtures [GGHW08, MS09a, VKG09]. ML [BB02]. MLE [But00, Kim09]. MM [BB02, BK06]. MMC [JST06]. Mobility [FSZ04, Bie02, MRS07]. MOD [Cop01]. Model [BDW07, FS07, Wri08, AL04, Abr00, ACMP08, ALS04, AL01, AG09, BSJ01, BBP06, BV04, BGG08, Bek02, BH07, BSEY01, Ber01, BKPT09, BHH08, CC04, CP02, CMvS00, Che00b, CC01, CF06a, CZ07, CHS07, Cho02, CN06, CGM00, Cop01, Cop07, CS04, DS04, Dav02c, DHL06, DS06, DF09, Dub07, EM06b, EG06, FFL08, FLS01, For05, FHLR04, GLR03, GV07, GSN08, GP02b, Gos08, Gre05, HN06, HOS05, Han03, HZ07a, Hau03, HLZ07, HKvD07a, HTV06, HL04, HL09, Hou03, HLR07, Hu06, HS02b, IvD05, Ink00, JJ00, Joh06, Jun08, Jun09, KLM08, Kru08, Kru14, LG02, LS08a, LSW08, fL00, LS08b, LL09, Lew00, Lew07, LP03, Lie01, Lon04a, LT06b, MM08, MPR00, MN00b, MJ05, Moo04, MT03, MT08b, PB09, Per00, Rac00, SC07, SYR09, Seo06, SL07b]. model [SNT06, SI04, Tse00, Tur00, WW02, Wes01, Yan06, Yu05, YS07, ZRT01, vDFP02, vSDG02]. model-selection [Rac00]. Modeling [CJ07, FPW04, FF07, GW02, Liu00, Jin09, KPS07a]. Modelling [Ano05u, BU05, Bow07, Cla07, CHP09, GMA08, Gir00, HZ07a, Lüt07, Sko07, Tho00]. models [Aba03, ALS01, AC07, ABD07, And02, AG08, AL01, AC03b, Art04, AM09, AHR09, BN01, BRT03, BC08a, BK08, BSK03, BEP07, BS03, BN03, BMM09, BU07a, BU07b, BIKY07, BS06, Bha08, Bie07, BGW02, Bow07, BD07, BLS06, BK06, CDXW06, Cai07, CLP09, CGM08, Car02, CCFG07, Car07, CG08, CT07a, CDPX07, CP03, CMvS01, CK00, Che00a, CK03, CCZ05, CF06b, CD06, CF06a, CG07, CHS07, CP09, rMCGJL08, CGGT03, CH06, CIN07, CHJ09, CH00, CNS02, CH02, CNS06, CW07a, CP04b, CM01, CW07b, CT07b, CFP09, CS07b, CRW04, DGR03, Das03a, Das04, Das05, DM02, Dav02d, DE07, DS08a, Des00, DP08b, DLM06, Dis08, Dom01, DP02, DIN09, DGR07, DMN05, DSS07b, DJ06, DV09, Dur06, EHL06, ET07, Esc08, FG05, FG06, FV09, FVRR04, Fit04, FFP07]. models [FZ01, FHH07, GG07, GC08, GGJ04, GRS06, GW04, GP02b, GW05, GM07, GJ08, GL01, GS04, GS08, HHK07, HI03, HI07, HL05, Han00, HS02a, HL06, Han07b, Hay09, HPU06, HCL08, HN05, Hid03, HZ07b, HZ07c, Hid08, Hil05, HS03, HT03, HS06, HH04, HKvD07b, HPT02, Hu08, IS02, IK06, Ino08, JPR04, JJP07, JKS03, JMR09, JL08, JST06, JX05b, KKP07, KS08a, KLV04, KP08, KP01b, Kha01, KT07, KT09, Kim04, KAR07, KPS08, Kim09, Kle04, Kle09, KY09, KV00, Koo01, KT06a, KYH09, Kue01, KH09b, jLV06, Lee07, fL07a, fL07b, fL07c, Leh09, LP08c, Li02, LHZ03, LH04, Li08, Li09, Lin07, Lug06, LT02, MK06, MM07, MR04, MME08, MW08, Nak00, NS07a, NS03, Nic06, Nob00, Oku09, OP04, Ots08, PWT08, PS08, PSS03a]. models [PSS03b, PSS07a, PP04a, PQ06, PSS00, PT05, Phi00, RHSP05, Ram02, Rob01, Rot09, Sai05, SW01a, Sar03, SF01, SCF08, Seo07, SL07a, ST01a, SL03, SW01b, SW07a, SWZ08, Sin01, SV07b, SU08, SS06, Tab00, Tim00, Tod09, VV05, WZ03a, WS09, Wan01, Xia09a, XP08, Yam08, Zd03, Zaf04, Zaf07, Zaf09, dLJ06, vGLP00, vG01, vGK02]. moderate [PM07]. Modified [HLT06, HLT12, HW06, OK00, SO04]. moment [AC07, ASM08b, AL01, CCFG07, CMvS00, CCK00, CHS07, CC07, CP09, DIN03, DIN09, HIJS07, JP09b, KT09, MS09b, Phi00, Ros08, Smi07, vAJM02]. Moments [Tim00, BZ02, BZ04, GG03, HI03, HI07, HIJS07, IS08, IS02, LM02, Oku09, OT05, PQT01, SS09]. monetary [Bar07]. money [HS02b]. Monitoring [AG06]. monopolistic [DF08c, vdB07]. monotone [MM07]. monotonic [JX09]. monotonicity [McC08a]. Monte [BM00a, BBvDvO04, CNS02, CS06b, DGR06, Duf06, Dur06, GRZ07, KW06b, LG02, MN00b, VI02, Vil01]. Monte-Carlo [VI02]. monthly [RR02]. Moran [KP01a]. mortality [BvS08, Bha03b, Cos03]. most [GSCV06]. movement [ASWY01]. movements [CHP09]. moving [Hid03]. MSE [AL04, CS07a, LS09a]. multi [Bha05, BP07c, Dav02d, JKS03, KLM08, McA08]. multi-factor [JKS03]. multi-object [BP07c]. multi-stage [Bha05]. multi-state [KLM08]. multi-unit [McA08]. multi-way [Dav02d]. multicollinearity [SM02]. multifractal [CF01]. multifrequency [CFT06]. Multilateral [Bar07, Fré09]. multilevel [Han07b]. multinomial [BHH08, IvD05, MPR00, Nob00]. multiple [AHR09, Bas08, EG06, FKS00, FKS05, GP02b, HN00, KP08, KAR07, Kle07, MM08, MT03, SW01b, SWZ08, WZ03a]. multiple-equation [SWZ08]. multiple-output [FKS00, FKS05]. multiplicative [CK00, HS07]. multistep [FPW04, MSW06]. Multivariate [GJ06, MS09a, AM09, BC08a, BIKY07, Bow07, CF06a, CNS06, CHP09, DS07, DK02b, Eic07, FS07, For08, FZ01, GJS09, Haf08, HM03, LL09, PY08b, SC07, SA04, Shi07, Tse00, ZBK09].

n [CP02, YdJfL08]. Nagar [Phi00]. Nagar-type [Phi00]. narrow [CN06]. narrow-band [CN06]. national [JDGZ03]. Native [LP03]. Natural [HKvD07a, Whi06]. NBER [IV06]. near [ES01, SM02]. near-multicollinearity [SM02]. Nearly [Can08, AL04, Cho02]. Nearly-singular [Can08]. negativity [MT08a]. neglected [DGR03]. Neglecting [Hil05]. Nelson [CDN08, DLY08, Kim08, Nel08, OZC08]. neonatal [BvS08]. Nested [Ant01, BSJ01, CK02, CK07, CM01, CW07b, Dub07, LG02, Lug06, RHSP05, RS02]. network [MMS08a, MMS^{+}08b, SL04]. networks [BDF09, GG00, HKvD07b]. Neural [MMS^{+}08b, GG00, HKvD07b, MMS08a, SL04]. neutral [BS03, Bon03, CGM00]. No [ABD07, Moe08, Seo06, Wes01]. No-arbitrage [ABD07, Moe08]. Nobile [MR00]. noise [ABD07, Art04, JO08, LM08b, PY09a]. noises [DF09]. noisy [BR09]. nominal [Kon05]. nominal-to-real [Kon05]. Non [Bre07, KPS07b, MS06, WW08, AL04, AB00, BRT03, BR08, CT07a, CT08, CK02, CK07, DP08a, DV09, FFP07, GH09, GS06, HL05, HD08, IS08, Kar02, Lee08, LHMC09, Lug03, Lug06, MGM08, MT08a, MS04, RS02, vGLP00]. non-admissible [MS04]. Non-causality [MS06]. non-experimental [HL05]. non-Gaussian [AL04, GS06]. non-identifying [GH09]. non-ignorable [HD08]. non-linear [FFP07, Lug06, vGLP00]. non-local [DP08a]. non-negativity [MT08a]. non-nested [CK02, CK07, Lug06, RS02]. non-normal [BRT03]. non-normality [IS08]. Non-parametric [Bre07, KPS07b, WW08, LHMC09, Lug03, MGM08]. non-participants [BR08]. non-random [Lee08]. non-regular [DV09]. non-response [HD08]. non-stationary [CT07a, CT08, Kar02]. non-stochastic [AB00]. nondurable [RR02]. nonexperimental [ST05b]. Noninformative [NS03]. Nonlinear [BK08, Cha02, PPC04, SP03, dJ01, dJ02a, CP07b, CSW00, CS02, DFG04, Dav02a, Dav02b, DT02, DG02, GG07, GW04, GL00b, HT03, HS06, Hu08, KSS03, Li02, Par02, RSU96, RU05, Sai05, SW01a, ZRT01, vDFP02]. Nonlinearity [Mar08, BK07, DGR03, GKvD07, PT07a]. nonnested [DM02]. Nonparametric [ASL00, ASD03, Bre02, BT03a, CW08, CFS02, CCZ05, CHS07, CDSV08, DS07, DH00, DE07, FG05, Frö07, GGV08, HCL08, KPST07, LMPX09, LM08b, MFY07, OFRP05, RL04, SL04, SK00, YH06, ZW09, BM00a, BL02b, BDGT00, CK00, Cop03, Das04, Das05, DEG01, EV08, EE00, FW08b, FS05, For08, FNS04, GG07, GL00b, GL01, HHP09, HK03, HL09, HN00, JX05a, Lav01, LP08a, LHZ03, LMR09, LS09b, MFY08, Sic05, SU08, Xu09]. nonregular [Sar03]. Nonresponse [Nic06, KMR07]. nonseparable [CIN07, Jun09]. nonsimilar [AMS08]. nonsmooth [CP09]. nonstandard [Duf06]. Nonstationarity [ADG07, Hay09, Mar08, SL07b]. Nonstationarity-extended [ADG07]. Nonstationary [CP07b, HP04, Par02, PJH07, AV09, Bai04, BP07a, CLP09, Cho02, DT00, Jin09, Moo04, NS07b, PP04a, Rob05, SO04]. normal [BRT03, CW07b, Hor05, SA04]. normality [BM05, CN06, IS08, dJ02b]. normalization [Nob00, WZ03a]. normalizing [SvG09]. normals [MS09a]. normative [Man00]. Norwegian [AHV05]. note [GL02, dJ02b]. Notes [Tau01]. November [Ano00j, Ano02p, Ano03y, Ano04s, Ano06s, Ano09p]. NQ [FS08a]. nuisance [Duf06]. null [Jan05, KP09b, Seo06, SK06b]. number [AC03a, AB01, CP04b, DGR08, Dir01, Ros08, SH07]. numbers [Hil06, JP09a]. numerical [HW08a]. numerically [fL00].

object [BP07c]. objectives [Zel07c]. observations [Dir01, Pot03, SW01a]. observed [BN06a, CL00, GM06b, JO08]. obtaining [Phi00]. occasion [Dir01]. OCS [LPV00]. October [Ano00o, Ano04r, Ano05-27, Ano06w, Ano07r]. offerings [Reb04]. offs [BP08]. oil [Ham03]. older [Cos03]. OLS [LS09a]. omission [CES08]. omnibus [ALS04]. one [AD07, BBP06, Cop04, Dur06, ES01, fL00, YH00]. one- [Dur06]. one-factor [fL00]. one-sided [AD07]. one-way [BBP06, YH00]. Online [San07]. Open [Ano01j, EL03, Hsi01, Kri01]. Opening [Gra06]. operations [WW08]. operators [Tur00]. opinion [Lüt07]. Oprah [Bas03b]. Optimal [ET04, For05, Kue01, RV02, SV07b, WZ03b, vG01, Jan05, Oku09, PT07a, SK06b, vdB07]. optimality [Cla07]. Optimally [DT09]. Option [CHJ06, ASWY01, ASBS01, ASD03, BRY08, Bat00, Bat03, BGW00, rMCGJL08, GLR03]. options [BCC00, BDGT00, Jon03]. oracle [LP08b]. order [ALM06, BU07b, CC07, DP02, GP02a, HK08a, Hid02, Maa01, PQT01, RSU96, RU05, RH03, XP02b]. ordered [Cop07, HZ07a, LS07, MME08, MT08b]. oriented [KPS07a, KT06b]. origin [KS07]. Origins [And05]. Ornstein [GS06]. other [HIM05, OZC08, RM07, Zaf09]. out-of-sample [CW06, HLZ07, McC00]. outcome [Man07]. outcomes [AHV05, CL09, Chi07, CJ07, HPT08, MT03, SS04]. outcry [EL03]. outlier [Bas03b]. outliers [GKvD07, HMS05]. output [FKS00, FKS05, Kim08, Pes07]. outputs [AD05, GO07, LP03]. overcome [ST05b]. overidentifying [MS09b]. overview [BU05, GMA08, Gol07, McA07]. overweight [LHMC09]. owner [EL03].

P500 [BKPT09]. Pages [Ano00s, Ano00r, Ano00p, Ano00n, Ano00q, Ano00l, Ano00m, Ano00k, Ano00j, Ano00o, Ano00i, Ano01o, Ano01m, Ano01p, Ano01r, Ano01q, Ano01n, Ano01l, Ano01k, Ano02s, Ano02q, Ano02r, Ano02p, Ano02m, Ano02o, Ano02n, Ano02l, Ano03s, Ano03t, Ano03x, Ano03w, Ano03v, Ano03y, Ano03u, Ano03r, Ano04s, Ano04u, Ano04t, Ano04v, Ano04r, Ano04q, Ano05z, Ano05-28, Ano05x, Ano05v, Ano05y, Ano05-27, Ano05w, Ano06x, Ano06u, Ano06p, Ano06v, Ano06r, Ano06t, Ano06s, Ano06w, Ano06q, Ano07p, Ano07q, Ano07o, Ano07n, Ano07r, Ano08r, Ano08q, Ano08s, Ano08p, Ano08o, Ano09n, Ano09t, Ano09u, Ano09s, Ano09o, Ano09r, Ano09q, Ano09p]. pair [Gre06, Pes07]. pair-wise [Pes07]. Panel [BKN09, CKP01, KKP07, PW08, Sic05, SBES05, ALS01, AL01, Ant01, AC03b, Bai04, BSK03, BSJK07, Bha08, Biø04, BK06, CC04, Car07, CMvS00, CMvS01, CC01, CJ07, Das03a, Das04, DS06, FV09, GV07, Gew03, Gre05, HHK07, HOS05, Han07a, Han07b, Hay09, HCL08, HH04, HL04, HPT02, Ink00, Ino08, IL09, Kru08, Kru14, LLC02, MPP07, MW08, Nic06, Oku09, PSS03a, PSS03b, PSS07a, PS07a, PS09b, SYR09, SK06a, WW02, Wan01, Yam08, YdJfL08]. panel-sample-selection [GV07]. panels [Cha02, Cha04b, IPS03, Jin09, McK04, MP04, PY08a]. Paper [Gra07]. papers [Zel01]. paradox [Kle04]. Parameter [TC09, Che00b, Cho00, DF02, Han06, Hil05, Lan00, PS07b]. Parameterization [FGHM08]. parameters [CI08, Duf06, FV09, GH09, GL01, GLS02, Hu06, Kle07, Mac04, MPR00, OFRP05, Ros08]. Parametric [FS05, KY09, LHMC09, ALM06, BP07a, Bre07, CG07, CHS07, CHMR08, DH05, DP08b, GL00b, HOS05, HL09, KPS07b, jLV06, Lug03, MGM08, Sic05, SW07a, Wha00, WW08]. Parental [WKV03]. paribus [BS00]. parking [FTC03]. part [SW01b]. Partial [KT07, McA08, Mol08, Cho00, JP09b, Kri08, PSZ08]. partially [GH09, JX05b, Ros08]. participants [BR08]. participation [SSVY09]. pass [PW08]. passenger [MMS^{+}08b]. Past [Bol01, Gra01, Gra05, HIM05]. PAT [EL05]. Patent [BMS07, HL07, Mac04, SV07a]. Patenting [GO07]. patents [MCM07]. path [Eic07, Oul08]. path-dependence [Oul08]. paths [AHM^{+}03b]. pathways [ACM03]. Patient [Mal08, CS04]. Paul [Bas03b]. pay [Cap04]. Pearson [BSW04]. peer [BDF09, Kra06]. Performance [AMS07, AT06, PN05, Rot04]. periodic [BT01]. periodogram [HMV06, Phi07, SP03]. periods [HPT02]. permanent [OZC08]. permanent-transitory [OZC08]. permutation [Lug06, MGM08]. Perron [MOC05]. Perron-type [MOC05]. Persistence [AT06, BT04b, CT08, HLT06, HLT12, Kim00, KA02, Kim08]. persistency [BM06]. persistent [Gos02, HP08]. personal [Gra05]. perspective [DP08a, MN00a, SW08b, Sim00]. perturbed [SP03]. Philosophy [Zel07c]. Pierce [HLNS06]. piped [JR03]. Pitfalls [KW06b]. pivotal [Can07]. plan [CGS08]. planning [Bha03b]. plus [Art04]. Point [Jan05, Bow07, CC04, LS08b, SK06b, Vil06]. points [CCZ05]. Poland [OP04]. polarization [And04]. policy [BDW07, CR08, DG08, FW08a, HL05, Hec03, JLWY08, KPS07a, SW08a]. policy-oriented [KPS07a]. polluting [FGNW05]. pollution [AD05]. polynomial [CH03a, SU08]. popular [OZC08]. pork [Kar02]. Portfolio [Stu03, MS09a, OS06]. portfolios [Lam01, PS00, Sen04]. Portmanteau [EL09]. positive [Bon03]. possibly [AC07, AM08, CP09, DV09, FL03]. Post [Bat00, AG09]. Post-’87 [Bat00]. post-consistent [AG09]. postal [HW08b]. posterior [HKvD07b]. posteriors [KP02]. potential [CL09, Chi07, GS08]. potentially [BP07a]. poverty [Bie02, DF07, Zhe01, vdK08]. power [AL04, CM05, DM06, DP08a, For05, JX09, KP09a, MPR02, MPP07, Mül05, PWT08, Pes04, ZZW09]. power-GARCH [MPR02]. power-transformed [PWT08]. powerful [HLT07, HLT08]. Practical [Deh05a]. practice [CG08, FGNW05, GC08, IL08b]. Pradesh [Bha03b]. pre [RV02, Rei07, WZ03b]. pre-test [RV02, Rei07]. pre-tests [WZ03b]. predetermined [AC03b]. predictability [LW07, MR02]. Predictable [WW09]. Predicting [GSCV06, HIM05, BS06]. prediction [CW06, DHL06, FPW04, vG01]. predictions [CC04, Rot04]. Predictive [CSO01, CS06c, CDS09, Gra06, SEGG06, CM05, CW07b, CS02]. predictors [BN08, DGR08, LY06]. preference [ABW07, GLS02]. preferences [BS08, FW08b, Slo08, vSDG02]. premia [Kle09]. premium [BL02a]. prenatal [EL05]. presence [AD05, AB00, BSEY01, BK07, BT01, BT03b, CT08, CK00, CH03a, CM05, CS06a, Dav06b, Gre06, KL08, Mac04, MT08a, PT07b, Phi02, dJ02a]. present [Phi01]. Presentation [Gra06]. preserving [WZ03a]. pretesting [DM04]. price [ASWY01, CCL03, CGGT03, FFP07, HH09, HM08, Li05, Lie01, SH07, YH06, ZBK09]. prices [BGW00, rMCGJL08, DF09, EL03, HN06, HW08b, JKS03, JO08, Li05, vdB07]. Pricing [BCC00, GG00, ASD03, BRY08, Bat03, Ber01, Che03, GLR03, KS02, Sin01]. primal [FGHM08, KW06a]. prime [HSWW08]. prime-age [HSWW08]. principal [DGR08, DS06]. prior [CP02, CHP09, Des06]. Priors [KP02, FLS01, HKvD07a, NS03]. private [CR08, DSK07, HKL05, LPV00]. probabilities [ASWY01, ABR07, SA04]. probability [EHL06, GH06b, Kar02, Mol08, OK00, Woo07, Zel07a, Zel07b]. probit [BHH08, CZ07, FV09, HZ07a, IvD05, Ink00, MPR00, MT08b, Nob00]. problem [CHMR08, CF07, Deh05b, FRW05, Lan00, Oul08, SP01, SM02]. problems [Man00, MZ04, PS09a, Woo07]. procedure [Biø04, CDSV08, Kim09, fL00]. procedures [AL01, AG09, AHR09, Dav06a]. process [Bow07, Cor00, DP08b, GJ06, GJS09, HW06, LP08d, LS00, LST03]. processes [ASY06, ALM06, BK08, BNS06, BH02, CV03, CSW00, CS05, Dav02a, Dav02b, DS05, Dav06a, DGR06, Dis08, DG02, DT00, EZ01, FMZ08, GG04, Gos02, GW02, GH06a, GS06, Haf08, HP08, Hid05, HK06, Hid09, HHS07, Laz05, LM02, MC06, MPR02, MP08, Par06a, PY09b, Pos09, Rob05, RT04, Shi07, SW07a, SP03, TC09, TC00, XP02b, ZW09]. processing [WW08, Zel02, Zel07d]. Product [KAR07, PN05, Zel07a, Zel07b]. Production [Mos01, BGKT04, FKS00, FKS05, GO05, KT06b, PN05, RR02, SW07a]. productive [KPS07b]. Productivity [FS08a, AP02, AD05, DK05, DSK07, SBES05]. profile [MRS07]. Program [Deh05b, CvdK08]. programming [CDSV08, Tab00]. programs [AHV05, CR05, HIM05]. PROGRESA [DS08b]. Progress [FvD07, DF08c]. project [CH09, CR05]. projection [DT07]. projection-based [DT07]. projections [KS02]. propensity [Deh05a, SSVY09]. Properties [DG02, PT07a, ZZW09, ACMP08, BU07a, CC07, DP08a, DGR06, DS06, FMZ08, HP08, Han07a, HK08a, KH09b, LLC02, OK00, OS06, PT05]. property [GO07, LP08b, McA07, dMF07]. proportional [HL04]. prospect [ABW07, Die01, Nel08]. prospects [Koe00]. proxy [DG07]. Pseudo [KS08a, CK02, CK07, Ino08, McK04]. Pseudo-likelihood [KS08a]. pseudo-panel [Ino08]. pseudo-panels [McK04]. pseudo-true [CK02, CK07]. PSID [GK00]. Psychology [Coo08]. public [MC08, Reb04, SW08a]. publications [FF07]. publishing [Dir01]. Pure [EM06b]. Purebred [ADGG03]. puzzles [BZ06].

QML [CI08]. QMLE [ACMP08]. quadratic [ASM08a, KL08, MJ05, Plo04]. quadrature [fL00]. qualitative [Lew00]. Quality [AG08, BS04]. Quantile [HKP02, MK06, Xia09b, CH06, CH08, CHJ09, DS07, Gal09, GJ08, HL09, Jun08, Jun09, Kha01, Koe00, LY06, Lee07, MFY08, Ots08, WW08, YS07]. quantile-regression [HL09]. Quantiles [DH09, BCY00, BP07b, Kom05, Qu08]. quantilogram [LW07]. quantity [Dav06b]. Quasi [Kom05, YdJfL08, Gir00, Lin07]. quasi-differencing [Gir00]. Quasi-maximum [Kom05, YdJfL08, Lin07].

R [CDN08, Muk07]. R-estimators [Muk07]. radial [BBvDvO04]. radial-based [BBvDvO04]. random [Ant01, BSJK07, BH07, DGR03, Dis08, EHL06, HLZ07, HKP02, HN00, JP09a, JMR09, KLV04, fL00, Lee08, Lug03, MW08, RHSP05, Rot02, SS01]. random-component [fL00]. Randomization [HKL05]. Randomized [Lee08, AAD05, Mal08]. randomly [ASM08b]. range [CG08, CP07a, HHS07, MV02, MvD07, Zd03]. range-based [CP07a]. Rank [Abr00, Che00b, Has01, And02, And06, ABG05, AV09, BH02, For08, HKvD07b, Kha01, KT07, KP06, LS00, LST03, NS07b, RY02, Shi01]. ranking [HL06, Hor05, LP04, SS09]. rate [GZ06, HSWW08, Jin09, LT06b, Wri08]. Rates [CK03, BGW00, CFK06, Das02, FHLR04, GT09, Hon01, HLZ07, JKS03, OP04, PW08, dJ02b]. rating [KLM08]. ratio [CHS07, SP01, Tay05]. rational [CDSV08, JL08, Leh09, MN00b, Pol00]. re [BR05b]. re-employment [BR05b]. reaction [LW01]. real [CS01, Gre06, Kon05]. real-time [CS01]. realignment [Yu07]. realised [BNS06]. realization [Dom01]. Realized [BRY08, CP07a, BKPT09, CDS09, DHL06, MvD07, PY09b]. recession [IV06]. recipients [AAD05, CR05, CH09, HSWW08]. Reconciling [ABW07]. Reconsidering [Cor00, Gre05]. recovery [MhL02]. recurrent [Xu09]. Recursive [ST01b, MK06, SS00, SS01]. reduce [JR03]. Reduced [And02, And06, HKvD07b, KP06, Li08]. reduced-form [Li08]. reducing [AD05]. reduction [DS04]. redundancy [PS09a, PS09b]. Referee [Ano07s]. Reference [Sar03]. refinements [CI08]. Regime [BGW00, Pel06, DI01, HN06, HS02a, Jin09, KPS08, LPS05, MM08, MP08]. regime-switching [KPS08, MP08]. regimes [DS05]. region [vGK02]. regional [BEP08]. regions [GH09]. Regression [Biø04, IL08b, LC08, Ram02, VKG09, Abr00, ASBS01, And02, And06, AMS07, AM08, AMS08, AHR09, BSJ01, BSK03, BMM09, BR08, BH07, BCY00, BP07b, BT01, CMvS00, CK00, CK03, CG07, CH06, CH08, CHJ09, CP07b, Coo08, DGR03, DM02, EV08, Fit04, For05, FRW05, GL00b, GL01, HM03, Hid03, Hid08, HKP02, HH04, HKvD07a, HKvD07b, HL09, Hu06, IL08a, KP08, KP01b, KT09, KPS08, KZ03, Kle04, Koe00, KP04b, KH09b, Lal08, Laz05, Lee07, LM08a, LS09b, Lug06, MK06, MFY07, McC08b, MMS^{+}08b, Ots08, PSZ08, Phi07, PY08b, Qu08, RL04, RV02, Rei07, SW01a, SL07a, SK00, ST01b, SP03, TC00, Tur00, WZ03b, Wha00, XP02a, Xia09b, YS07, ZBK09, vGK02]. regression-based [BT01, ST01b]. regression-discontinuity [BR08, Coo08]. regressions [BZ06, CPS06, CHO08, DK02a, DK02b, DT07, EM07, FL03, GK07, Lav01, SO04, XP02b, XL04]. regressive [fL07a, fL07c]. regressor [Hu06]. regressors [BP07b, Das05, IS08, Rot09, SYR09, SO04]. regret [Man07, Sto09]. regular [Des00, DV09, McC07a]. regularity [Bar02]. rehabilitation [AHV05]. Rejoinder [Hau01b, ST05a]. related [GKLT03, GKLT05, MZ04]. relation [DS04]. relations [Joh06, dJ01, dJ02a]. relationship [For08, OZC08]. relationships [Gra03, HMV06]. Relative [HW08b, Bas08, CL00, KP09a, Zhe01]. relevance [Maa01]. relevant [HS00]. Reliable [Dav09]. Renewal [LPS05]. rental [CR08]. repeated [GH06b, HM08, HSWW08, Mor04, VV05]. repetitions [AB01]. replacement [CR08]. Reply [MR00, Deh05a]. report [Ano07s, CGS08]. representation [Cop04]. representing [Sen04]. reproducible [Vin01]. Reproducing [ADGG03]. Resampling [DP06, BCY00]. Rescaled [GKLT03, GKLT05]. research [AT07, CDN08, LP04]. reservation [Li05]. Reserve [WW08, EL03]. Residual [BLS06, HMV06, IS08]. residuals [CP09, XP02a]. Response [AHM^{+}03a, Hor01a, Aba03, CK03, CDS02, Chi07, CJ08, Cop01, Cop07, FFP07, HD08, LS08b, Lew00, Man00, PW08, Ram02, Rot09, vDP08]. responses [AHV05]. Restricted [PFC08, PQ06]. restriction [AC07, DIN09, Sak07]. restrictions [ASD03, ABD07, CMvS00, CCK00, CC07, Che09, Das03a, DIN03, GH09, GSN08, HS02b, JP09b, McA08, McC08a, OFRP05, Smi07]. results [But00, CS06b, DT07, PS09a, Sai05]. retirement [vdKW08]. Retrospect [Die01, Nel08]. retrospection [Bat03]. return [BP08, BZ06, CCZ08, GSCV06, LP08b, Mor04]. return-volatility [BZ06]. Returns [CL00, Arc04, BL02a, BNS06, BPT01, BKPT09, DF08c, GO05, JLS05, MR02, PQT01, STW01, WW09]. revealed [ABW07]. reverse [Che03]. reversibility [CCK00, DFG04, McC07a, RM07]. revisited [MS00, PFC08, Sch04, SV07a, SM02]. Revolution [GO07]. rich [Moe08]. Riesz [AHT07]. rights [GO07]. Risk [BLT08, ASL00, AG08, BS03, BP08, BL02a, Bon03, CDPX07, CGM00, DF09, GMA08, GT06, HLW09, Kle09, KYH09, Mos01, Rei07, Yu07]. risk-neutral [Bon03]. risk-return [BP08]. risks [NS03]. risky [PS00]. rival [Rob05]. Robust [CHO08, GTU05, Li02, LH04, McC00, OT05, RT01, BBvDvO04, BS04, CFS02, CH08, CM07b, Han07a, HLT07, HLT08, HLNS06, Jun08, Kle04, Kle07, LMPX09, Mül07, RM07, WW02, dJAS07]. Robustifying [Hen06, Im00]. Robustness [PS09b]. role [BW00, HKL05, Leh09]. rolling [ST01b]. Root [GV07, HR07, AB00, BGG08, BT04a, Cha02, Cha04b, CS06b, Dis08, FMZ08, Gal09, Gos08, HMS05, Has01, KSS03, KLN02, KP09b, Kru08, Kru14, LLC02, MP04, MPP07, PPP06, PR03, Phi02, PM07, Phi07, RT04, RT07, SK06a, ST01b, Tay05]. Root- [HR07, GV07]. roots [Bre02, BT03a, BT01, BT03b, CT07a, Cha04a, Cha08, EJ03, EM06a, FdV07, Gre06, IPS03, MN00a, SS00, SL03]. Rotterdam [Gra07]. rule [Gol01, Zel07a, Zel07b]. rules [AC03a]. run [ADG09, BP08, BC06, BC08b, DPR06, EM06b, HMV06, Mül07, Phi09]. running [McC08b]. rural [JR03].

S [BKPT09, SW01a]. S&P [Bat00, BPT01]. S-estimation [SW01a]. Saddlepoint [ASY06]. Sample [Nev02, PS00, ASM08a, ACMP08, BU07a, BIKY07, CHJ09, CW06, CS02, Das04, Des00, DS06, DT00, DK02b, DKBG04, DJ06, Duf06, GV07, HI03, HI07, HLZ07, Ink00, Joh02, LS09a, LLC02, McC00, McC07b, OK00, PT05, SB02, Win05]. sampled [ASM08b, GSCV06, ZW09]. samples [Bha05, Gas03, Ram02, RS00, Sto09, ZC01]. sampling [BBvDvO04, But00, Cha04a, Cha08, HKvD07b, KSC09, MC06, Nak00, RZ07, SA04]. San [Gra06]. SAT [Rot04]. satisfy [Ros08]. saving [Hou03]. savings [vdKW08]. scale [DF08c, GO05, MS09a]. scanner [FFP07]. scanner-data-based [FFP07]. Scheinkman [ASM08b]. scheme [rMCGJL08]. schemes [DS05, KLV04]. school [Mat08, Reb04]. schooling [BMM09, HHM04, JLS05]. schools [MC08]. scientific [FF07]. score [CK02, CK07, Deh05a, FS08b, Hor02, Moo04, MPS09, SSVY09]. scores [HHM04]. search [DF08b, EvdB07, GSN08, HSWW08, Koo01]. Seasonal [HS02b, BT01, BT04a, DHL06, HMS05, RT04, RT07, SS00, SL03, SO04, SL07b, ST01b, Tay05]. Seasonality [FFP07, BT03b]. seasonals [KV00]. second [BU07b, CC07, RSU96, RU05]. second-order [BU07b, CC07, RSU96, RU05]. secret [EL03]. section [Bai04, CH00, GH06b, HM08, PS07a, SYR09]. sectional [Cha02, Cha04b, HH04, KP04a, Mor04, Rob08a, SK06a, vGLP00]. sections [Gir00, VV05]. sector [WvDW03]. sectoral [GO07]. securities [GG00]. Security [LSW08, Bow07, vdKW08]. seek [AAD05]. seemingly [DK02a, FRW05, SK00, Tur00]. Selection [AT07, Mal08, MOC05, PT07b, AL01, AG09, CF06a, CHS07, DG07, GV07, Gol01, GP02b, HIJS07, Hid02, Hor05, IK06, jLV06, Lee08, Lew07, Li09, MZ04, MT03, MT08b, Nev02, PS08, PS00, Rac00, SNT06, ZBK09]. selective [BR05b]. Self [Lin07, CH09, MT03, CR05]. self-selection [MT03]. Self-Sufficiency [CR05]. Self-weighted [Lin07]. sellers [vdB07]. Semi [BL02b, ABD07, CHMR08, jLV06, LHMC09, SW07a]. semi-martingale [ABD07]. Semi-nonparametric [BL02b]. semi-parametric [CHMR08, jLV06, LHMC09, SW07a]. Semiparametric [Aba03, BMM09, BW00, BP07c, CH02, FLS07, GS04, HL04, JP09b, KT06a, LS08b, Lew00, MR01, PSS03a, PSS03b, Par03, PSS07a, Rot09, SS06, Tab00, AC07, Art04, CMvS01, CC01, CF06b, CF06a, CP09, Dav02a, Dav02b, GV07, HY05, JLWY08, KP01b, KP04b, LHZ03, Lin02, PS08, RH03, ST01a, Shi07, SO04, Sic05, Yan06]. Semiparametric-efficient [PSS03a, PSS03b]. semiparametric/nonparametric [LHZ03]. sensitivity [BM00b]. separability [FW08b]. September [Ano00i, Ano01r, Ano02l, Ano04q, Ano05w, Ano06q, Ano08o]. sequences [Deo00a]. Sequential [Pal09, AHR09, BP06, Bre07, Das03a]. serial [BSJK07, DFH06, EL09, Han07b, Yam08]. serially [Cai07, KH09b]. series [AT05, AC03a, ALS04, AHR09, BN01, BN08, BU07b, BS06, Cai07, CLP09, CT07a, CP03, CC01, CCZ05, CF06b, CG07, Dah09, Dav02a, Dav02b, DT02, DV05, DEG01, DFH06, DPR06, Eic07, EM07, FPW04, Gir00, GTP06, Gre06, GS08, HP08, HTV07, Hid03, HT09, Hsi09, IS06, Kim00, KA02, LY06, LHZ03, LT06a, LP08d, LS09b, LT06b, MSW06, MV02, MS06, Mül05, Phi01, PP04b, Rob08a, SL03, SL07b, Whi06, XP02b, dJ02b]. service [KM09]. services [HW08b, HSWW08]. set [CS01, vdB07]. sets [EM07, HKvD07b, JM07, PY08b, Ros08]. setting [BL08, PS09b]. settings [Sic05]. several [Cop04]. severely [Cop03]. shadow [AP02]. shape [ASD03, BMM09, HKvD07b, OFRP05]. share [Dub07]. sharpen [Mul06, Sim00]. shift [LST03]. shock [Ham03]. shocks [GM06a]. Short [DPR06, Tho00, ADG09, BC06, GT09, HPT02, Kri01]. short- [BC06]. short-term [GT09]. shortfall [CW08]. Should [WW09]. shrinkage [DGR08]. sided [AD07, AMS08, DT00]. Siegel [DLY08]. Sign [DV05, SS01]. signal [Art04]. significance [SV07a]. similar [For05]. Simple [BCY00, Das04, BP07a, BZ06, CS06b, EE00, HLT07, HLT08, HT03, LS07, Shi01]. simplified [ST01a]. simulated [CS07b, GG03, Nie00]. Simulation [DK02b, DKBG04, Kra06, Li09, CW07a, DJ06, GLS02, MN00b, MT03].

Simulation-based [DKBG04, Kra06, DJ06, MT03]. simultaneous [And06, CP02, GL02, GGV08, Gol01, KP04a, Phi00, RT06, SU08]. since [Kim08, Lan00]. single [EM07, GP02b, HN05]. singular [Bie07, Can08, KP06]. SIPP [LSW08]. SIPP/Administrative [LSW08]. Size [Mül05, AG09, BL02a, GM06a, Mor09, SV07a]. skeptical [WW09]. skewed [FS07]. skewness [CHJ06, MS09a, RJ02]. Skill [BR05a]. Skill-biased [BR05a]. skills [CL00]. slope [KP09a, PY08a]. slopes [VF05]. Slutsky [HHP09]. Small [PT05, Des00, GM06a, Ink00, Joh02, PS07b, RS00]. small-sample [Des00]. Smith [Deh05a]. SML [Ink00]. smoking [JDGZ03]. smooth [GJ06, KT06a, MFY08, MM08, VKG09]. smoothed [Hor02, SL07a]. smoother [CCZ08]. smoothing [Dur06, LMNT01]. Smoothly [GK07]. SNP [CG02]. Social [LSW08, vdKW08, BDF09, BD07, LM08a, Mor04, dP09]. Socio [ACM03]. Socio-economic [ACM03]. socioeconomic [AHM^{+}03b]. solution [Hu08, Oul08]. solutions [MZ04]. Some [BDW07, Dir01, Flo03, Gra03, Hoo03, Phi01, Pot03, Rot02, Zel07d, ABG05, BBvDvO04, CMP09, Gra05, KP04b, SP06, SV07b]. sorting [Arc04]. space [GG04, Vil06]. space-time [GG04]. spaces [Sch04]. spacing [BvS08]. Sparse [LP08b, HW08a]. Spatial [CM07b, BSK03, BSJK07, BU07a, CC01, KP07, KS07, LP07, fL07a, fL07c, PSS07b, Rob08a, SC07, YdJfL08]. spatially [BKP07, KKP07, KP04a]. Special [ABCD07, Ano07m, HL05, IL08a, SW08a, CDN08]. specific [CKL06, Lüt07, Sic05]. specific-to-general [Lüt07]. Specification [CCL03, Del08, EV08, GC08, Hid08, CGM00, CS05, DS08a, EE00, Esc08, FG05, GM07, Hau03, HL07, HK06, HLR07, LP07, LC08, LHZ03, MME08, SSVY09]. specifications [CLP03, GT06]. specified [CD06]. spectra [XL04]. Spectral [CV03, Deo00b, Ber01, Dah09, EV06]. speculative [MN00b]. spell [SW01b]. spells [GH06b]. spherical [Bek02]. spillover [Hon01, HLW09]. spillovers [CS04]. splines [DH09]. spurious [CHO08, Mar08, NS07a, TC00]. square [Pol00]. square-wave [Pol00]. squared [BU07b, BSW04, CW06, RSU96, RU05]. squares [And05, CN06, DP08a, GL00b, Han07b, Han08, HK08a, HW06, IS08, PS07b, Sch04, SL07a, Wil04, YH06]. Stability [KLV04, Sai05, DF02]. stable [Deo00a, fL00, LPS05, MPR02, ZW09, vGK02]. stage [And05, Bha05, HW06, Kha01, PY09b, PS07b, SW07a]. standard [Zel07a, Zel07b]. STAR [KSS03]. State [DG08, BGKT04, CP04b, HH09, Hsi01, HSWW08, JJP07, KLM08, Lüt01, McC07a, YH06, ZBK09]. state-price [ZBK09]. states [AAD05]. static [vGK02]. Stationarity [FZ01, LM02, MPR02, BT04b, CSW00, Kur02, MV02, Mül05]. stationarity-ergodicity [CSW00]. stationary [CT07a, CT08, CN06, Dah09, Des00, DT00, EJ03, Hid05, HK06, HT09, Kar02, Kru08, Kru14, McC07a, PPP06, PY09a]. statistic [HKZ06, KP01a, vGK02]. Statistical [Bas03b, Joh06, Zhe01, ZC01, CFP09, FGS03, FSZ04, HM06, SV07b].

Statistics [Coo08, Bas03a, BHPW04, Das03b, DP02, Kle07, Lin02, Per00, SV07a, XP02b]. status [AHM^{+}03b, ACM03, Pot03]. step [AB01, KP01b, Kim09, Win05]. stepwise [Biø04]. Stochastic [AP02, GGJ04, HML02, OCSN07, Art04, AM09, AB00, BRT03, Bai04, BKN09, BV04, Bie07, BZ02, BZ04, BDGT00, BT03b, CMP09, rMCGJL08, CNS02, CNS06, CS07b, DHL06, DP08b, DGR07, DV09, Dur06, FS07, FMZ08, GSN08, GM07, GJS09, Gre05, GS04, GS06, GO05, HKZ06, Hou03, JPR04, Jon03, KT06b, KPST07, PZ05, Rob01, Tod09, WS09, Yu05]. stock [BU05, BM06, BGKT04, CGGT03, HS02b, Lie01, Liu00, MR02, PQT01, STW01]. stocks [ACMP08, SCF08]. storage [MN00b]. strategies [AT06]. strategy [Bas08, KM09]. strength [MN00a]. strike [EL05]. Strong [AC03a, Can07]. strongly [BK08, Hid02]. Structural [Car02, GG03, GM06b, Han03, PZ05, PSS00, Pos09, AG08, Ano05u, BU05, BH07, BIKY07, Can07, CCZ05, CH06, CHO08, CM05, DH00, DP08a, DP02, DF09, FV09, Fit04, GTU05, GGV08, GKvD07, Han00, Hau03, HW06, Jun09, KS08a, KP08, KP09a, Laz05, Li09, LST03, MK06, MN00a, MJ05, PR03, PQ06, PT05, Qu08, dLJ06, vSDG02]. structure [ADGG03, AM09, BS03, CFK06, DL06, DMN05, For05, GT06, HMSA03, HTV06, LaF08, WZ03b]. structures [Dav02d]. student [Mat08, vdK08, Des06]. Student- [Des06]. Studies [HP01, Ehr04]. study [BM00a, CR08, VI02]. Studying [CHP09]. subject [ABD07]. Subsampling [BHPW04, Cho05, GW05, AG09, HS06]. subsidy [CH09]. subspace [BW02, CP04a, DS04]. substitution [HW08b, fL07c]. sufficiency [CH09, CR05]. sufficient [Bar02]. summer [Mat08]. super [Hil06]. Superconsistent [DP02]. Superlative [Hil06]. supply [CES08, Hou03, JS08, Reb04, vSDG02]. support [Bas03a, Bas03b]. surprise [Das02]. survey [Bha07, EvdB07, JDGZ03]. surveys [KMR07]. swap [JKS03, JO08]. swaption [JKS03]. Sweden [EL03]. switching [BGW00, Car02, Des06, DI01, FZ01, HN06, Kim04, KPS08, Kim08, Kim09, LPS05, MP08, Pel06, SWZ08, SNT06, Tim00]. symmetric [DFH06]. Symmetry [BL08, BN01, Che00a, DE07, HHP09]. Symmetry-based [BL08]. Synchronization [HP06, dP09]. synthesis [BB02, Gol07]. System [vS04, KT05, KW06a, MMS08a]. systematic [SM02]. systems [AV09, BS08, BW02, Biø04, CM07a, CLP03, HHP09, Hen06, HPT08, HN05, KP04a, LP04, MT08a, NS07b, Par06b, PN05, PP04b, RY02, RI05, Slo08].

T [Han07a, YdJfL08]. Table [Ano04w]. tailed [CDPX07]. Tails [ST02, JPR04]. talented [Reb04]. tamed [DF08a]. TAR [Gos08]. TAR-GARCH [Gos08]. target [LT06b]. targeted [BN08]. Tastes [Bar02]. teachers [Reb04]. technical [BR05a, BMS07, DF08c, Flo03, KS08b, KT05, KT06b, KW06a, WS09]. techniques [JST06, KP04b]. technology [AP02, Bar02, FGNW05]. tell [APW06]. Temporal [Haf08, MR04, HOS05]. temporally [CM07a]. tenant [EL03]. tennis [KM09]. Term [GT06, ADGG03, BCC00, BS03, CH09, CFK06, DL06, DMN05, GT09, HTV06, RT06]. term-structure [HTV06]. terminology [Lüt07]. test [ALS04, AV09, BN01, BBP06, BSW04, CK02, CG07, CK07, CHO08, CW06, CS02, CS06b, DH05, DEG01, EL09, FW08b, FMZ08, GH09, HHM04, HLT07, HLT08, Hid05, HZ07b, HZ07c, Hon01, HLNS06, HLR07, Im00, JX05a, KP01a, Kem01, Lav01, LT06a, LMR09, Lin02, MS00, MGM08, McC08b, MPS09, PW08, RM07, RV02, Rei07, SYR09, SSVY09, Shi01, SL04, SS01, SK06b, SW07b, Tse00, XP02a, Yam08, dJAS07, vGK02]. testable [ABD07, McA08]. testimony [FGS03]. Testing [AD07, ABG05, AS07, BC08a, BSK03, BSJK07, BK07, BM05, BC06, BT03b, CT07a, CT08, Cha04a, Cha08, CCK00, CSW00, DGR03, DP08b, Dis08, DMN05, ET07, EJ03, GL01, HHP09, Han00, HS02a, HL09, HKZ06, HP01, IPS03, JO08, JX05b, KSS03, Kon05, KV00, KSC09, Kur02, Laz05, LS00, MP04, PY08a, PY08b, Qu08, VF05, dMF07, Ano01f, BL02b, BLS06, Dav06a, Del08, Deo00a, DSS07b, DJ06, EM06a, EE00, Gal09, GGV08, GC08, HMS05, HY05, HK08b, HCL08, Hid08, LP08a, Li08, MME08, PPP06, Pes07, Rob08a, Rob08b, SF01, Seo06, ZC01]. Tests [BSEY01, BT04b, CM01, JX09, Kle09, Mor09, ASBS01, AMS07, AMS08, AB00, BM00b, BGG08, BIKY07, Bre02, BT03a, BT01, BT04a, Can07, CGM08, Cha02, Cha04b, CHS07, CDSV08, Cho05, CM05, CW07b, CS05, CS06a, CS06b, CS06c, DM02, Dav02c, DM06, DF02, DV05, DE07, DS08a, DP08a, Deo00b, DIN03, DK02a, DK02b, DKBG04, Duf06, DV09, EV08, EV06, Esc08, FW08a, FG05, For05, For08, GTU05, GG03, GKLT03, GKLT05, Gos08, Gre06, GS08, HLT06, HLT12, Has01, HK08a, HK06, HS00, Hor01b, Hor02, Jan05, Joh02, Jun08, JP09b, KLN02, KS08b, KP09a, KP09b, KP06, KPS07b, LLC02, LHZ03, Lug03, Lug06, LST03, McC07b, MOC05, MPP07, Mül05, PWT08, PR03, Pes04, Plo04, RS02, RT04, RT07, SS00, SL03, SK06a, ST01b, Tay05]. tests [Vil01, WZ03b, Wes01, Wha00, AHM^{+}03b]. Thailand [WKV03]. their [AT05, JDGZ03]. them [Hil06]. Theorem [Zel07a, Zel07b, Dav02a, Dav02b, PSS07b]. theorems [JP09a]. theoretic [Bar07, GLS02, LBPA02, Nev02, Smi07]. theoretically [CLP03]. Theory [AT07, ABD07, CFK06, CG08, CMP09, GC08, IL08a, MCM07, PSZ08, ABW07, ABCD07, Bas08, BS00, BDW07, Cha00, DGR03, FGNW05, KPS07a, Koo01, Maa01, McK04, Mül07, Par06a, PM07, Rot02, Zel07a, Zel07b, vS04]. there [BEP07]. thick [GZ06]. third [BEP07]. third-country [BEP07]. Thirty [Ame09]. Thirty-five [Ame09]. Three [Sim00, AB01, MN00b]. three-step [AB01]. Threshold [Car02, BGG08, DSS07b, GP02b, GW05, HS02a, LS08b, PWT08, Seo06, SL07a]. Time [GRZ07, HP08, McC07a, SS04, Whi06, ALS01, ASY06, AC03a, ABD07, ALS04, AHR09, BN01, BN08, BU07b, BNS06, BS06, BKPT09, Bow07, Cai07, CLP09, CT07a, CV03, CP03, CCK00, CC01, CCZ05, CF06b, CG07, Cor00, CSW00, CS01, Dah09, DFG04, Dav02a, Dav02b, DT02, DV05, DP08b, DEG01, DFH06, DPR06, ELX03, Eic07, EM07, FPW04, FHH07, GG04, Gir00, GTP06, Gre06, GS08, HTV07, Hid03, HT09, HPT02, IS06, Kim00, KA02, KP09b, LY06, LT06a, Li08, LS09b, LHMC09, Lon04a, LT06b, LS00, MSW06, MV02, MS06, Mül05, OFRP05, Phi01, PP04b, RM07, Rob08a, SL03, SL07b, Tod09, WW09, XP02b, XP08]. time-deformed [BNS06]. time-inhomogeneous [ELX03]. time-reversibility [RM07]. Time-series [Whi06, BU07b, LT06a, PP04b]. Time-use [SS04]. time-varying [ALS01, Cai07, Li08, XP08]. times [AHR09]. timing [GRZ07]. timing-of-events [GRZ07]. Title [vdK08]. tobacco [HZ07a]. Tobit [YS07]. Todd [Deh05a]. top [KM09]. toxic [FGS03]. tracking [Lam01]. Trade [BGKT04, BEP08, BP08]. trade-offs [BP08]. trademark [Gas03]. trading [DF09, Lie01]. training [BGKT04, HIM05]. transaction [DMN05, GW02, ZRT01]. transformation [CS06b, Kon05, LM08b]. transformations [DG02]. transformed [HT09, PWT08]. transforms [AT05, SvG09]. transition [FSZ04, Kar02, MM08]. transitions [GJ06, KLM08]. transitory [OZC08]. translog [KT05]. treatment [AHV05, Aba03, CH06, CH00, CH02, Chi07, CJ07, CJ08, Frö07, HN07, Lew07, Man00, Man07, Sto09]. Trend [MV02, MP08, Pol00, AB00, BT03b, CMP09, HLT07, HLT08, KP09b, KV00, LS00, VF05]. Trend/cycle [MP08]. Trending [Cai07, Phi01, Pol00]. Trends [HTV07, Bai04, BKN09, CH03a, FS08a, JX05a, MPP07, Par06b, PZ05, PY09a, Phi02, PS07a, RI05]. trials [AAD05, Mal08]. Triangular [Hau03]. true [CK02, CK07]. Truncated [DG01, Hor05, PSZ08]. truncation [CGM08]. Turning [SW08b, CC04]. Two [ADG09, Ano02t, KP01b, Kha01, SW01b, And05, AMS08, DT00, Dur06, HS02a, HW06, Kim09, OK00, PY09b, PS07b, SW07a, Win05]. two-factor [Dur06]. Two-part [SW01b]. two-regime [HS02a]. two-sided [AMS08, DT00]. Two-stage [Kha01, And05, HW06, PY09b, PS07b, SW07a]. Two-step [KP01b, Kim09, Win05]. type [CHO08, Das03b, HKZ06, MOC05, OP04, Phi00].

U.S. [AAD05, FS08a, Lee08, PN05, RT06, SBES05]. Uhlenbeck [GS06]. unattended [BT03b]. unbalanced [Ant01, BSJ01, Biø04, Dav02d]. unbiased [Gos02]. uncensored [DT09]. uncertainty [AB00, BDW07, Han06]. uncompleted [GH06b]. unconditional [WW08]. uncorrelated [LL09]. uncorrelatedness [HLNS06]. underlying [ASWY01, Jon03]. Understanding [Cos03]. undesirable [AD05]. unemployment [AAD05, GZ06, GH06b, Jur02, Lal08, RS00, vDFP02, vdBvdK01]. ungrouped [DLM06]. unidentified [Can07]. unification [SR02]. Unified [HY05, GKvD07]. uniform [JP09a, dJ02b]. unilateral [CCL03]. Union [Cos03]. uniqueness [vdB07]. Unit [AB00, Gal09, KLN02, KP09b, LLC02, PPP06, Phi07, BGG08, Bre02, BT03a, BT01, BT04a, BT03b, CT07a, Cha04a, Cha08, Cha02, Cha04b, CS06b, CLP03, Dis08, EJ03, EM06a, FdV07, FMZ08, Gos08, Gre06, HMS05, Has01, IPS03, KSS03, KMR07, Kru08, Kru14, MN00a, McA08, MP04, MPP07, PR03, Phi02, PM07, RT04, RT07, SS00, SL03, SK06a, ST01b, Tay05]. unit-root [BGG08, FMZ08]. University [Gra07, BGKT04]. unknown [BK07, Fit04, GK04, KP09b, Lew00, Lug03]. unobservables [Bha08]. unobserved [CL00]. unrelated [DK02a, FRW05, SK00, Tur00]. unrestricted [Kle07]. updating [Can07]. use [ABR07, CDS09, SW08a, SS04]. useful [CR08]. usefulness [BS06]. Uses [Ull02]. Using [CW06, GL00b, Sim00, AP02, BN08, BSJK07, BW02, BZ02, BZ04, Bon03, CL09, CMvS00, CMvS01, Cop01, Dav02c, DGR08, DHL06, Dom01, DP02, DS06, EG06, FFL08, Gal09, Gas03, Gew03, GH06b, HL07, HKvD07b, HM06, HKZ06, HIM05, Hu06, Hu08, IS08, IvD05, IV06, JKS03, JMR09, JX05b, KT09, KP09a, KP06, KV00, LSW08, MGM08, MT03, PY08b, SSVY09, Sin01, SNT06, Tod09, VKG09, XP02a, XL04, YH06, dJ01, vDP08]. usual [BM00b]. utilities [AD05]. utility [ABW07, HN00, Stu03]. utilization [Ano07s]. Uttar [Bha03b].

v [Bas03b]. Valid [KS08b, DGR08]. validated [CG02]. Validating [EHL06]. validation [GRS06, Rac00]. validatory [Rac00]. validity [MPS09, PS09b]. valuation [CHJ06, CGM00]. value [CDPX07, CGS08, CLP03, HKL05, HL07, KP06, KYH09, SV07a]. value-at-risk [CDPX07]. values [Che07, CF07, HS00, Hor02, WZ03b]. Valuing [Dub07]. VAR [CC04, DJ06, GSN08, HPU06, LS09a, LS00, LST03, Moe08, Par06b, Sch05, VI02, CW08]. variable [Aba03, BV04, BL08, CIN07, CH08, CHMR08, Cop04, Gol01, GO05, HT03, HKvD07b, HL09, Jun09, KZ03, Lin02, McC08b, PPC04, RHSP05, Ram02, Sak07, SL03, SK06a, WW02]. variables [AC07, AM08, AC03b, CDXW06, CL09, CP02, Cho02, Cop04, CSO01, Das05, DIN09, FL03, GLR03, HHK07, HKvD07a, Hu08, JST06, Kim04, Kim09, Kue01, KPS07b, Lew00, Li02, MW08, OK00, PW08, Par03, PSS00]. Variance [EM06b, Tay05, XL04, ADG09, CP04a, CP07a, Esc08, GKLT03, GKLT05, Han07a, Has01, JO08, KLN02, LS09a, MS09a, Mül07, WZ03b, Win05]. variances [BNS06, XP08]. variants [KP04b, SP06]. variation [ASM08a, BGKT04, HOS05, HM08, KL08, Phi09]. variations [BKPT09]. VARs [GRS06]. vary [AHV05]. varying [ALS01, Cai07, Cha04a, Cha08, Li08, OFRP05, XP08]. vec [Tur00]. vector [BLS06, Cho05, ET07, HS02a, Han03, HW06, NS03, PSS00, Sch04, Seo06, Seo07, Shi01, Tur00]. vector-autoregressive [NS03]. vectors [Joh02]. versatile [RM07]. version [dJAS07]. versus [FMRT09, MV02, vGK02]. veterans [Cos03]. via [Cho00, MFY07, PPP06, Pol00]. Vietnam [WvDW03]. View [Gol02, Mar04]. viewing [GL00a]. vocational [AHV05]. volatilities [ASBS01, BPT01, EM06b]. Volatility [BZ06, CFT06, ADGG03, ABD07, Art04, AM09, BRT03, BRY08, BU05, BM06, BPT01, BW00, BZ02, BZ04, BDGT00, CF01, CG08, CT07a, CT08, CCZ05, CCZ08, rMCGJL08, CNS02, CNS06, CN06, CKL06, CDS09, DHL06, DP08b, DV09, Dur06, EG06, FNS04, GGJ04, GSCV06, GKLT03, GKLT05, GJS09, GW02, GM06b, GS06, HL06, Hon01, JPR04, Jon03, Lie01, Liu00, MvD07, MCM07, MR04, MMS^{+}08b, OCSN07, PY09b, PY08b, Rob01, SM02, Tod09, Yan06, Yu05, Zd03, Zaf07, Zaf09]. Volume [Ano07t, Dir01, SEGG06, Ano01e, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano03c, Ano03a, Ano03d, Ano03e, Ano03f, Ano03b, Ano04b, Ano04c, Ano04a, Ano04d, Ano04e, Ano04f, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Gol07, Lie01, PY08b]. vs [GM06a, SM02].

W [Dei01, SEGG06]. wage [BH07, CL09, CL00]. wages [CL00]. Waiting [Coo08]. Wald [AMS07, AV09, Kem01, Per00, XP02b]. walk [EHL06, HLZ07, Lug03, Rot02]. walks [SS01]. Washington [HSWW08]. water [JR03]. wave [Pol00]. wavelet [CCZ08]. wavelets [FMRT09]. way [BBP06, Dav02d, YH00]. Weak [Jun08, AMS07, AS07, AM08, AMS08, Can07, CS07a, DT07, FW08b, HR07, JP09b, Kle07, MPS09, Mor09]. weakly [NS07a, Par06a]. wealthy [AHM^{+}03b]. weighted [FS08b, For05, HKZ06, LS07, Lin07, Woo07]. weighting [DS06]. weights [But00, OS06]. welfare [CR05, CH09, HSWW08]. whether [KS08b]. white [CL00, LM08b]. Whittle [ADG07, NS07b, SP06, Zaf09]. wild [DF08a]. wildcat [LPV00]. window [PT07b]. Winfrey [Bas03b]. wise [AHM^{+}03b, Pes07]. Wishart [GJS09]. within [HN00]. without [CCK00, Chi07, Shi01]. work [AAD05, CvdK08]. workers [JLS05, Jur02]. working [Bas08]. worklife [MNRS03]. world [BEP08]. Worst [San07]. wrongful [Lew03].

years [Ame09]. Yield [LMNT01, APW06, CP04b, DRA06, DLY08, Moe08]. yields [DL06, EHL06]. youth [DG08]. Yuan [Yu07].

Zellner [Ano05a, Ano07t]. zero [HZ07a, NS07a]. zero-inflated [HZ07a]. zero-information-limit [NS07a]. zone [LT06b].

References

Ashenfelter:2005:DUI

[AAD05] Orley Ashenfelter, David Ashmore, and Olivier Deschênes. Do unemployment insurance recipients actively seek work? Evidence from randomized trials in four U.S. states. Journal of Econometrics, 125(1–2):53–75, March 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

            URL http://www.sciencedirect.com/science/article/pii/
            S0304407604000740.

Ayat:2000:URT

[AB00] Leila Ayat and Peter Burridge. Unit root tests in the presence of uncertainty about the non-stochastic trend. Journal of Econometrics, 95(1):71–96, March 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000305.

Andrews:2001:ETS

[AB01] Donald W. K. Andrews and Moshe Buchinsky. Evaluation of a three-step method for choosing the number of bootstrap repetitions. Journal of Econometrics, 103(1–2):345–386, July 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000471.

Abadie:2003:SIV

[Aba03] Alberto Abadie. Semiparametric instrumental variable estimation of treatment response models. Journal of Econometrics, 113(2): 231–263, April 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602002014.

Aliprantis:2007:SIE

[ABCD07] Charalambos D. Aliprantis, William A. Barnett, Bernard Cornet, and Steven Durlauf. Special issue editors’ introduction: The interface between econometrics and economic theory. Journal of Econometrics, 136(2):325–329, February 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002162.

Andersen:2007:NAS

[ABD07] Torben G. Andersen, Tim Bollerslev, and Dobrislav Dobrev. No-arbitrage semi-martingale restrictions for continuous-time volatility models subject to leverage effects, jumps and i.i.d. noise: Theory and testable distributional implications. Journal of Econometrics, 138(1):125–180, May 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000911.

Andrade:2005:TCR

[ABG05] Philippe Andrade, Catherine Bruneau, and Stéphane Gregoir. Testing for the cointegration rank when some cointegrating directions are changing. Journal of Econometrics, 124(2):269–310, February 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760400034X.

Abrevaya:2000:REG

[Abr00] Jason Abrevaya. Rank estimation of a generalized fixed-effects regression model. Journal of Econometrics, 95(1):1–23, March 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000275.

Antoine:2007:EUI

[ABR07] Bertille Antoine, Hélène Bonnal, and Eric Renault. On the efficient use of the informational content of estimating equations: Implied probabilities and Euclidean empirical likelihood. Journal of Econometrics, 138(2):461–487, June 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000789.

Abdellaoui:2007:RIU

[ABW07] Mohammed Abdellaoui, Carolina Barrios, and Peter P. Wakker. Reconciling introspective utility with revealed preference: Experimental arguments based on prospect theory. Journal of Econometrics, 138(1):356–378, May 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606000984.

Altissimo:2003:SRD

[AC03a] Filippo Altissimo and Valentina Corradi. Strong rules for detecting the number of breaks in a time series. Journal of Econometrics, 117(2):207–244, December 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001477.

Arellano:2003:BCP

[AC03b] Manuel Arellano and Raquel Carrasco. Binary choice panel data models with predetermined variables. Journal of Econometrics, 115(1):125–157, July 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603000952.

Ai:2007:EPM

[AC07] Chunrong Ai and Xiaohong Chen. Estimation of possibly misspecified semiparametric conditional moment restriction models with different conditioning variables. Journal of Econometrics, 141(1): 5–43, November 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607000061.

Adda:2003:SES

[ACM03] Jérôme Adda, Tarani Chandola, and Michael Marmot. Socioeconomic status and health: causality and pathways. Journal of Econometrics, 112(1):57–63, January 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760200146X.

Allen:2008:FSP

[ACMP08] David Allen, Felix Chan, Michael McAleer, and Shelton Peiris. Finite sample properties of the QMLE for the log-ACD model: Application to Australian stocks. Journal of Econometrics, 147(1):163–185, November 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001334.

Atkinson:2005:BMP

[AD05] Scott E. Atkinson and Jeffrey H. Dorfman. Bayesian measurement of productivity and efficiency in the presence of undesirable outputs: crediting electric utilities for reducing air pollution. Journal of Econometrics, 126(2):445–468, June 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001198.

Abadir:2007:TJH

[AD07] Karim M. Abadir and Walter Distaso. Testing joint hypotheses when one of the alternatives is one-sided. Journal of Econometrics, 140(2):695–718, October 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001527.

Abadir:2007:NEL

[ADG07] Karim M. Abadir, Walter Distaso, and Liudas Giraitis. Nonstationarityextended local Whittle estimation. Journal of Econometrics, 141(2):1353–1384, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407607000474.

Abadir:2009:TEL

[ADG09] Karim M. Abadir, Walter Distaso, and Liudas Giraitis. Two estimators of the long-run variance: Beyond short memory. Journal of Econometrics, 150(1):56–70, May 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000451.

Ahn:2003:PHR

[ADGG03] Dong-Hyun Ahn, Robert F. Dittmar, A. Ronald Gallant, and Bin Gao. Purebred or hybrid?: Reproducing the volatility in term structure dynamics. Journal of Econometrics, 116(1–2):147–180, September 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603001064.

Andreou:2006:MDF

[AG06] Elena Andreou and Eric Ghysels. Monitoring disruptions in financial markets. Journal of Econometrics, 135(1–2):77–124, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001739.

Andreou:2008:QCS

[AG08] Elena Andreou and Eric Ghysels. Quality control for structural credit risk models. Journal of Econometrics, 146(2):364–375, October 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608001139.

Andrews:2009:IAS

[AG09] Donald W. K. Andrews and Patrik Guggenberger. Incorrect asymptotic size of subsampling procedures based on postconsistent model selection estimators. Journal of Econometrics, 152(1):19–27, September 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609000517.

Adams:2003:R

[AHM^{+}03a] P. Adams, M. D. Hurd, D. McFadden, A. Merrill, and T. Ribeirio. Response. Journal of Econometrics, 112(1):129–133, January 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602001562.

Adams:2003:HWW

[AHM^{+}03b] Peter Adams, Michael D. Hurd, Daniel McFadden, Angela Merrill, and Tiago Ribeiro. Healthy, wealthy, and wise? Tests for direct causal paths between health and socioeconomic status. Journal of Econometrics, 112(1):3–56, January 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001458.

Aue:2009:DTS

[AHR09] Alexander Aue, Lajos Horváth, and Matthew L. Reimherr. Delay times of sequential procedures for multiple time series regression models. Journal of Econometrics, 149(2):174–190, April 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760900027X.

Aliprantis:2007:RE

[AHT07] Charalambos D. Aliprantis, David Harris, and Rabee Tourky. Riesz estimators. Journal of Econometrics, 136(2):431–456, February 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605002204.

Aakvik:2005:ETE

[AHV05] Arild Aakvik, James J. Heckman, and Edward J. Vytlacil. Estimating treatment effects for discrete outcomes when responses to treatment vary: an application to Norwegian vocational rehabilitation programs. Journal of Econometrics, 125(1–2):15– 51, March 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604000739.

Anderson:2006:CF

[AIV06] Heather M. Anderson, João Victor Issler, and Farshid Vahid. Common features. Journal of Econometrics, 132(1):1–5, May 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605000382.

Andrews:2001:CMM

[AL01] Donald W. K. Andrews and Biao Lu. Consistent model and moment selection procedures for GMM estimation with application to dynamic panel data models. Journal of Econometrics, 101(1): 123–164, March 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000774.

Abadir:2004:CMM

[AL04] Karim M. Abadir and André Lucas. A comparison of minimum MSE and maximum power for the nearly integrated non-Gaussian model. Journal of Econometrics, 119(1):45–71, March 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001556.

Andrews:2006:HOI

[ALM06] Donald W. K. Andrews, Offer Lieberman, and Vadim Marmer. Higher-order improvements of the parametric bootstrap for long-memory Gaussian processes. Journal of Econometrics, 133(2):673–702, August 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605001302.

Ahn:2001:GEL

[ALS01] Seung Chan Ahn, Young Hoon Lee, and Peter Schmidt. GMM estimation of linear panel data models with time-varying individual effects. Journal of Econometrics, 101(2):219–255, April 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760000083X.

Anderson:2004:OTT

[ALS04] T. W. Anderson, R. A. Lockhart, and M. A. Stephens. An omnibus test for the time series model AR(1). Journal of Econometrics, 118(1–2):111–127, January 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001374.

Andrews:2008:EDF

[AM08] Donald W. K. Andrews and Vadim Marmer. Exactly distributionfree inference in instrumental variables regression with possibly weak instruments. Journal of Econometrics, 142(1):183–200, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607001261.

Asai:2009:SDC

[AM09] Manabu Asai and Michael McAleer. The structure of dynamic correlations in multivariate stochastic volatility models. Journal of Econometrics, 150(2):182–192, June 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002170.

Amemiya:2009:TFY

[Ame09] Takeshi Amemiya. Thirty-five years of journal of econometrics. Journal of Econometrics, 148(2):179–185, February 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002042.

Andrews:2007:PCW

[AMS07] Donald W. K. Andrews, Marcelo J. Moreira, and James H. Stock. Performance of conditional Wald tests in IV regression with weak instruments. Journal of Econometrics, 139(1):116–132, July 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407606001059.

Andrews:2008:ETS

[AMS08] Donald W. K. Andrews, Marcelo J. Moreira, and James H. Stock. Efficient two-sided nonsimilar invariant tests in IV regression with weak instruments. Journal of Econometrics, 146(2):241–254, October 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608001061.

Anderson:2002:RRR

[And02] T. W. Anderson. Reduced rank regression in cointegrated models. Journal of Econometrics, 106(2):203–216, February 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000951.

Anderson:2004:TEA

[And04] Gordon Anderson. Toward an empirical analysis of polarization. Journal of Econometrics, 122(1):1–26, September 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002677.

Anderson:2005:OLI

[And05] T. W. Anderson. Origins of the limited information maximum likelihood and two-stage least squares estimators. Journal of Econometrics, 127(1):1–16, July 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604001745.

Anderson:2006:RRR

[And06] T. W. Anderson. Reduced rank regression for blocks of simultaneous equations. Journal of Econometrics, 135(1–2):55–76, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001788.

Anonymous:2000:C

[Ano00a] Anonymous. Conference. Journal of Econometrics, 95(2):221, April 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000470.

Anonymous:2000:Ea

[Ano00b] Anonymous. Editorial. Journal of Econometrics, 94(1–2):1– 7, January 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407699000159.

Anonymous:2000:Eb

[Ano00c] Anonymous. Editorial. Journal of Econometrics, 95(2):219–220, April 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000378.

Anonymous:2000:Ia

[Ano00d] Anonymous. Index. Journal of Econometrics, 95(2):463, April 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000676.

Anonymous:2000:Ib

[Ano00e] Anonymous. Index. Journal of Econometrics, 96(2):395, June 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000178.

Anonymous:2000:Ic

[Ano00f] Anonymous. Index. Journal of Econometrics, 97(2):383, August 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000373.

Anonymous:2000:Id

[Ano00g] Anonymous. Index. Journal of Econometrics, 98(2):385, October 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000385.

Anonymous:2000:Ie

[Ano00h] Anonymous. Index. Journal of Econometrics, 99(2):387, December 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760000066X.

Anonymous:2000:PS

[Ano00i] Anonymous. Pages 1–186 (September 2000). Journal of Econometrics, 98(1):??, September 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2000:PN

[Ano00j] Anonymous. Pages 1–194 (November 2000). Journal of Econometrics, 99(1):??, November 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2000:PMb

[Ano00k] Anonymous. Pages 1–200 (May 2000). Journal of Econometrics, 96(1):??, May 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2000:PJc

[Ano00l] Anonymous. Pages 1–206 (July 2000). Journal of Econometrics, 97(1):??, July 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2000:PMa

[Ano00m] Anonymous. Pages 1–218 (March 2000). Journal of Econometrics, 95(1):??, March 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2000:PJa

[Ano00n] Anonymous. Pages 1–320 (January 2000). Journal of Econometrics, 94(1–2):??, January 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2000:PO

[Ano00o] Anonymous. Pages 187–386 (October 2000). Journal of Econometrics, 98(2):??, October 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2000:PD

[Ano00p] Anonymous. Pages 195–388 (December 2000). Journal of Econometrics, 99(2):??, December 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2000:PJb

[Ano00q] Anonymous. Pages 201–396 (June 2000). Journal of Econometrics, 96(2):??, June 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2000:PAb

[Ano00r] Anonymous. Pages 207–384 (August 2000). Journal of Econometrics, 97(2):??, August 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2000:PAa

[Ano00s] Anonymous. Pages 219–464 (April 2000). Journal of Econometrics, 95(2):??, April 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2001:AIa

[Ano01a] Anonymous. Author index. Journal of Econometrics, 101(2): 383, April 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000379.

Anonymous:2001:AIb

[Ano01b] Anonymous. Author index. Journal of Econometrics, 102(2): 399, June 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000653.

Anonymous:2001:AIc

[Ano01c] Anonymous. Author index. Journal of Econometrics, 103(1– 2):387, July 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000665.

Anonymous:2001:AId

[Ano01d] Anonymous. Author index. Journal of Econometrics, 104(2):407, September 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000975.

Anonymous:2001:AIV

[Ano01e] Anonymous. Author index to volume 105. Journal of Econometrics, 105(2):413, December 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407601001099.

Anonymous:2001:ET

[Ano01f] Anonymous. (estimation and testing). Journal of Econometrics, 103(1–2):??, July 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2001:Ia

[Ano01g] Anonymous. Index. Journal of Econometrics, 100(1):153–275, January 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000816.

Anonymous:2001:Ib

[Ano01h] Anonymous. Index. Journal of Econometrics, 100(1):277–318, January 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000804.

Anonymous:2001:Ic

[Ano01i] Anonymous. Index. Journal of Econometrics, 100(2):429, February 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000828.

Anonymous:2001:OF

[Ano01j] Anonymous. (open forum). Journal of Econometrics, 100(1):??, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2001:PMb

[Ano01k] Anonymous. Pages 1–142 (May 2001). Journal of Econometrics, 102(1):??, May 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2001:PMa

[Ano01l] Anonymous. Pages 1–194 (March 2001). Journal of Econometrics, 101(1):??, March 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2001:PAb

[Ano01m] Anonymous. Pages 1–208 (August 2001). Journal of Econometrics, 104(1):??, August 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2001:PJ

[Ano01n] Anonymous. Pages 143–400 (June 2001). Journal of Econometrics, 102(2):??, June 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2001:PAa

[Ano01o] Anonymous. Pages 195–384 (April 2001). Journal of Econometrics, 101(2):??, April 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2001:PD

[Ano01p] Anonymous. Pages 309–414 (December 2001). Journal of Econometrics, 105(2):??, December 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2001:PF

[Ano01q] Anonymous. Pages 319–430 (February 2001). Journal of Econometrics, 100(2):??, February 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2001:PEE

[Ano01r] Anonymous. Pages EX1–EX2, 209–412 (September 2001). Journal of Econometrics, 104(2):??, September 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2002:AIVa

[Ano02a] Anonymous. Author index to volume 106. Journal of Econometrics, 106(2):401, February 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407601001294.

Anonymous:2002:AIVb

[Ano02b] Anonymous. Author index to volume 107. Journal of Econometrics, 107(1–2):375, March 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602000660.

Anonymous:2002:AIVc

[Ano02c] Anonymous. Author index to volume 108. Journal of Econometrics, 108(2):395, June 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602000726.

Anonymous:2002:AIVd

[Ano02d] Anonymous. Author index to volume 109. Journal of Econometrics, 109(1):EX1–EX2, July 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602000866.

Anonymous:2002:AIVe

[Ano02e] Anonymous. Author index to volume 109. Journal of Econometrics, 109(2):393, August 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602001318.

Anonymous:2002:AIVf

[Ano02f] Anonymous. Author index to volume 110. Journal of Econometrics, 110(2):437, October 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602001793.

Anonymous:2002:AIVg

[Ano02g] Anonymous. Author index to volume 111. Journal of Econometrics, 111(2):385, December 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602002142.

Anonymous:2002:EM

[Ano02h] Anonymous. (EC2 meeting). Journal of Econometrics, 111(2): ??, December 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2002:EB

[Ano02i] Anonymous. Editorial Board. Journal of Econometrics, 111(2): ifc, December 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602002105.

Anonymous:2002:IIFa

[Ano02j] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 110(2):ifc, October 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001744.

Anonymous:2002:IIFb

[Ano02k] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 111(1):ifc, November 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001860.

Anonymous:2002:PS

[Ano02l] Anonymous. Pages 1–104 (September 2002). Journal of Econometrics, 110(1):??, September 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2002:PJa

[Ano02m] Anonymous. Pages 1–202 (January 2002). Journal of Econometrics, 106(1):??, January 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2002:PM

[Ano02n] Anonymous. Pages 1–202 (May 2002). Journal of Econometrics, 108(1):??, May 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2002:PJb

[Ano02o] Anonymous. Pages 203–396 (June 2002). Journal of Econometrics, 108(2):??, June 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2002:PEEd

[Ano02p] Anonymous. Pages EX1–EX2, 1–134 (November 2002). Journal of Econometrics, 111(1):??, November 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2002:PEEb

[Ano02q] Anonymous. Pages EX1–EX2, 1–194 (July 2002). Journal of Econometrics, 109(1):??, July 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2002:PEEc

[Ano02r] Anonymous. Pages EX1–EX2, 195–397 (August 2002). Journal of Econometrics, 109(2):??, August 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2002:PEEa

[Ano02s] Anonymous. Pages EX1–EX2, 203–402 (February 2002). Journal of Econometrics, 106(2):??, February 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2002:TAE

[Ano02t] Anonymous. Two adverts: Economics direct, authors. Journal of Econometrics, 109(2):EX1–EX2, August 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001355.

Anonymous:2003:AIVb

[Ano03a] Anonymous. Author index to volume. Journal of Econometrics, 113(2):399, April 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603000708.

Anonymous:2003:AIVf

[Ano03b] Anonymous. Author index to volume. Journal of Econometrics, 117(2):405, December 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603002331.

Anonymous:2003:AIVa

[Ano03c] Anonymous. Author index to volume 112. Journal of Econometrics, 112(2):401–402, February 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602002452.

Anonymous:2003:AIVc

[Ano03d] Anonymous. Author index to volume 114. Journal of Econometrics, 114(2):395, June 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603001271.

Anonymous:2003:AIVd

[Ano03e] Anonymous. Author index to volume 115. Journal of Econometrics, 115(2):391, August 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603001726.

Anonymous:2003:AIVe

[Ano03f] Anonymous. Author index to volume 116. Journal of Econometrics, 116(1–2):405, September 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603001830.

Anonymous:2003:EBa

[Ano03g] Anonymous. Editorial Board. Journal of Econometrics, 112(1): ifc, January 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602002269.

Anonymous:2003:EBb

[Ano03h] Anonymous. Editorial Board. Journal of Econometrics, 112(2): ifc, February 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602002427.

Anonymous:2003:EBc

[Ano03i] Anonymous. Editorial Board. Journal of Econometrics, 113(1): ifc, March 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602002506.

Anonymous:2003:EBd

[Ano03j] Anonymous. Editorial Board. Journal of Econometrics, 114 (1):ifc, May 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603000769.

Anonymous:2003:IIFa

[Ano03k] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 113(2):ifc, April 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603000678.

Anonymous:2003:IIFb

[Ano03l] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 114(2):ifc, June 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001246.

Anonymous:2003:IIFc

[Ano03m] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 115(1):ifc, July 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760300160X.

Anonymous:2003:IIFd

[Ano03n] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 115(2):ifc, August 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001696.

Anonymous:2003:IIFe

[Ano03o] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 116(1–2):ifc, September 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001775.

Anonymous:2003:IIFf

[Ano03p] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 117(1):ifc, November 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002197.

Anonymous:2003:IIFg

[Ano03q] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 117(2):ifc, December 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002306.

Anonymous:2003:PM

[Ano03r] Anonymous. Pages 1–196 (May 2003). Journal of Econometrics, 114(1):??, May 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2003:PA

[Ano03s] Anonymous. Pages 199–392 (August 2003). Journal of Econometrics, 115(2):??, August 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2003:PD

[Ano03t] Anonymous. Pages 207–406 (December 2003). Journal of Econometrics, 117(2):??, December 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2003:PF

[Ano03u] Anonymous. Pages 241–402 (February 2003). Journal of Econometrics, 112(2):??, February 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2003:PEEc

[Ano03v] Anonymous. Pages EX1–EX2, 1–198 (July 2003). Journal of Econometrics, 115(1):??, July 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2003:PEEb

[Ano03w] Anonymous. Pages EX1–EX2, 197–396 (June 2003). Journal of Econometrics, 114(2):??, June 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2003:PEEa

[Ano03x] Anonymous. Pages EX1–EX2, 201-400 (april 2003). Journal of Econometrics, 113(2):??, April 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2003:PEEd

[Ano03y] Anonymous. Pages EX1–EX6, 1–206 (November 2003). Journal of Econometrics, 117(1):??, November 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2004:AIVc

[Ano04a] Anonymous. Author index to volume. Journal of Econometrics, 120(2):361, June 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604000570.

Anonymous:2004:AIVa

[Ano04b] Anonymous. Author index to volume 118. Journal of Econometrics, 118(1–2):375, January 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603002446.

Anonymous:2004:AIVb

[Ano04c] Anonymous. Author index to volume 119. Journal of Econometrics, 119(2):435, April 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604000181.

Anonymous:2004:AIVd

[Ano04d] Anonymous. Author index to volume 121. Journal of Econometrics, 121(1–2):427, July 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604000685.

Anonymous:2004:AIVe

[Ano04e] Anonymous. Author index to volume 122. Journal of Econometrics, 122(2):405, October 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604000983.

Anonymous:2004:AIVf

[Ano04f] Anonymous. Author index to volume 123. Journal of Econometrics, 123(2):393, December 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604001289.

Anonymous:2004:IIFa

[Ano04g] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 118(1–2):ifc, January 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002380.

Anonymous:2004:IIFb

[Ano04h] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 119(1):ifc, March 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760400003X.

Anonymous:2004:IIFc

[Ano04i] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 119(2):ifc, April 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000144.

Anonymous:2004:IIFd

[Ano04j] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 120(1):ifc, May 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000442.

Anonymous:2004:IIFe

[Ano04k] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 120(2):ifc, June 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000545.

Anonymous:2004:IIFg

[Ano04l] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 122(1):ifc, September 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000867.

Anonymous:2004:IIFh

[Ano04m] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 122(2):ifc, October 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000958.

Anonymous:2004:IIFi

[Ano04n] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 123(1):ifc, November 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/S0304407604001034.

Anonymous:2004:IIFj

[Ano04o] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 123(2):ifc, December 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001253.

Anonymous:2004:IIFf

[Ano04p] Anonymous. IFC: Inside front cover editorial board. Journal of Econometrics, 121(1–2):ifc, July 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604000624.

Anonymous:2004:PS

[Ano04q] Anonymous. Pages 1–212 (September 2004). Journal of Econometrics, 122(1):??, September 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2004:PO

[Ano04r] Anonymous. Pages 213–406 (October 2004). Journal of Econometrics, 122(2):??, October 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2004:PEE

[Ano04s] Anonymous. Pages EX1–EX2, 1–194 (November 2004). Journal of Econometrics, 123(1):??, November 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2004:PEEb

[Ano04t] Anonymous. Pages EX1–EX2, 1–206 (May 2004). Journal of Econometrics, 120(1):??, May 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2004:PEEa

[Ano04u] Anonymous. Pages EX1–EX2, 1–222 (March 2004). Journal of Econometrics, 119(1):??, March 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2004:PEEc

[Ano04v] Anonymous. Pages EX1–EX2, 207–362 (June 2004). Journal of Econometrics, 120(2):??, June 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2004:TC

[Ano04w] Anonymous. Table of contents. Journal of Econometrics, 123(2): 195, December 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603002963.

Anonymous:2005:AZA

[Ano05a] Anonymous. The Arnold zellner award. Journal of Econometrics, 126(2):573, June 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604002143.

Anonymous:2005:AIVa

[Ano05b] Anonymous. Author index to volume 124. Journal of Econometrics, 124(2):395, February 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604001903.

Anonymous:2005:AIVb

[Ano05c] Anonymous. Author index to volume 125. Journal of Econometrics, 125(1–2):377, March 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604002015.

Anonymous:2005:AIVc

[Ano05d] Anonymous. Author index to volume 126. Journal of Econometrics, 126(2):587–588, June 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760500028X.

Anonymous:2005:AIVd

[Ano05e] Anonymous. Author index to volume 127. Journal of Econometrics, 127(2):253, August 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605001004.

Anonymous:2005:AIVe

[Ano05f] Anonymous. Author index to volume 128. Journal of Econometrics, 128(2):325, October 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605001855.

Anonymous:2005:AIVf

[Ano05g] Anonymous. Author index to volume 129. Journal of Econometrics, 129(1–2):373, November 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760500196X.

Anonymous:2005:DJA

[Ano05h] Anonymous. The Dennis J. Aigner Award. Journal of Econometrics, 127(1):129, July 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605000023.

Anonymous:2005:E

[Ano05i] Anonymous. Erratum. Journal of Econometrics, 128(1):i, September 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605001193.

Anonymous:2005:FJE

[Ano05j] Anonymous. Fellows of the Journal of Econometrics as of 2005. Journal of Econometrics, 126(2):575–586, June 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604002155.

Anonymous:2005:IIFa

[Ano05k] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 124(1):ifc, January 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001769.

Anonymous:2005:IIFb

[Ano05l] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 124(2):ifc, February 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001873.

Anonymous:2005:IIFc

[Ano05m] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 125(1–2):ifc, March 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001952.

Anonymous:2005:IIFd

[Ano05n] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 126(1):ifc, May 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604002064.

Anonymous:2005:IIFe

[Ano05o] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 126(2):CO2, June 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000242.

Anonymous:2005:IIFf

[Ano05p] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 127(1):CO2, July 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000680.

Anonymous:2005:IIFg

[Ano05q] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 127(2):CO2, August 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000977.

Anonymous:2005:IIFh

[Ano05r] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 128(1):CO2, September 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001119.

Anonymous:2005:IIFi

[Ano05s] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 128(2):CO2, October 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760500182X.

Anonymous:2005:IIFj

[Ano05t] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 129(1–2):CO2, November 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001909.

Anonymous:2005:MSB

[Ano05u] Anonymous. Modelling structural breaks. Journal of Econometrics, 129(1–2):??, November 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2005:PJb

[Ano05v] Anonymous. Pages 1–130 (July 2005). Journal of Econometrics, 127(1):??, July 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2005:PS

[Ano05w] Anonymous. Pages 1–194 (September 2005). Journal of Econometrics, 128(1):??, September 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2005:PJa

[Ano05x] Anonymous. Pages 1–204 (January 2005). Journal of Econometrics, 124(1):??, January 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2005:PM

[Ano05y] Anonymous. Pages 1–232 (May 2005). Journal of Econometrics, 126(1):??, May 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2005:PA

[Ano05z] Anonymous. Pages 131–254 (August 2005). Journal of Econometrics, 127(2):??, August 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2005:PO

[Ano05-27] Anonymous. Pages 195–326 (October 2005). Journal of Econometrics, 128(2):??, October 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2005:PEE

[Ano05-28] Anonymous. Pages EX1–EX2, 205–396 (February 2005). Journal of Econometrics, 124(2):??, February 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2006:AIVa

[Ano06a] Anonymous. Author index to volume 130. Journal of Econometrics, 130(2):385, February 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606000066.

Anonymous:2006:AIVb

[Ano06b] Anonymous. Author index to volume 131. Journal of Econometrics, 131(1–2):611–612, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760600025X.

Anonymous:2006:AIVc

[Ano06c] Anonymous. Author index to volume 132. Journal of Econometrics, 132(2):545–546, June 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606001151.

Anonymous:2006:AIVd

[Ano06d] Anonymous. Author index to volume 133. Journal of Econometrics, 133(2):887–888, August 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606001357.

Anonymous:2006:AIVe

[Ano06e] Anonymous. Author index to volume 134. Journal of Econometrics, 134(2):665–666, October 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606001758.

Anonymous:2006:EBa

[Ano06f] Anonymous. Editorial Board. Journal of Econometrics, 130(2): CO2, February 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000030.

Anonymous:2006:EBb

[Ano06g] Anonymous. Editorial Board. Journal of Econometrics, 131 (1–2):CO2, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606000194.

Anonymous:2006:EBc

[Ano06h] Anonymous. Editorial Board. Journal of Econometrics, 132(1): CO2, May 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000571.

Anonymous:2006:EBd

[Ano06i] Anonymous. Editorial Board. Journal of Econometrics, 132(2): CO2, June 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606001114.

Anonymous:2006:EBe

[Ano06j] Anonymous. Editorial Board. Journal of Econometrics, 133(1): CO2, July 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606001217.

Anonymous:2006:EBf

[Ano06k] Anonymous. Editorial Board. Journal of Econometrics, 133(2): CO2, August 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760600131X.

Anonymous:2006:EBg

[Ano06l] Anonymous. Editorial Board. Journal of Econometrics, 134(1):CO2, September 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606001631.

Anonymous:2006:EBh

[Ano06m] Anonymous. Editorial Board. Journal of Econometrics, 134(2): CO2, October 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606001722.

Anonymous:2006:EBi

[Ano06n] Anonymous. Editorial Board. Journal of Econometrics, 135(1– 2):CO2, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606001801.

Anonymous:2006:IIF

[Ano06o] Anonymous. IFC — inside front cover — Editorial Board. Journal of Econometrics, 130(1):CO2, January 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002083.

Anonymous:2006:PJa

[Ano06p] Anonymous. Pages 1–208 (January 2006). Journal of Econometrics, 130(1):??, January 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2006:PS

[Ano06q] Anonymous. Pages 1–316 (September 2006). Journal of Econometrics, 134(1):??, September 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2006:PJc

[Ano06r] Anonymous. Pages 1–410 (July 2006). Journal of Econometrics, 133(1):??, July 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2006:PND

[Ano06s] Anonymous. Pages 1–568 (November–December 2006). Journal of Econometrics, 135(1–2):??, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2006:PMA

[Ano06t] Anonymous. Pages 1–612 (March–April 2006). Journal of Econometrics, 131(1–2):??, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2006:PF

[Ano06u] Anonymous. Pages 209–386 (February 2006). Journal of Econometrics, 130(2):??, February 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2006:PJb

[Ano06v] Anonymous. Pages 305–546 (June 2006). Journal of Econometrics, 132(2):??, June 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2006:PO

[Ano06w] Anonymous. Pages 317–666 (October 2006). Journal of Econometrics, 134(2):??, October 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2006:PA

[Ano06x] Anonymous. Pages 411–888 (August 2006). Journal of Econometrics, 133(2):??, August 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2007:EBa

[Ano07a] Anonymous. Editorial Board. Journal of Econometrics, 136(1): CO2, January 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606001916.

Anonymous:2007:EBb

[Ano07b] Anonymous. Editorial Board. Journal of Econometrics, 136(2): CO2, February 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606002405.

Anonymous:2007:EBc

[Ano07c] Anonymous. Editorial Board. Journal of Econometrics, 137(1): CO2, March 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607000188.

Anonymous:2007:EBd

[Ano07d] Anonymous. Editorial Board. Journal of Econometrics, 137(2): CO2, April 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607000309.

Anonymous:2007:EBe

[Ano07e] Anonymous. Editorial Board. Journal of Econometrics, 138 (1):ifc, May 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607000504.

Anonymous:2007:EBf

[Ano07f] Anonymous. Editorial Board. Journal of Econometrics, 138(2): ifc, June 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607000693.

Anonymous:2007:EBg

[Ano07g] Anonymous. Editorial Board. Journal of Econometrics, 139(1):ifc, July 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S030440760700084X.

Anonymous:2007:EBh

[Ano07h] Anonymous. Editorial Board. Journal of Econometrics, 139(2): ifc, August 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760700098X.

Anonymous:2007:EBi

[Ano07i] Anonymous. Editorial Board. Journal of Econometrics, 140(1): ifc, September 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001340.

Anonymous:2007:EBj

[Ano07j] Anonymous. Editorial Board. Journal of Econometrics, 140(2): ifc, October 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001492.

Anonymous:2007:EBk

[Ano07k] Anonymous. Editorial Board. Journal of Econometrics, 141(1): ifc, November 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001881.

Anonymous:2007:EBl

[Ano07l] Anonymous. Editorial Board. Journal of Econometrics, 141(2): ifc, December 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607002084.

Anonymous:2007:ECL

[Ano07m] Anonymous. Error in contents listing of special issue. Journal of Econometrics, 141(2):1420, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607002072.

Anonymous:2007:PM

[Ano07n] Anonymous. Pages 1–276 (March 2007). Journal of Econometrics, 137(1):??, March 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2007:PJ

[Ano07o] Anonymous. Pages 1–324 (January 2007). Journal of Econometrics, 136(1):??, January 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2007:PA

[Ano07p] Anonymous. Pages 277–728 (April 2007). Journal of Econometrics, 137(2):??, April 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2007:PD

[Ano07q] Anonymous. Pages 323–1420 (December 2007). Journal of Econometrics, 141(2):??, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2007:PO

[Ano07r] Anonymous. Pages 333–948 (October 2007). Journal of Econometrics, 140(2):??, October 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2007:RUR

[Ano07s] Anonymous. Referee utilization report. Journal of Econometrics, 141(1):320–322, November 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760700200X.

Anonymous:2007:VHA

[Ano07t] Anonymous. A volume in honor of Arnold zellner. Journal of Econometrics, 138(2):??, June 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2008:AA

[Ano08a] Anonymous. Aigner award. Journal of Econometrics, 142(2): i, February 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607002382.

Anonymous:2008:EBa

[Ano08b] Anonymous. Editorial Board. Journal of Econometrics, 142(1): ifc, January 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607002187.

Anonymous:2008:EBb

[Ano08c] Anonymous. Editorial Board. Journal of Econometrics, 142(2): CO2, February 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607002321.

Anonymous:2008:EBc

[Ano08d] Anonymous. Editorial Board. Journal of Econometrics, 143(1): ifc, March 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607002461.

Anonymous:2008:EBd

[Ano08e] Anonymous. Editorial Board. Journal of Econometrics, 143(2): ifc, April 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000109.

Anonymous:2008:EBe

[Ano08f] Anonymous. Editorial Board. Journal of Econometrics, 144 (1):ifc, May 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000249.

Anonymous:2008:EBf

[Ano08g] Anonymous. Editorial Board. Journal of Econometrics, 144(2): ifc, June 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760800064X.

Anonymous:2008:EBg

[Ano08h] Anonymous. Editorial Board. Journal of Econometrics, 145(1– 2):ifc, July 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000857.

Anonymous:2008:EBh

[Ano08i] Anonymous. Editorial Board. Journal of Econometrics, 146(1): ifc, September 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001565.

Anonymous:2008:EBi

[Ano08j] Anonymous. Editorial Board. Journal of Econometrics, 146(2): ifc, October 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001711.

Anonymous:2008:EBj

[Ano08k] Anonymous. Editorial Board. Journal of Econometrics, 147(1): ifc, November 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760800184X.

Anonymous:2008:EBk

[Ano08l] Anonymous. Editorial Board. Journal of Econometrics, 147(2): ifc, December 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001942.

Anonymous:2008:FL

[Ano08m] Anonymous. Fellows list. Journal of Econometrics, 142(2):III– VIII, February 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607002400.

Anonymous:2008:NF

[Ano08n] Anonymous. New fellows. Journal of Econometrics, 142(2):ii, February 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607002394.

Anonymous:2008:PS

[Ano08o] Anonymous. Pages 1–198 (September 2008). Journal of Econometrics, 146(1):??, September 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2008:PM

[Ano08p] Anonymous. Pages 1–324 (May 2008). Journal of Econometrics, 144(1):??, May 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2008:PJa

[Ano08q] Anonymous. Pages 1–610 (January 2008). Journal of Econometrics, 142(1):??, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2008:PA

[Ano08r] Anonymous. Pages 227–398 (April 2008). Journal of Econometrics, 143(2):??, April 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2008:PJb

[Ano08s] Anonymous. Pages 325–528 (June 2008). Journal of Econometrics, 144(2):??, June 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2009:A

[Ano09a] Anonymous. Announcement. Journal of Econometrics, 148(1): v–x, January 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609000116.

Anonymous:2009:EBa

[Ano09b] Anonymous. Editorial Board. Journal of Econometrics, 148(1): ifc, January 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609000037.

Anonymous:2009:EBb

[Ano09c] Anonymous. Editorial Board. Journal of Econometrics, 148(2): ifc, February 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609000359.

Anonymous:2009:EBc

[Ano09d] Anonymous. Editorial Board. Journal of Econometrics, 149(1): ifc, April 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609000608.

Anonymous:2009:EBd

[Ano09e] Anonymous. Editorial Board. Journal of Econometrics, 149(2): ifc, April 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609000876.

Anonymous:2009:EBe

[Ano09f] Anonymous. Editorial Board. Journal of Econometrics, 150 (1):ifc, May 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001018.

Anonymous:2009:EBf

[Ano09g] Anonymous. Editorial Board. Journal of Econometrics, 150(2): ifc, June 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001146.

Anonymous:2009:EBg

[Ano09h] Anonymous. Editorial Board. Journal of Econometrics, 151 (1):ifc, July 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001304.

Anonymous:2009:EBh

[Ano09i] Anonymous. Editorial Board. Journal of Econometrics, 151(2): ifc, August 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001535.

Anonymous:2009:EBi

[Ano09j] Anonymous. Editorial Board. Journal of Econometrics, 152(1): ifc, September 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001699.

Anonymous:2009:EBj

[Ano09k] Anonymous. Editorial Board. Journal of Econometrics, 152(2): ifc, October 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001808.

Anonymous:2009:EBk

[Ano09l] Anonymous. Editorial Board. Journal of Econometrics, 153(1): ifc, November 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609002000.

Anonymous:2009:EBl

[Ano09m] Anonymous. Editorial Board. Journal of Econometrics, 153(2): ifc, December 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609002218.

Anonymous:2009:PAa

[Ano09n] Anonymous. Pages 1–100 (April 2009). Journal of Econometrics, 149(1):??, April 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2009:PJa

[Ano09o] Anonymous. Pages 1–100 (January 2009). Journal of Econometrics, 148(1):??, January 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2009:PN

[Ano09p] Anonymous. Pages 1–104 (November 2009). Journal of Econometrics, 153(1):??, November 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2009:PM

[Ano09q] Anonymous. Pages 1–116 (May 2009). Journal of Econometrics, 150(1):??, May 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2009:PJb

[Ano09r] Anonymous. Pages 1–98 (July 2009). Journal of Econometrics, 151(1):??, July 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2009:PF

[Ano09s] Anonymous. Pages 101–200 (February 2009). Journal of Econometrics, 148(2):??, February 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2009:PAb

[Ano09t] Anonymous. Pages 101–208 (April 2009). Journal of Econometrics, 149(2):??, April 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2009:PD

[Ano09u] Anonymous. Pages 105–210 (December 2009). Journal of Econometrics, 153(2):??, December 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Antweiler:2001:NRE

[Ant01] Werner Antweiler. Nested random effects estimation in unbalanced panel data. Journal of Econometrics, 101(2):295–313, April 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000865.

Atkinson:2002:SEF

[AP02] Scott E. Atkinson and Daniel Primont. Stochastic estimation of firm technology, inefficiency, and productivity growth using shadow cost and distance functions. Journal of Econometrics, 108(2):203–225, June 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407601001336.

Ang:2006:WDY

[APW06] Andrew Ang, Monika Piazzesi, and Min Wei. What does the yield curve tell us about GDP growth? Journal of Econometrics, 131(1–2):359–403, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000163.

Arcidiacono:2004:ASR

[Arc04] Peter Arcidiacono. Ability sorting and the returns to college major. Journal of Econometrics, 121(1–2):343–375, July 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760300263X.

Arteche:2004:GSE

[Art04] Josu Arteche. Gaussian semiparametric estimation in long memory in stochastic volatility and signal plus noise models. Journal of Econometrics, 119(1):131–154, March 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001581.

Andrews:2007:TMW

[AS07] Donald W. K. Andrews and James H. Stock. Testing with many weak instruments. Journal of Econometrics, 138(1):24–46, May 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760600087X.

Ait-Sahalia:2001:GFT

[ASBS01] Yacine Aït-Sahalia, Peter J. Bickel, and Thomas M. Stoker. Goodness-of-fit tests for kernel regression with an application to option implied volatilities. Journal of Econometrics, 105(2):363– 412, December 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000914.

Ait-Sahalia:2003:NOP

[ASD03] Yacine Aït-Sahalia and Jefferson Duarte. Nonparametric option pricing under shape restrictions. Journal of Econometrics, 116(1–2):9–47, September 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603001027.

Ait-Sahalia:2000:NRM

[ASL00] Yacine Aït-Sahalia and Andrew W. Lo. Nonparametric risk management and implied risk aversion. Journal of Econometrics, 94(1–2):9–51, January 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407699000160.

Ait-Sahalia:2008:SFQ

[ASM08a] Yacine Aït-Sahalia and Loriano Mancini. Out of sample forecasts of quadratic variation. Journal of Econometrics, 147(1):17–33, November 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001218.

Ait-Sahalia:2008:AHS

[ASM08b] Yacine Aiẗ-Sahalia and Per A. Mykland. An analysis of Hansenscheinkman moment estimators for discretely and randomly sampled diffusions. Journal of Econometrics, 144(1):1–26, May 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760700228X.

Ait-Sahalia:2001:DOM

[ASWY01] Yacine Aït-Sahalia, Yubo Wang, and Francis Yared. Do option markets correctly price the probabilities of movement of the underlying asset? Journal of Econometrics, 102(1):67–110, May 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000919.

Ait-Sahalia:2006:SAC

[ASY06] Yacine Aiẗ-Sahalia and Jialin Yu. Saddlepoint approximations for continuous-time Markov processes. Journal of Econometrics, 134(2):507–551, October 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605001545.

Abadir:2005:AFS

[AT05] Karim M. Abadir and Gabriel Talmain. Autocovariance functions of series and of their transforms. Journal of Econometrics, 124(2):227–252, February 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604000326.

Aiolfi:2006:PFP

[AT06] Marco Aiolfi and Allan Timmermann. Persistence in forecasting performance and conditional combination strategies. Journal of Econometrics, 135(1–2):31–53, November/December 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001661.

Ahlin:2007:SAC

[AT07] Christian Ahlin and Robert M. Townsend. Selection into and across credit contracts: Theory and field research. Journal of Econometrics, 136(2):665–698, February 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002289.

Avarucci:2009:WTC

[AV09] Marco Avarucci and Carlos Velasco. A Wald test for the cointegration rank in nonstationary fractional systems. Journal of Econometrics, 151(2):178–189, August 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000815.

Bai:2004:ECS

[Bai04] Jushan Bai. Estimating cross-section common stochastic trends in nonstationary panel data. Journal of Econometrics, 122(1):137– 183, September 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603002720.

Banerjee:2007:MEA

[Ban07] Anurag Banerjee. A method of estimating the average derivative. Journal of Econometrics, 136(1):65–88, January 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001600.

Barnett:2002:TTC

[Bar02] William A. Barnett. Tastes and technology: curvature is not sufficient for regularity. Journal of Econometrics, 108(1):199– 202, May 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601001312.

Barnett:2007:MAT

[Bar07] William A. Barnett. Multilateral aggregation-theoretic monetary aggregation over heterogeneous countries. Journal of Econometrics, 136(2):457–482, February 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002216.

Basmann:2003:ISE

[Bas03a] Robert L. Basmann. Introduction to statistics and econometrics in litigation support. Journal of Econometrics, 113(1):1– 2, March 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760200163X.

Basmann:2003:SOA

[Bas03b] Robert L. Basmann. Statistical outlier analysis in litigation support: the case of Paul F. Engler and Cactus Feeders, Inc., v. Oprah Winfrey et al. Journal of Econometrics, 113(1):159–200, March 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602001719.

Basmann:2008:CSM

[Bas08] R. L. Basmann. Chamberlin’s strategy of multiple working hypotheses and a relative frequency theory of market demand. Journal of Econometrics, 147(2):225–231, December 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001383.

Bates:2000:PCF

[Bat00] David S. Bates. Post-’87 crash fears in the S&P 500 futures option market. Journal of Econometrics, 94(1–2):181–238, January 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000214.

Bates:2003:EOP

[Bat03] David S. Bates. Empirical option pricing: a retrospection. Journal of Econometrics, 116(1–2):387–404, September 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001131.

Bera:2002:MMM

[BB02] Anil K. Bera and Yannis Bilias. The MM, ME, ML, EL, EF and GMM approaches to estimation: a synthesis. Journal of Econometrics, 107(1–2):51–86, March 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001130.

Baltagi:2006:JLT

[BBP06] Badi H. Baltagi, Georges Bresson, and Alain Pirotte. Joint LM test for homoskedasticity in a one-way error component model. Journal of Econometrics, 134(2):401–417, October 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001508.

Bauwens:2006:CEE

[BBU06] Luc Bauwens, H. Peter Boswijk, and Jean-Pierre Urbain. Causality and exogeneity in econometrics. Journal of Econometrics, 132(2):305–309, June 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605000564.

Bauwens:2004:ARB

[BBvDvO04] Luc Bauwens, Charles S. Bos, Herman K. van Dijk, and Rutger D. van Oest. Adaptive radial-based direction sampling: some flexible and robust Monte Carlo integration methods. Journal of Econometrics, 123(2):201–225, December 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002884.

Breitung:2006:TSL

[BC06] Jörg Breitung and Bertrand Candelon. Testing for shortand long-run causality: A frequency-domain approach. Journal of Econometrics, 132(2):363–378, June 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760500059X.

Bai:2008:TMD

[BC08a] Jushan Bai and Zhihong Chen. Testing multivariate distributions in GARCH models. Journal of Econometrics, 143(1):19– 36, March 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001650.

Browning:2008:LRC

[BC08b] Martin Browning and Thomas F. Crossley. The long-run cost of job loss as measured by consumption changes. Journal of Econometrics, 145(1–2):109–120, July 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000560.

Bakshi:2000:PHL

[BCC00] Gurdip Bakshi, Charles Cao, and Zhiwu Chen. Pricing and hedging long-term options. Journal of Econometrics, 94(1–2):277– 318, January 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407699000238.

Bilias:2000:SRM

[BCY00] Yannis Bilias, Songnian Chen, and Zhiliang Ying. Simple resampling methods for censored regression quantiles. Journal of Econometrics, 99(2):373–386, December 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000427.

Brockwell:2004:GLD

[BD04] P. J. Brockwell and R. Dahlhaus. Generalized Levinson–Durbin and Burg algorithms. Journal of Econometrics, 118(1–2):129– 149, January 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603001386.

Brock:2007:IBC

[BD07] William A. Brock and Steven N. Durlauf. Identification of binary choice models with social interactions. Journal of Econometrics, 140(1):52–75, September 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002235.

Bramoulle:2009:IPE

[BDF09] Yann Bramoullé, Habiba Djebbari, and Bernard Fortin. Identification of peer effects through social networks. Journal of Econometrics, 150(1):41–55, May 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000335.

Broadie:2000:AOS

[BDGT00] Mark Broadie, Jérôme Detemple, Eric Ghysels, and Olivier Torrés. American options with stochastic dividends and volatility: A nonparametric investigation. Journal of Econometrics, 94(1–2): 53–92, January 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407699000172.

Brock:2007:MUP

[BDW07] William A. Brock, Steven N. Durlauf, and Kenneth D. West. Model uncertainty and policy evaluation: Some theory and empirics. Journal of Econometrics, 136(2):629–664, February 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002277.

Bekker:2002:EIL

[Bek02] Paul A. Bekker. Exact inference for the linear model with groupwise heteroscedastic spherical disturbances. Journal of Econometrics, 111(2):285–302, December 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001070.

Baltagi:2007:EMC

[BEP07] Badi H. Baltagi, Peter Egger, and Michael Pfaffermayr. Estimating models of complex FDI: Are there third-country effects? Journal of Econometrics, 140(1):260–281, September 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002314.

Baltagi:2008:ERT

[BEP08] Badi H. Baltagi, Peter Egger, and Michael Pfaffermayr. Estimating regional trade agreement effects on FDI in an interdependent world. Journal of Econometrics, 145(1–2):194–208, July 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608000559.

Berkowitz:2001:GSE

[Ber01] Jeremy Berkowitz. Generalized spectral estimation of the consumption-based asset pricing model. Journal of Econometrics, 104(2):269–288, September 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407601000811.

Bec:2008:ACU

[BGG08] Frédérique Bec, Alain Guay, and Emmanuel Guerre. Adaptive consistent unit-root tests based on autoregressive threshold model. Journal of Econometrics, 142(1):94–133, January 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001236.

Bound:2004:TUT

[BGKT04] John Bound, Jeffrey Groen, Gábor Kézdi, and Sarah Turner. Trade in university training: cross-state variation in the production and stock of college-educated labor. Journal of Econometrics, 121(1–2):143–173, July 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002604.

Bollen:2000:RSF

[BGW00] Nicolas P. B. Bollen, Stephen F. Gray, and Robert E. Whaley. Regime switching in foreign exchange rates:: Evidence from currency option prices. Journal of Econometrics, 94(1–2):239– 276, January 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407699000226.

Blundell:2002:IED

[BGW02] Richard Blundell, Rachel Griffith, and Frank Windmeijer. Individual effects and dynamics in count data models. Journal of Econometrics, 108(1):113–131, May 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001087.

Breitung:2002:ICR

[BH02] Jörg Breitung and Uwe Hassler. Inference on the cointegration rank in fractionally integrated processes. Journal of Econometrics, 110(2):167–185, October 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760200091X.

Belzil:2007:SAC

[BH07] Christian Belzil and Jörgen Hansen. A structural analysis of the correlated random coefficient wage regression model. Journal of Econometrics, 140(2):827–848, October 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001576.

Bhargava:2003:ADH

[Bha03a] Alok Bhargava. Analysis of data on health: 2. Journal of Econometrics, 112(1):1, January 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602001446.

Bhargava:2003:FPG

[Bha03b] Alok Bhargava. Family planning, gender differences and infant mortality: evidence from uttar pradesh, India. Journal of Econometrics, 112(1):225–240, January 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001628.

Bhattacharya:2005:AIM

[Bha05] Debopam Bhattacharya. Asymptotic inference from multistage samples. Journal of Econometrics, 126(1):145–171, May 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001332.

Bhattacharya:2007:IIH

[Bha07] Debopam Bhattacharya. Inference on inequality from household survey data. Journal of Econometrics, 137(2):674–707, April 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606000728.

Bhattacharya:2008:IPD

[Bha08] Debopam Bhattacharya. Inference in panel data models under attrition caused by unobservables. Journal of Econometrics, 144(2):

430–446, June 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000389.

Burda:2008:BML

[BHH08] Martin Burda, Matthew Harding, and Jerry Hausman. A Bayesian mixed logit-probit model for multinomial choice. Journal of Econometrics, 147(2):232–246, December 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001395.

Bertail:2004:SDD

[BHPW04] Patrice Bertail, Christian Haefke, Dimitris N. Politis, and Halbert White. Subsampling the distribution of diverging statistics with applications to finance. Journal of Econometrics, 120(2): 295–326, June 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760300215X.

Biewen:2002:BII

[Bie02] Martin Biewen. Bootstrap inference for inequality, mobility and poverty measurement. Journal of Econometrics, 108(2):317–342, June 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601001385.

Bierens:2007:EAL

[Bie07] Herman J. Bierens. Econometric analysis of linearized singular dynamic stochastic general equilibrium models. Journal of Econometrics, 136(2):595–627, February 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002265.

Bernard:2007:FSM

[BIKY07] Jean-Thomas Bernard, Nadhem Idoudi, Lynda Khalaf, and Clément Yélou. Finite sample multivariate structural change tests with application to energy demand models. Journal of Econometrics, 141(2):1219–1244, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000413.

Biorn:2004:RSU

[Biø04] Erik Biørn. Regression systems for unbalanced panel data: a stepwise maximum likelihood procedure. Journal of Econometrics, 122(2):281–291, October 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603002768.

Bun:2006:EDF

[BK06] Maurice J. G. Bun and Jan F. Kiviet. The effects of dynamic feedbacks on LS and MM estimator accuracy in panel data models. Journal of Econometrics, 132(2):409–444, June 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000618.

Blake:2007:TAP

[BK07] Andrew P. Blake and George Kapetanios. Testing for ARCH in the presence of nonlinearity of unknown form in the conditional mean. Journal of Econometrics, 137(2):472–488, April 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606000558.

Baillie:2008:NMS

[BK08] Richard T. Baillie and George Kapetanios. Nonlinear models for strongly dependent processes with financial applications. Journal of Econometrics, 147(1):60–71, November 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001243.

Bai:2009:PCG

[BKN09] Jushan Bai, Chihwa Kao, and Serena Ng. Panel cointegration with global stochastic trends. Journal of Econometrics, 149(1): 82–99, April 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608002066.

Baltagi:2007:ASD

[BKP07] Badi H. Baltagi, Harry H. Kelejian, and Ingmar R. Prucha. Analysis of spatially dependent data. Journal of Econometrics, 140(1): 1–4, September 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606002211.

Bollerslev:2009:DTM

[BKPT09] Tim Bollerslev, Uta Kretschmer, Christian Pigorsch, and George Tauchen. A discrete-time model for daily s and p500 returns and realized variations: Jumps and leverage effects. Journal of Econometrics, 150(2):151–166, June 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002157.

Bansal:2002:MEA

[BL02a] Ravi Bansal and Christian Lundblad. Market efficiency, asset returns, and the size of the risk premium in global equity markets. Journal of Econometrics, 109(2):195–237, August 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602000672.

Boswijk:2002:SNC

[BL02b] H. Peter Boswijk and André Lucas. Semi-nonparametric cointegration testing. Journal of Econometrics, 108(2):253–280, June 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601001361.

Bekker:2008:SBI

[BL08] Paul A. Bekker and Steve Lawford. Symmetry-based inference in an instrumental variable setting. Journal of Econometrics, 142(1): 28–49, January 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001212.

Bruggemann:2006:RAT

[BLS06] Ralf Brüggemann, Helmut Lütkepohl, and Pentti Saikkonen. Residual autocorrelation testing for vector error correction models. Journal of Econometrics, 134(2):579–604, October 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001569.

Bollerslev:2008:RJD

[BLT08] Tim Bollerslev, Tzuo Hann Law, and George Tauchen. Risk, jumps, and diversification. Journal of Econometrics, 144(1):234– 256, May 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000080.

Bauwens:2004:RAB

[BLvD04] Luc Bauwens, Michel Lubrano, and Herman K. van Dijk. Recent advances in Bayesian econometrics. Journal of Econometrics, 123(2):197–199, December 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603002872.

Baker:2000:DDN

[BM00a] Michael Baker and Angelo Melino. Duration dependence and nonparametric heterogeneity: A Monte Carlo study. Journal of Econometrics, 96(2):357–393, June 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000640.

Banerjee:2000:SUT

[BM00b] Anurag N. Banerjee and Jan R. Magnus. On the sensitivity of the usual t- and F -tests to covariance misspecification. Journal of Econometrics, 95(1):157–176, March 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000342.

Bontemps:2005:TNG

[BM05] Christian Bontemps and Nour Meddahi. Testing normality: a GMM approach. Journal of Econometrics, 124(1):149–186, January 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604000296.

Beltratti:2006:BPM

[BM06] A. Beltratti and C. Morana. Breaks and persistency: macroeconomic causes of stock market volatility. Journal of Econometrics, 131(1–2):151–177, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605000096.

Banerjee:2009:SBR

[BMM09] Moulinath Banerjee, Debasri Mukherjee, and Santosh Mishra. Semiparametric binary regression models under shape constraints with an application to Indian schooling data. Journal of Econometrics, 149(2):101–117, April 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002091.

Basmann:2007:PAT

[BMS07] Robert L. Basmann, Michael McAleer, and Daniel Slottje. Patent activity and technical change. Journal of Econometrics, 139(2):355–375, August 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606002193.

Bai:2001:CTC

[BN01] Jushan Bai and Serena Ng. A consistent test for conditional symmetry in time series models. Journal of Econometrics, 103(1–2): 225–258, July 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000446.

Bandi:2003:FEJ

[BN03] Federico M. Bandi and Thong H. Nguyen. On the functional estimation of jump-diffusion models. Journal of Econometrics, 116(1–2):293–328, September 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603001106.

Bai:2006:ELO

[BN06a] Jushan Bai and Serena Ng. Evaluating latent and observed factors in macroeconomics and finance. Journal of Econometrics, 131(1–2):507–537, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000205.

Boivin:2006:MDA

[BN06b] Jean Boivin and Serena Ng. Are more data always better for factor analysis? Journal of Econometrics, 132(1):169–194, May 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760500045X.

Bai:2008:FET

[BN08] Jushan Bai and Serena Ng. Forecasting economic time series using targeted predictors. Journal of Econometrics, 146(2):304– 317, October 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001085.

Barndorff-Nielsen:2006:IJR

[BNS06] Ole E. Barndorff-Nielsen and Neil Shephard. Impact of jumps on returns and realised variances: econometric analysis of time-deformed Lévy processes. Journal of Econometrics, 131 (1–2):217–252, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605000114.

Bollerslev:2001:FEP

[Bol01] Tim Bollerslev. Financial econometrics: Past developments and future challenges. Journal of Econometrics, 100(1):41–51, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760000052X.

Bondarenko:2003:ERN

[Bon03] Oleg Bondarenko. Estimation of risk-neutral densities using positive convolution approximation. Journal of Econometrics, 116(1–2):85–112, September 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603001040.

Bowsher:2007:MSM

[Bow07] Clive G. Bowsher. Modelling security market events in continuous time: Intensity based, multivariate point process models. Journal of Econometrics, 141(2):876–912, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760600251X.

Brendstrup:2006:IES

[BP06] Bjarne Brendstrup and Harry J. Paarsch. Identification and estimation in sequential, asymmetric, English auctions. Journal of Econometrics, 134(1):69–94, September 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001387.

Bandi:2007:SAP

[BP07a] Federico M. Bandi and Peter C. B. Phillips. A simple approach to the parametric estimation of potentially nonstationary diffusions. Journal of Econometrics, 137(2):354–395, April 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000492.

Blundell:2007:CRQ

[BP07b] Richard Blundell and James L. Powell. Censored regression quantiles with endogenous regressors. Journal of Econometrics, 141(1):65–83, November 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407607000085.

Brendstrup:2007:SIE

[BP07c] Bjarne Brendstrup and Harry J. Paarsch. Semiparametric identification and estimation in multi-object, English auctions. Journal of Econometrics, 141(1):84–108, November 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000097.

Bandi:2008:LRR

[BP08] Federico M. Bandi and Benoît Perron. Long-run risk-return tradeoffs. Journal of Econometrics, 143(2):349–374, April 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607002291.

Blair:2001:FSV

[BPT01] Bevan J. Blair, Ser-Huang Poon, and Stephen J. Taylor. Forecasting S&P 100 volatility: the incremental information content of implied volatilities and high-frequency index returns. Journal of Econometrics, 105(1):5–26, November 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000689.

Baltagi:2005:SBT

[BR05a] Badi H. Baltagi and Daniel P. Rich. Skill-biased technical change in US manufacturing: a general index approach. Journal of Econometrics, 126(2):549–570, June 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760400123X.

Bijwaard:2005:CSC

[BR05b] Govert E. Bijwaard and Geert Ridder. Correcting for selective compliance in a re-employment bonus experiment. Journal of Econometrics, 125(1–2):77–111, March 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000752.

Battistin:2008:IEN

[BR08] Erich Battistin and Enrico Rettore. Ineligibles and eligible non-participants as a double comparison group in regressiondiscontinuity designs. Journal of Econometrics, 142(2):715–730, February 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001145.

Bonhomme:2009:CNI

[BR09] Stéphane Bonhomme and Jean-Marc Robin. Consistent noisy independent component analysis. Journal of Econometrics, 149(1): 12–25, April 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760900030X.

Breitung:2002:NTU

[Bre02] Jörg Breitung. Nonparametric tests for unit roots and cointegration. Journal of Econometrics, 108(2):343–363, June 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0304407601001397. See corrigendum [BT03a].

Brendstrup:2007:NPE

[Bre07] Bjarne Brendstrup. Non-parametric estimation of sequential English auctions. Journal of Econometrics, 141(2):460–481, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606002028.

Bai:2003:KGS

[BRT03] Xuezheng Bai, Jeffrey R. Russell, and George C. Tiao. Kurtosis of GARCH and stochastic volatility models with non-normal innovations. Journal of Econometrics, 114(2):349–360, June 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603000885.

Bandi:2008:RVF

[BRY08] Federico M. Bandi, Jeffrey R. Russell, and Chen Yang. Realized volatility forecasting and option pricing. Journal of Econometrics, 147(1):34–46, November 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760800122X.

Bierens:2000:ECC

[BS00] Herman J. Bierens and Norman R. Swanson. The econometric consequences of the ceteris paribus condition in economic theory. Journal of Econometrics, 95(2):223–253, April 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769900038X.

Bams:2003:DER

[BS03] Dennis Bams and Peter C. Schotman. Direct estimation of the risk neutral factor dynamics of Gaussian term structure models. Journal of Econometrics, 117(1):179–206, November 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001222.

Black:2004:HRE

[BS04] Dan A. Black and Jeffrey A. Smith. How robust is the evidence on the effects of college quality? Evidence from matching. Journal of Econometrics, 121(1–2):99–124, July 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002562.

Bhardwaj:2006:EIU

[BS06] Geetesh Bhardwaj and Norman R. Swanson. An empirical investigation of the usefulness of ARFIMA models for predicting macroeconomic and financial time series. Journal of Econometrics, 131(1–2):539–578, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000217.

Barnett:2008:CPD

[BS08] William A. Barnett and Apostolos Serletis. Consumer preferences and demand systems. Journal of Econometrics, 147(2):210–224, December 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001371.

Bera:2001:TEC

[BSEY01] Anil K. Bera, Walter Sosa-Escudero, and Mann Yoon. Tests for the error component model in the presence of local misspecification. Journal of Econometrics, 101(1):1–23, March 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000713.

Baltagi:2001:UNE

[BSJ01] Badi H. Baltagi, Seuck Heun Song, and Byoung Cheol Jung. The unbalanced nested error component regression model. Journal of Econometrics, 101(2):357–381, April 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000890.

Baltagi:2007:TSC

[BSJK07] Badi H. Baltagi, Seuck Heun Song, Byoung Cheol Jung, and Won Koh. Testing for serial correlation, spatial autocorrelation and random effects using panel data. Journal of Econometrics, 140(1):5–51, September 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002223.

Baltagi:2003:TPD

[BSK03] Badi H. Baltagi, Seuck Heun Song, and Won Koh. Testing panel data regression models with spatial error correlation.

Journal of Econometrics, 117(1):123–150, November 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001209.

Boero:2004:DPC

[BSW04] Gianna Boero, Jeremy Smith, and Kenneth F. Wallis. Decompositions of Pearson’s chi-squared test. Journal of Econometrics, 123(1):189–193, November 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407604000235.

Burridge:2001:RBT

[BT01] Peter Burridge and A. M. Robert Taylor. On regression-based tests for seasonal unit roots in the presence of periodic heteroscedasticity. Journal of Econometrics, 104(1):91–117, August 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601000604.

Breitung:2003:CNT

[BT03a] Jörg Breitung and A. M. Robert Taylor. Corrigendum to “Nonparametric tests for unit roots and cointegration” [J. Econom. 108 (2002) 343–363]. Journal of Econometrics, 117(2):401–404, December 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603002173. See [Bre02].

Busetti:2003:TAS

[BT03b] Fabio Busetti and A. M. Robert Taylor. Testing against stochastic trend and seasonality in the presence of unattended breaks and unit roots. Journal of Econometrics, 117(1):21–53, November 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603001179.

Burridge:2004:BHS

[BT04a] Peter Burridge and A. M. Robert Taylor. Bootstrapping the HEGY seasonal unit root tests. Journal of Econometrics, 123(1):67–87, November 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407603002835.

Busetti:2004:TSA

[BT04b] Fabio Busetti and A. M. Robert Taylor. Tests of stationarity against a change in persistence. Journal of Econometrics, 123(1):33–66, November 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407603002823.

Banerjee:2005:MSB

[BU05] Anindya Banerjee and Giovanni Urga. Modelling structural breaks, long memory and stock market volatility: an overview. Journal of Econometrics, 129(1–2):1–34, November 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001630.

Bao:2007:FSP

[BU07a] Yong Bao and Aman Ullah. Finite sample properties of maximum likelihood estimator in spatial models. Journal of Econometrics, 137(2):396–413, April 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000509.

Bao:2007:SOB

[BU07b] Yong Bao and Aman Ullah. The second-order bias and mean squared error of estimators in time-series models. Journal of Econometrics, 140(2):650–669, October 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001503.

Butler:2000:ERM

[But00] J. S. Butler. Efficiency results of MLE and GMM estimation with sampling weights. Journal of Econometrics, 96(1):25–37, May 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000494.

Bauwens:2004:SCD

[BV04] Luc Bauwens and David Veredas. The stochastic conditional duration model: a latent variable model for the analysis of financial durations. Journal of Econometrics, 119(2):381–412, April 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760300201X.

Bhalotra:2008:BSF

[BvS08] Sonia Bhalotra and Arthur van Soest. Birth-spacing, fertility and neonatal mortality in India: Dynamics, frailty, and fecundity. Journal of Econometrics, 143(2):274–290, April 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607002138.

Bollerslev:2000:SEL

[BW00] Tim Bollerslev and Jonathan H. Wright. Semiparametric estimation of long-memory volatility dependencies: The role of highfrequency data. Journal of Econometrics, 98(1):81–106, September 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000792.

Bauer:2002:ECS

[BW02] Dietmar Bauer and Martin Wagner. Estimating cointegrated systems using subspace algorithms. Journal of Econometrics, 111(1):47–84, November 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602001197.

Bollerslev:2002:ESV

[BZ02] Tim Bollerslev and Hao Zhou. Estimating stochastic volatility diffusion using conditional moments of integrated volatility. Journal of Econometrics, 109(1):33–65, July 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0304407601001415. See corrigendum [BZ04].

Bollerslev:2004:CES

[BZ04] Tim Bollerslev and Hao Zhou. Corrigendum to “Estimating stochastic volatility diffusion using conditional moments of integrated volatility” [J. Econom. 109 (2002) 33–65]. Journal of Econometrics, 119(1):221–222, March 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002665. See [BZ02].

Bollerslev:2006:VPS

[BZ06] Tim Bollerslev and Hao Zhou. Volatility puzzles: a simple framework for gauging return-volatility regressions. Journal of Econometrics, 131(1–2):123–150, March/April 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000084.

Cai:2007:TTV

[Cai07] Zongwu Cai. Trending time-varying coefficient time series models with serially correlated errors. Journal of Econometrics, 136(1):163–188, January 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605002058.

Caner:2007:BPS

[Can07] Mehmet Caner. Boundedly pivotal structural change tests in continuous updating GMM with strong, weak identification and completely unidentified cases. Journal of Econometrics, 137(1):28– 67, March 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000388.

Caner:2008:NSD

[Can08] Mehmet Caner. Nearly-singular design in GMM and generalized empirical likelihood estimators. Journal of Econometrics, 144(2): 511–523, June 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000444.

Cappelli:2004:WDE

[Cap04] Peter Cappelli. Why do employers pay for college? Journal of Econometrics, 121(1–2):213–241, July 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002641.

Carrasco:2002:MSC

[Car02] Marine Carrasco. Misspecified structural change, threshold, and Markov-switching models. Journal of Econometrics, 109(2):239– 273, August 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602001124.

Carro:2007:EDP

[Car07] Jesus M. Carro. Estimating dynamic panel data discrete choice models with fixed effects. Journal of Econometrics, 140(2):503– 528, October 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606001461.

Chen:2001:NSS

[CC01] Xiaoheng Chen and Timothy G. Conley. A new semiparametric spatial model for panel time series. Journal of Econometrics, 105(1):59–83, November 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000707.

Canova:2004:FTP

[CC04] Fabio Canova and Matteo Ciccarelli. Forecasting and turning point predictions in a Bayesian panel VAR model. Journal of Econometrics, 120(2):327–359, June 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002161.

Chen:2007:SOP

[CC07] Song Xi Chen and Hengjian Cui. On the second-order properties of empirical likelihood with moment restrictions. Journal of Econometrics, 141(2):492–516, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002041.

Carrasco:2007:EEG

[CCFG07] Marine Carrasco, Mikhail Chernov, Jean-Pierre Florens, and Eric Ghysels. Efficient estimation of general dynamic models with a continuum of moment conditions. Journal of Econometrics, 140(2):529–573, October 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606001473.

Chen:2000:TTR

[CCK00] Yi-Ting Chen, Ray Y. Chou, and Chung-Ming Kuan. Testing time reversibility without moment restrictions. Journal of Econometrics, 95(1):199–218, March 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000366.

Capps:2003:SIC

[CCL03] Oral Capps, Jr., Jeffrey Church, and H. Alan Love. Specification issues and confidence intervals in unilateral price effects analysis. Journal of Econometrics, 113(1):3–31, March 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001641.

Chen:2005:NES

[CCZ05] Gongmeng Chen, Yoon K. Choi, and Yong Zhou. Nonparametric estimation of structural change points in volatility models for time series. Journal of Econometrics, 126(1):79–114, May 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604000417.

Chen:2008:DCR

[CCZ08] Gongmeng Chen, Yoon K. Choi, and Yong Zhou. Detections of changes in return by a wavelet smoother with conditional heteroscedastic volatility. Journal of Econometrics, 143(2):227–262, April 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607002114.

Chen:2006:EMSa

[CD06] Willa W. Chen and Rohit S. Deo. Estimation of mis-specified long memory models. Journal of Econometrics, 134(1):257–281, September 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001454.

Cogley:2008:IJE

[CDN08] Timothy Cogley, Steven N. Durlauf, and James M. Nason. Introduction: Journal of Econometrics special issue honoring the research contributions of Charles R. Nelson. Journal of Econometrics, 146(2):199–201, October 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001012.

Chan:2007:IEV

[CDPX07] Ngai Hang Chan, Shi-Jie Deng, Liang Peng, and Zhendong Xia. Interval estimation of value-at-risk based on GARCH models with heavy-tailed innovations. Journal of Econometrics, 137(2):556– 576, April 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000686.

Chesher:2002:DRM

[CDS02] Andrew Chesher, Montezuma Dumangane, and Richard J. Smith. Duration response measurement error. Journal of Econometrics, 111(2):169–194, December 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001033.

Corradi:2009:PDE

[CDS09] Valentina Corradi, Walter Distaso, and Norman R. Swanson. Predictive density estimators for daily volatility based on the use of realized measures. Journal of Econometrics, 150(2):119–138, June 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608002133.

Cherchye:2008:NTC

[CDSV08] Laurens Cherchye, Bram De Rock, Jeroen Sabbe, and Frederic Vermeulen. Nonparametric tests of collectively rational consumption behavior: An integer programming procedure. Journal of Econometrics, 147(2):258–265, December 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001413.

Chesher:2007:EII

[CDvD07] Andrew Chesher, Geert Dhaene, and Herman van Dijk. Endogeneity, instruments and identification. Journal of Econometrics, 139 (1):1–3, July 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407606000996.

Cai:2006:FCI

[CDXW06] Zongwu Cai, Mitali Das, Huaiyu Xiong, and Xizhi Wu. Functional coefficient instrumental variables models. Journal of Econometrics, 133(1):207–241, July 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000898.

Chen:2008:DSE

[CES08] Jiawei Chen, Susanna Esteban, and Matthew Shum. Demand and supply estimation biases due to omission of durability. Journal of Econometrics, 147(2):247–257, December 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001401.

Calvet:2001:FMV

[CF01] Laurent Calvet and Adlai Fisher. Forecasting multifractal volatility. Journal of Econometrics, 105(1):27–58, November 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000690.

Chen:2006:EMSb

[CF06a] Xiaohong Chen and Yanqin Fan. Estimation and model selection of semiparametric copula-based multivariate dynamic models under copula misspecification. Journal of Econometrics, 135(1–2):125–154, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001776.

Chen:2006:ECB

[CF06b] Xiaohong Chen and Yanqin Fan. Estimation of copula-based semiparametric time series models. Journal of Econometrics, 130(2):307–335, February 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605000783.

Cowell:2007:IDI

[CF07] Frank A. Cowell and Emmanuel Flachaire. Income distribution and inequality measurement: The problem of extreme values. Journal of Econometrics, 141(2):1044–1072, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000036.

Carriero:2006:FFM

[CFK06] Andrea Carriero, Carlo A. Favero, and Iryna Kaminska. Financial factors, macroeconomic information and the expectations theory of the term structure of interest rates. Journal of Econometrics, 131(1–2):339–358, March/April 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000151.

Consolo:2009:SID

[CFP09] Agostino Consolo, Carlo A. Favero, and Alessia Paccagnini. On the statistical identification of DSGE models. Journal of Econometrics, 150(1):99–115, May 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000487.

Cazals:2002:NFE

[CFS02] Catherine Cazals, Jean-Pierre Florens, and Léopold Simar. Nonparametric frontier estimation: a robust approach. Journal of Econometrics, 106(1):1–25, January 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760100080X.

Calvet:2006:VCM

[CFT06] Laurent E. Calvet, Adlai J. Fisher, and Samuel B. Thompson. Volatility comovement: a multifrequency approach. Journal of Econometrics, 131(1–2):179–215, March/April 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000102.

Coppejans:2002:CVS

[CG02] Mark Coppejans and A. Ronald Gallant. Cross-validated SNP density estimates. Journal of Econometrics, 110(1):27–65, September 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602001215.

Chen:2007:AEL

[CG07] Song Xi Chen and Jiti Gao. An adaptive empirical likelihood test for parametric time series regression models. Journal of Econometrics, 141(2):950–972, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002533.

Casas:2008:EEL

[CG08] Isabel Casas and Jiti Gao. Econometric estimation in longrange dependent volatility models: Theory and practice. Journal of Econometrics, 147(1):72–83, November 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001255.

Chernov:2003:AMS

[CGGT03] Mikhail Chernov, A. Ronald Gallant, Eric Ghysels, and George Tauchen. Alternative models for stock price dynamics. Journal of Econometrics, 116(1–2):225–257, September 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001088.

Clement:2000:ESR

[CGM00] E. Clement, C. Gourieroux, and A. Monfort. Econometric specification of the risk neutral valuation model. Journal of Econometrics, 94(1–2):117–143, January 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000196.

Cao:2008:GFT

[CGM08] Ricardo Cao and Wenceslao González-Manteiga. Goodness-of-fit tests for conditional models under censoring and truncation. Journal of Econometrics, 143(1):166–190, March 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001662.

Chernew:2008:LVI

[CGS08] Michael Chernew, Gautam Gowrisankaran, and Dennis P. Scanlon. Learning and the value of information: Evidence from health plan report cards. Journal of Econometrics, 144(1):156–174, May 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608000031.

Chib:2000:BAC

[CH00] Siddhartha Chib and Barton H. Hamilton. Bayesian analysis of cross-section and clustered data treatment models. Journal of Econometrics, 97(1):25–50, July 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000652.

Chib:2002:SBA

[CH02] Siddhartha Chib and Barton H. Hamilton. Semiparametric Bayes analysis of longitudinal data treatment models. Journal of Econometrics, 110(1):67–89, September 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001227.

Chen:2003:EFC

[CH03a] Willa W. Chen and Clifford M. Hurvich. Estimating fractional cointegration in the presence of polynomial trends. Journal of Econometrics, 117(1):95–121, November 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001192.

Chernozhukov:2003:MAC

[CH03b] Victor Chernozhukov and Han Hong. An MCMC approach to classical estimation. Journal of Econometrics, 115(2):293–346, August 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603001003.

Chernozhukov:2006:IQR

[CH06] Victor Chernozhukov and Christian Hansen. Instrumental quantile regression inference for structural and treatment effect models. Journal of Econometrics, 132(2):491–525, June 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000643.

Chernozhukov:2008:IVQ

[CH08] Victor Chernozhukov and Christian Hansen. Instrumental variable quantile regression: A robust inference approach. Journal of Econometrics, 142(1):379–398, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001455.

Card:2009:DEE

[CH09] David Card and Dean R. Hyslop. The dynamic effects of an earnings subsidy for long-term welfare recipients: Evidence from the self sufficiency project applicant experiment. Journal of Econometrics, 153(1):1–20, November 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000748.

Chamberlain:2000:EDT

[Cha00] Gary Chamberlain. Econometrics and decision theory. Journal of Econometrics, 95(2):255–283, April 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000391.

Chang:2002:NIU

[Cha02] Yoosoon Chang. Nonlinear IV unit root tests in panels with cross-sectional dependency. Journal of Econometrics, 110(2):261– 292, October 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602000957.

Chambers:2004:TUR

[Cha04a] Marcus J. Chambers. Testing for unit roots with flow data and varying sampling frequency. Journal of Econometrics, 119(1):1– 18, March 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603001520. See corrigendum

[Cha08].

Chang:2004:BUR

[Cha04b] Yoosoon Chang. Bootstrap unit root tests in panels with crosssectional dependency. Journal of Econometrics, 120(2):263–293, June 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603002148.

Chambers:2008:CTU

[Cha08] Marcus J. Chambers. Corrigendum to: “Testing for unit roots with flow data and varying sampling frequency” [J. Econom. 119(1) (2004) 1–18]. Journal of Econometrics, 144(2):524–525, June 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608000225. See [Cha04a].

Chen:2000:EEB

[Che00a] Songnian Chen. Efficient estimation of binary choice models under symmetry. Journal of Econometrics, 96(1):183–199, May 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000597.

Chen:2000:REL

[Che00b] Songnian Chen. Rank estimation of a location parameter in the binary choice model. Journal of Econometrics, 98(2):317– 334, October 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760000021X.

Chernov:2003:ERE

[Che03] Mikhail Chernov. Empirical reverse engineering of the pricing kernel. Journal of Econometrics, 116(1–2):329–364, September 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603001118.

Chesher:2007:IV

[Che07] Andrew Chesher. Instrumental values. Journal of Econometrics, 139(1):15–34, July 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407606001011.

Chesher:2009:EHE

[Che09] Andrew Chesher. Excess heterogeneity, endogeneity and index restrictions. Journal of Econometrics, 152(1):37–45, September 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609000530.

Chib:2007:ATR

[Chi07] Siddhartha Chib. Analysis of treatment response data without the joint distribution of potential outcomes. Journal of Econometrics, 140(2):401–412, October 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001412.

Christoffersen:2006:OVC

[CHJ06] Peter Christoffersen, Steve Heston, and Kris Jacobs. Option valuation with conditional skewness. Journal of Econometrics, 131(1–2):253–284, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605000126.

Chernozhukov:2009:FSI

[CHJ09] Victor Chernozhukov, Christian Hansen, and Michael Jansson. Finite sample inference for quantile regression models. Journal of Econometrics, 152(2):93–103, October 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000220.

Conley:2008:SPB

[CHMR08] Timothy G. Conley, Christian B. Hansen, Robert E. McCulloch, and Peter E. Rossi. A semi-parametric Bayesian approach to the instrumental variable problem. Journal of Econometrics, 144(1): 276–305, May 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000171.

Chong:2000:EDP

[Cho00] Terence Tai-Leung Chong. Estimating the differencing parameter via the partial autocorrelation function. Journal of Econometrics, 97(2):365–381, August 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407699000767.

Choi:2002:IVE

[Cho02] In Choi. Instrumental variables estimation of a nearly nonstationary, heterogeneous error component model. Journal of Econometrics, 109(1):1–32, July 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407601001403.

Choi:2005:SVA

[Cho05] In Choi. Subsampling vector autoregressive tests of linear constraints. Journal of Econometrics, 124(1):55–89, January 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604000260.

Choi:2008:RES

[CHO08] Chi-Young Choi, Ling Hu, and Masao Ogaki. Robust estimation for structural spurious regressions and a hausman-type cointegration test. Journal of Econometrics, 142(1):327–351, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607001431.

Cubadda:2009:SCM

[CHP09] Gianluca Cubadda, Alain Hecq, and Franz C. Palm. Studying co-movements in large multivariate data prior to multivariate modelling. Journal of Econometrics, 148(1):25–35, January 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608001188.

Chen:2007:NLR

[CHS07] Xiaohong Chen, Han Hong, and Matthew Shum. Nonparametric likelihood ratio model selection tests between parametric likelihood and moment condition models. Journal of Econometrics, 141(1):109–140, November 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000103.

Corradi:2008:BRQ

[CI08] Valentina Corradi and Emma M. Iglesias. Bootstrap refinements for QML estimators of the GARCH(1,1) parameters. Journal of Econometrics, 144(2):500–510, June 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000432.

Chernozhukov:2007:IVE

[CIN07] Victor Chernozhukov, Guido W. Imbens, and Whitney K. Newey. Instrumental variable estimation of nonseparable models. Journal of Econometrics, 139(1):4–14, July 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/S030440760600100X.

Chib:2007:MCE

[CJ07] Siddhartha Chib and Liana Jacobi. Modeling and calculating the effect of treatment at baseline from panel outcomes. Journal of Econometrics, 140(2):781–801, October 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001552.

Chib:2008:ATR

[CJ08] Siddhartha Chib and Liana Jacobi. Analysis of treatment response data from eligibility designs. Journal of Econometrics, 144(2): 465–478, June 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000407.

Chen:2000:ECR

[CK00] Songnian Chen and Shakeeb Khan. Estimating censored regression models in the presence of nonparametric multiplicative heteroskedasticity. Journal of Econometrics, 98(2):283–316, October 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000208.

Chen:2002:PTS

[CK02] Yi-Ting Chen and Chung-Ming Kuan. The pseudo-true score encompassing test for non-nested hypotheses. Journal of Econometrics, 106(2):271–295, February 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0304407601001014. See erratum [CK07].

Chen:2003:RCE

[CK03] Songnian Chen and Shakeeb Khan. Rates of convergence for estimating regression coefficients in heteroskedastic discrete response models. Journal of Econometrics, 117(2):245–278, December 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603001489.

Chen:2007:CPT

[CK07] Yi-Ting Chen and Chung-Ming Kuan. Corrigendum to “The pseudo-true score encompassing test for non-nested hypotheses”: [Journal of Econometrics 106, 271–295]. Journal of Econometrics, 141(2):1412–1417, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000024. See [CK02].

Connor:2006:CSC

[CKL06] Gregory Connor, Robert A. Korajczyk, and Oliver Linton. The common and specific components of dynamic volatility. Journal of Econometrics, 132(1):231–255, May 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000473.

Chintagunta:2001:PDA

[CKP01] Pradeep Chintagunta, Ekaterini Kyriazidou, and Josef Perktold. Panel data analysis of household brand choices. Journal of Econometrics, 103(1–2):111–153, July 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000410.

Chay:2000:CRW

[CL00] Kenneth Y. Chay and David S. Lee. Changes in relative wages in the 1980s returns to observed and unobserved skills and blackwhite wage differentials. Journal of Econometrics, 99(1):1–38, November 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000294.

Carneiro:2009:EDP

[CL09] Pedro Carneiro and Sokbae Lee. Estimating distributions of potential outcomes using local instrumental variables with an application to changes in college enrollment and wage inequality. Journal of Econometrics, 149(2):191–208, April 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000281.

Clarke:2007:IOB

[Cla07] Bertrand Clarke. Information optimality and Bayesian modelling. Journal of Econometrics, 138(2):405–429, June 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000765.

Crawford:2003:EHD

[CLP03] Ian Crawford, François Laisney, and Ian Preston. Estimation of household demand systems with theoretically compatible Engel curves and unit value specifications. Journal of Econometrics, 114(2):221–241, June 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603000836.

Cai:2009:FCM

[CLP09] Zongwu Cai, Qi Li, and Joon Y. Park. Functional-coefficient models for nonstationary time series data. Journal of Econometrics, 148(2):101–113, February 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760800167X.

Clark:2001:TEF

[CM01] Todd E. Clark and Michael W. McCracken. Tests of equal forecast accuracy and encompassing for nested models. Journal of Econometrics, 105(1):85–110, November 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000719.

Clark:2005:PTP

[CM05] Todd E. Clark and Michael W. McCracken. The power of tests of predictive ability in the presence of structural breaks. Journal of Econometrics, 124(1):1–31, January 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000247.

Chambers:2007:FDE

[CM07a] Marcus J. Chambers and J. Roderick McCrorie. Frequency domain estimation of temporally aggregated Gaussian cointegrated systems. Journal of Econometrics, 136(1):1–29, January 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606000364.

Conley:2007:SCR

[CM07b] Timothy G. Conley and Francesca Molinari. Spatial correlation robust inference with errors in location or distance. Journal of Econometrics, 140(1):76–96, September 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002247.

Chang:2009:ECS

[CMP09] Yoosoon Chang, J. Isaac Miller, and Joon Y. Park. Extracting a common stochastic trend: Theory with some applications. Journal of Econometrics, 150(2):231–247, June 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002212.

Charlier:2000:ECR

[CMvS00] Erwin Charlier, Bertrand Melenberg, and Arthur van Soest. Estimation of a censored regression panel data model using conditional moment restrictions efficiently. Journal of Econometrics, 95(1): 25–56, March 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407699000287.

Charlier:2001:AHE

[CMvS01] Erwin Charlier, Bertrand Melenberg, and Arthur van Soest. An analysis of housing expenditure using semiparametric models and panel data. Journal of Econometrics, 101(1):71–107, March 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000749.

Christensen:2006:ANN

[CN06] Bent Jesper Christensen and Morten Ørregaard Nielsen. Asymptotic normality of narrow-band least squares in the stationary fractional cointegration model and volatility forecasting. Journal of Econometrics, 133(1):343–371, July 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760500093X.

Chib:2002:MCM

[CNS02] Siddhartha Chib, Federico Nardari, and Neil Shephard. Markov chain Monte Carlo methods for stochastic volatility models. Journal of Econometrics, 108(2):281–316, June 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001373.

Chib:2006:AHD

[CNS06] Siddhartha Chib, Federico Nardari, and Neil Shephard. Analysis of high dimensional multivariate stochastic volatility models. Journal of Econometrics, 134(2):341–371, October 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001478.

Cook:2008:WLA

[Coo08] Thomas D. Cook. “waiting for life to arrive”: A history of the regression-discontinuity design in psychology, statistics and economics. Journal of Econometrics, 142(2):636–654, February 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607001108.

Coppejans:2001:EBR

[Cop01] Mark Coppejans. Estimation of the binary response model using a mixture of distributions estimator (MOD). Journal of Econometrics, 102(2):231–269, June 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000549.

Coppejans:2003:ENE

[Cop03] Mark Coppejans. Effective nonparametric estimation in the case of severely discretized data. Journal of Econometrics, 117(2):331– 367, December 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603001519.

Coppejans:2004:KRF

[Cop04] Mark Coppejans. On Kolmogorov’s representation of functions of several variables by functions of one variable. Journal of Econometrics, 123(1):1–31, November 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/S030440760300280X.

Coppejans:2007:EEO

[Cop07] Mark Coppejans. On efficient estimation of the ordered response model. Journal of Econometrics, 137(2):577–614, April 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000698.

Corradi:2000:RCT

[Cor00] Valentina Corradi. Reconsidering the continuous time limit of the GARCH(1, 1) process. Journal of Econometrics, 96(1):145– 153, May 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407699000536.

Costa:2003:UML

[Cos03] Dora L. Costa. Understanding mid-life and older age mortality declines: evidence from union Army veterans. Journal of Econometrics, 112(1):175–192, January 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001598.

Chao:2002:JPA

[CP02] John C. Chao and Peter C. B. Phillips. Jeffreys prior analysis of the simultaneous equations model in the case with n +1 endogenous variables. Journal of Econometrics, 111(2):251–283, December 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602001069.

Chang:2003:IMI

[CP03] Yoosoon Chang and Joon Y. Park. Index models with integrated time series. Journal of Econometrics, 114(1):73–106, May 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602002208.

Chiuso:2004:AVS

[CP04a] Alessandro Chiuso and Giorgio Picci. The asymptotic variance of subspace estimates. Journal of Econometrics, 118(1–2):257– 291, January 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760300143X.

Chopin:2004:BIS

[CP04b] Nicolas Chopin and Florian Pelgrin. Bayesian inference and state number determination for hidden Markov models: an application to the information content of the yield curve about inflation. Journal of Econometrics, 123(2):327–344, December 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002896.

Christensen:2007:RRB

[CP07a] Kim Christensen and Mark Podolskij. Realized range-based estimation of integrated variance. Journal of Econometrics, 141(2):323–349, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606001989.

Chung:2007:NNH

[CP07b] Heetaik Chung and Joon Y. Park. Nonstationary nonlinear heteroskedasticity in regression. Journal of Econometrics, 137(1): 230–259, March 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000467.

Chen:2009:EES

[CP09] Xiaohong Chen and Demian Pouzo. Efficient estimation of semiparametric conditional moment models with possibly nonsmooth residuals. Journal of Econometrics, 152(1):46–60, September 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609000529.

Chang:2006:BCR

[CPS06] Yoosoon Chang, Joon Y. Park, and Kevin Song. Bootstrapping cointegrating regressions. Journal of Econometrics, 133(2):703– 739, August 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001314.

Card:2005:HIE

[CR05] David Card and Philip K. Robins. How important are “entry effects” in financial incentive programs for welfare recipients? Experimental evidence from the Self-Sufficiency Project. Journal of Econometrics, 125(1–2):113–139, March 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000764.

Cho:2008:EUP

[CR08] Sungjin Cho and John Rust. Is econometrics useful for private policy making? A case study of replacement policy at an auto rental company. Journal of Econometrics, 145(1–2):243–257, July 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608000602.

Croux:2004:DFM

[CRW04] Christophe Croux, Eric Renault, and Bas Werker. Dynamic factor models. Journal of Econometrics, 119(2):223–230, April 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603001957.

Croushore:2001:RTD

[CS01] Dean Croushore and Tom Stark. A real-time data set for macroeconomists. Journal of Econometrics, 105(1):111–130, November 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601000720.

Corradi:2002:CTN

[CS02] Valentina Corradi and Norman R. Swanson. A consistent test for nonlinear out of sample predictive accuracy. Journal of Econometrics, 110(2):353–381, October 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602000994.

Coscelli:2004:EML

[CS04] Andrea Coscelli and Matthew Shum. An empirical model of learning and patient spillovers in new drug entry. Journal of Econometrics, 122(2):213–246, October 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002744.

Corradi:2005:BST

[CS05] Valentina Corradi and Norman R. Swanson. Bootstrap specification tests for diffusion processes. Journal of Econometrics, 124(1):117–148, January 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604000284.

Corradi:2006:BCD

[CS06a] Valentina Corradi and Norman R. Swanson. Bootstrap conditional distribution tests in the presence of dynamic misspecification. Journal of Econometrics, 133(2):779–806, August 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001338.

Corradi:2006:EDT

[CS06b] Valentina Corradi and Norman R. Swanson. The effect of data transformation on common cycle, cointegration, and unit root tests: Monte Carlo results and a simple test. Journal of Econometrics, 132(1):195–229, May 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000461.

Corradi:2006:PDC

[CS06c] Valentina Corradi and Norman R. Swanson. Predictive density and conditional confidence interval accuracy tests. Journal of Econometrics, 135(1–2):187–228, November/December 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001764.

Chao:2007:AAB

[CS07a] John Chao and Norman R. Swanson. Alternative approximations of the bias and MSE of the IV estimator under weak identification with an application to bias correction. Journal of Econometrics, 137(2):515–555, April 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000674.

Corradi:2007:EDS

[CS07b] Valentina Corradi and Norman R. Swanson. Evaluation of dynamic stochastic general equilibrium models based on distributional comparison of simulated and historical data. Journal of Econometrics, 136(2):699–723, February 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002290.

Corradi:2001:PAC

[CSO01] Valentina Corradi, Norman R. Swanson, and Claudia Olivetti. Predictive ability with cointegrated variables. Journal of Econometrics, 104(2):315–358, September 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000860.

Corradi:2000:TSE

[CSW00] Valentina Corradi, Norman R. Swanson, and Halbert White. Testing for stationarity-ergodicity and for comovements between nonlinear discrete time Markov processes. Journal of Econometrics, 96(1):39–73, May 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407699000500.

Cavaliere:2007:TUR

[CT07a] Giuseppe Cavaliere and A. M. Robert Taylor. Testing for unit roots in time series models with non-stationary volatility. Journal of Econometrics, 140(2):919–947, October 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001618.

Conley:2007:EDL

[CT07b] Timothy G. Conley and Giorgio Topa. Estimating dynamic local interactions models. Journal of Econometrics, 140(1):282–303, September 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606002326.

Cavaliere:2008:TCP

[CT08] Giuseppe Cavaliere and A. M. Robert Taylor. Testing for a change in persistence in the presence of non-stationary volatility. Journal of Econometrics, 147(1):84–98, November 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001267.

Chacko:2003:SGE

[CV03] George Chacko and Luis M. Viceira. Spectral GMM estimation of continuous-time processes. Journal of Econometrics, 116(1–2):259–292, September 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760300109X.

Chen:2008:WDE

[CvdK08] Susan Chen and Wilbert van der Klaauw. The work disincentive effects of the disability insurance program in the 1990s. Journal of Econometrics, 142(2):757–784, February 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001169.

Clark:2006:USM

[CW06] Todd E. Clark and Kenneth D. West. Using out-of-sample mean squared prediction errors to test the martingale difference hypothesis. Journal of Econometrics, 135(1–2):155–186, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001648.

Chiou:2007:MID

[CW07a] Lesley Chiou and Joan L. Walker. Masking identification of discrete choice models under simulation methods. Journal of Econometrics, 141(2):683–703, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002119.

Clark:2007:ANT

[CW07b] Todd E. Clark and Kenneth D. West. Approximately normal tests for equal predictive accuracy in nested models. Journal of Econometrics, 138(1):291–311, May 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000960.

Cai:2008:NEC

[CW08] Zongwu Cai and Xian Wang. Nonparametric estimation of conditional VaR and expected shortfall. Journal of Econometrics, 147(1):120–130, November 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407608001292.

Chen:2007:EGC

[CZ07] Songnian Chen and Yahong Zhou. Estimating a generalized correlation coefficient for a generalized bivariate probit model. Journal of Econometrics, 141(2):1100–1114, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000255.

Dahlhaus:2009:LIL

[Dah09] Rainer Dahlhaus. Local inference for locally stationary time series based on the empirical spectral measure. Journal of Econometrics, 151(2):101–112, August 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760900075X.

Das:2002:SEJ

[Das02] Sanjiv R. Das. The surprise element: jumps in interest rates. Journal of Econometrics, 106(1):27–65, January 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000859.

Das:2003:ISE

[Das03a] M. Das. Identification and sequential estimation of panel data models with insufficient exclusion restrictions. Journal of Econometrics, 114(2):297–328, June 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603000861.

Dastoor:2003:ECE

[Das03b] Naorayex K. Dastoor. The equality of comparable extended families of classical-type and hausman-type statistics. Journal of Econometrics, 117(2):313–330, December 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001507.

Das:2004:SEN

[Das04] M. Das. Simple estimators for nonparametric panel data models with sample attrition. Journal of Econometrics, 120(1):159– 180, May 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603002100.

Das:2005:IVE

[Das05] M. Das. Instrumental variables estimators of nonparametric models with discrete endogenous regressors. Journal of Econometrics, 124(2):335–361, February 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000363.

Davidson:2002:CEC

[Dav02a] James Davidson. Corrigendum to “Establishing conditions for the functional central limit theorem in nonlinear and semiparametric time series processes”: [Journal of Econometrics 106(2) (2002) 243–269]. Journal of Econometrics, 110(1):103–104, September 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602001264. See [Dav02b].

Davidson:2002:ECF

[Dav02b] James Davidson. Establishing conditions for the functional central limit theorem in nonlinear and semiparametric time series processes. Journal of Econometrics, 106(2):243–269, February 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601001002. See corrigendum [Dav02a].

Davidson:2002:MFC

[Dav02c] James Davidson. A model of fractional cointegration, and tests for cointegration using the bootstrap. Journal of Econometrics, 110(2):187–212, October 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602000921.

Davis:2002:EMW

[Dav02d] Peter Davis. Estimating multi-way error components models with unbalanced data structures. Journal of Econometrics, 106(1):67– 95, January 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000872.

Davidson:2006:ABP

[Dav06a] James Davidson. Alternative bootstrap procedures for testing cointegration in fractionally integrated processes. Journal of Econometrics, 133(2):741–777, August 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001326.

Davis:2006:EQG

[Dav06b] Peter Davis. Estimation of quantity games in the presence of indivisibilities and heterogeneous firms. Journal of Econometrics, 134(1):187–214, September 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001429.

Davidson:2009:RIG

[Dav09] Russell Davidson. Reliable inference for the Gini index. Journal of Econometrics, 150(1):30–40, May 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000323.

Delgado:2007:NTC

[DE07] Miguel A. Delgado and J. Carlos Escanciano. Nonparametric tests for conditional symmetry in dynamic models. Journal of Econometrics, 141(2):652–682, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002107.

Diebold:2006:EMF

[DEF^{+}06] F. X. Diebold, R. F. Engle, C. Favero, G. M. Gallo, and F. Schorfheide. The econometrics of macroeconomics, finance, and the interface. Journal of Econometrics, 131(1–2):1–2, March/ April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605000035.

Domowitz:2001:CNT

[DEG01] Ian Domowitz and Mahmoud A. El-Gamal. A consistent nonparametric test of ergodicity for time series with applications. Journal of Econometrics, 102(2):365–398, June 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000586.

Dehejia:2005:PPS

[Deh05a] Rajeev Dehejia. Practical propensity score matching: a reply to Smith and Todd. Journal of Econometrics, 125(1–2):355– 364, March 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604000831.

Dehejia:2005:PED

[Deh05b] Rajeev H. Dehejia. Program evaluation as a decision problem. Journal of Econometrics, 125(1–2):141–173, March 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000776.

Deistler:2001:CCC

[Dei01] M. Deistler. Comments on the contributions by C. W. J. Granger and J. J. Heckman. Journal of Econometrics, 100(1):71–72, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000567. See [Gra01, Hec01].

Delgado:2008:ST

[Del08] Miguel A. Delgado. Specification testing. Journal of Econometrics, 143(1):1–4, March 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407607001716.

Delgado:2009:EIa

[Del09a] Miguel A. Delgado. Editor’s introduction. Journal of Econometrics, 151(2):99–100, August 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609000839.

Delgado:2009:EIb

[Del09b] Miguel A. Delgado. Editor’s introduction. Journal of Econometrics, 152(1):1–2, September 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760900058X.

Deo:2000:ETG

[Deo00a] Rohit S. Deo. On estimation and testing goodness of fit for mdependent stable sequences. Journal of Econometrics, 99(2):349– 372, December 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000397.

Deo:2000:STM

[Deo00b] Rohit S. Deo. Spectral tests of the martingale hypothesis under conditional heteroscedasticity. Journal of Econometrics, 99(2):291–315, December 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407600000270.

Deschamps:2000:ESS

[Des00] Philippe J. Deschamps. Exact small-sample inference in stationary, fully regular, dynamic demand models. Journal of Econometrics, 97(1):51–91, July 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407699000664.

Deschamps:2006:FPD

[Des06] Philippe J. Deschamps. A flexible prior distribution for Markov switching autoregressions with Student-t errors. Journal of Econometrics, 133(1):153–190, July 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000874.

Delgado:2002:EBT

[DF02] Miguel A. Delgado and Inmaculada Fiteni. External bootstrap tests for parameter stability. Journal of Econometrics, 109(2):275–303, August 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760200115X.

Davidson:2007:ABI

[DF07] Russell Davidson and Emmanuel Flachaire. Asymptotic and bootstrap inference for inequality and poverty measures. Journal of Econometrics, 141(1):141–166, November 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000115.

Davidson:2008:WBT

[DF08a] Russell Davidson and Emmanuel Flachaire. The wild bootstrap, tamed at last. Journal of Econometrics, 146(1):162–169, September 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608000833.

Dey:2008:HSH

[DF08b] Matthew Dey and Christopher Flinn. Household search and health insurance coverage. Journal of Econometrics, 145(1–2): 43–63, July 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000481.

Diewert:2008:ERS

[DF08c] W. Erwin Diewert and Kevin J. Fox. On the estimation of returns to scale, technical progress and monopolistic markups. Journal of Econometrics, 145(1–2):174–193, July 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000584.

Duan:2009:ESC

[DF09] Jin-Chuan Duan and Andras Fulop. Estimating the structural credit risk model when equity prices are contaminated by trading noises. Journal of Econometrics, 150(2):288–296, June 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760800225X.

Darolles:2004:KBN

[DFG04] Serge Darolles, Jean-Pierre Florens, and Christian Gouriéroux. Kernel-based nonlinear canonical analysis and time reversibility. Journal of Econometrics, 119(2):323–353, April 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001994.

Dufour:2006:DFB

[DFH06] Jean-Marie Dufour, Abdeljelil Farhat, and Marc Hallin. Distributionfree bounds for serial correlation coefficients in heteroskedastic symmetric time series. Journal of Econometrics, 130(1):123– 142, January 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605000503.

Darolles:2001:TDE

[DG01] Serge Darolles and Christian Gouriéroux. Truncated dynamics and estimation of diffusion equations. Journal of Econometrics, 102(1):1–22, May 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407600000853.

Dittmann:2002:PNT

[DG02] Ingolf Dittmann and Clive W. J. Granger. Properties of nonlinear transformations of fractionally integrated processes. Journal of Econometrics, 110(2):113–133, October 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602000891.

Dias:2007:CEB

[DG07] Ronaldo Dias and Nancy L. Garcia. Consistent estimator for basis selection based on a proxy of the Kullback–Leibler distance. Journal of Econometrics, 141(1):167–178, November 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000127.

Doiron:2008:SDY

[DG08] Denise Doiron and Tue Gørgens. State dependence in youth labor market experiences, and the evaluation of policy interventions. Journal of Econometrics, 145(1–2):81–97, July 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000523.

Dahl:2003:TNN

[DGR03] Christian M. Dahl and Gloria González-Rivera. Testing for neglected nonlinearity in regression models based on the theory of random fields. Journal of Econometrics, 114(1):141–164, May 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602002221.

Detemple:2006:APM

[DGR06] Jérôme Detemple, René Garcia, and Marcel Rindisbacher. Asymptotic properties of Monte Carlo estimators of diffusion processes. Journal of Econometrics, 134(1):1–68, September 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605001375.

Dridi:2007:IIC

[DGR07] Ramdan Dridi, Alain Guay, and Eric Renault. Indirect inference and calibration of dynamic stochastic general equilibrium models. Journal of Econometrics, 136(2):397–430, February 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002198.

DeMol:2008:FUL

[DGR08] Christine De Mol, Domenico Giannone, and Lucrezia Reichlin. Forecasting using a large number of predictors: Is Bayesian shrinkage a valid alternative to principal components? Journal of Econometrics, 146(2):318–328, October 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001103.

Delgado:2000:NIS

[DH00] Miguel A. Delgado and Javier Hidalgo. Nonparametric inference on structural breaks. Journal of Econometrics, 96(1):113– 144, May 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407699000524.

Dalla:2005:PBT

[DH05] Violetta Dalla and Javier Hidalgo. A parametric bootstrap test for cycles. Journal of Econometrics, 129(1–2):219–261, November 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001708.

DeRossi:2009:QES

[DH09] Giuliano De Rossi and Andrew Harvey. Quantiles, expectiles and splines. Journal of Econometrics, 152(2):179–185, October 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609000153.

Deo:2006:FRV

[DHL06] Rohit Deo, Clifford Hurvich, and Yi Lu. Forecasting realized volatility using a long-memory stochastic volatility model: estimation, prediction and seasonal adjustment. Journal of Econometrics, 131(1–2):29–58, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000059.

Diebold:2001:LMR

[DI01] Francis X. Diebold and Atsushi Inoue. Long memory and regime switching. Journal of Econometrics, 105(1):131–159, November 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601000732.

Diebold:2001:ERP

[Die01] Francis X. Diebold. Econometrics: Retrospect and prospect. Journal of Econometrics, 100(1):73–75, January 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000592.

Donald:2003:ELE

[DIN03] Stephen G. Donald, Guido W. Imbens, and Whitney K. Newey. Empirical likelihood estimation and consistent tests with conditional moment restrictions. Journal of Econometrics, 117(1):55– 93, November 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603001180.

Donald:2009:CIV

[DIN09] Stephen G. Donald, Guido W. Imbens, and Whitney K. Newey. Choosing instrumental variables in conditional moment restriction models. Journal of Econometrics, 152(1):28–36, September 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609000566.

Dirkmaat:2001:SPF

[Dir01] Joop Dirkmaat. Some publishing facts, figures, and observations on the occasion of volume 100, number 1 of the journal of econometrics. Journal of Econometrics, 100(1):99–112, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000701.

Distaso:2008:TUR

[Dis08] Walter Distaso. Testing for unit root processes in random coefficient autoregressive models. Journal of Econometrics, 142(1):581– 609, January 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001765.

DeJong:2000:BAD

[DIW00] David N. DeJong, Beth F. Ingram, and Charles H. Whiteman. A Bayesian approach to dynamic macroeconomics. Journal of Econometrics, 98(2):203–223, October 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000191.

deJong:2001:NEU

[dJ01] Robert M. de Jong. Nonlinear estimation using estimated cointegrating relations. Journal of Econometrics, 101(1):109–122, March 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000750.

deJong:2002:NME

[dJ02a] Robert M. de Jong. Nonlinear minimization estimators in the presence of cointegrating relations. Journal of Econometrics, 110(2):241–259, October 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602000933.

deJong:2002:NCR

[dJ02b] Robert M. de Jong. A note on “convergence rates and asymptotic normality for series estimators”: uniform convergence rates. Journal of Econometrics, 111(1):1–9, November 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001136.

Dufour:2006:FSS

[DJ06] Jean-Marie Dufour and Tarek Jouini. Finite-sample simulationbased inference in VAR models with application to Granger causality testing. Journal of Econometrics, 135(1–2):229– 254, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605001752.

deJong:2007:RVK

[dJAS07] Robert M. de Jong, Christine Amsler, and Peter Schmidt. A robust version of the KPSS test based on indicators. Journal of Econometrics, 137(2):311–333, April 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000479.

Dufour:2002:ETC

[DK02a] Jean-Marie Dufour and Lynda Khalaf. Exact tests for contemporaneous correlation of disturbances in seemingly unrelated regressions. Journal of Econometrics, 106(1):143–170, January 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601000938.

Dufour:2002:SBF

[DK02b] Jean-Marie Dufour and Lynda Khalaf. Simulation based finite and large sample tests in multivariate regressions. Journal of Econometrics, 111(2):303–322, December 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001082.

Dorfman:2005:CDP

[DK05]  Jeffrey H. Dorfman and Gary Koop.  Current develop-

ments in productivity and efficiency measurement. Journal of Econometrics, 126(2):233–240, June 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001113.

Dufour:2004:SBF

[DKBG04] Jean-Marie Dufour, Lynda Khalaf, Jean-Thomas Bernard, and Ian Genest. Simulation-based finite-sample tests for heteroskedasticity and ARCH effects. Journal of Econometrics, 122(2):317– 347, October 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603002781.

Diebold:2006:FTS

[DL06] Francis X. Diebold and Canlin Li. Forecasting the term structure of government bond yields. Journal of Econometrics, 130(2):337– 364, February 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605000795.

deLuna:2006:ESE

[dLJ06] Xavier de Luna and Per Johansson. Exogeneity in structural equation models. Journal of Econometrics, 132(2):527–543, June 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605000655.

Dhrymes:2006:EMG

[DLM06] Phoebus J. Dhrymes and Adriana Lleras-Muney. Estimation of models with grouped and ungrouped data by means of “2SLS”. Journal of Econometrics, 133(1):1–29, July 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000813.

Diebold:2008:GYC

[DLY08] Francis X. Diebold, Canlin Li, and Vivian Z. Yue. Global yield curve dynamics and interactions: a dynamic Nelson–Siegel approach. Journal of Econometrics, 146(2):351–363, October 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608001127.

Davidson:2002:BJT

[DM02] Russell Davidson and James G. MacKinnon. Bootstrap j tests of nonnested linear regression models. Journal of Econometrics, 109(1):167–193, July 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407601001464.

Danilov:2004:HIP

[DM04] Dmitry Danilov and Jan R. Magnus. On the harm that ignoring pretesting can cause. Journal of Econometrics, 122(1):27–46, September 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603002689.

Davidson:2006:PBA

[DM06] Russell Davidson and James G. MacKinnon. The power of bootstrap and asymptotic tests. Journal of Econometrics, 133(2):421– 441, August 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001211.

deMatos:2007:TMP

[dMF07] João Amaro de Matos and Marcelo Fernandes. Testing the Markov property with high frequency data. Journal of Econometrics, 141(1):44–64, November 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000073.

Driessen:2005:TAT

[DMN05] Joost Driessen, Bertrand Melenberg, and Theo Nijman. Testing affine term structure models in case of transaction costs.

Journal of Econometrics, 126(1):201–232, May 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001368.

Dominitz:2001:EIE

[Dom01] Jeff Dominitz. Estimation of income expectations models using expectations and realization data. Journal of Econometrics, 102(2): 165–195, June 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000525.

Donald:2002:SEI

[DP02] Stephen G. Donald and Harry J. Paarsch. Superconsistent estimation and inference in structural econometric models using extreme order statistics. Journal of Econometrics, 109(2):305– 340, August 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602001161.

Dufour:2006:RME

[DP06] Jean-Marie Dufour and Benoît Perron. Resampling methods in econometrics. Journal of Econometrics, 133(2):411–419, August 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760500120X.

Deng:2008:NLP

[DP08a] Ai Deng and Pierre Perron. A non-local perspective on the power properties of the CUSUM and CUSUM of squares tests for structural change. Journal of Econometrics, 142(1):212–240, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607001285.

Dette:2008:TPF

[DP08b] Holger Dette and Mark Podolskij. Testing the parametric form of the volatility in continuous time diffusion models — a stochastic process approach. Journal of Econometrics, 143(1):56–73, March 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607001595.

dePaula:2009:ISG

[dP09] Áureo de Paula. Inference in a synchronization game with social interactions. Journal of Econometrics, 148(1):56–71, January 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608001541.

Dufour:2006:SRL

[DPR06] Jean-Marie Dufour, Denis Pelletier, and Éric Renault. Short run and long run causality in time series: inference. Journal of Econometrics, 132(2):337–362, June 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000588.

Diebold:2006:MYC

[DRA06] Francis X. Diebold, Glenn D. Rudebusch, and S. Borağan Aruoba. The macroeconomy and the yield curve: a dynamic latent factor approach. Journal of Econometrics, 131(1–2):309–338, March/ April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760500014X.

Dahlen:2004:RCS

[DS04] Anders Dahlén and Wolfgang Scherrer. The relation of the CCA subspace method to a balanced reduction of an autoregressive model. Journal of Econometrics, 118(1–2):293–312, January 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603001441.

Davidson:2005:GSL

[DS05] James Davidson and Philipp Sibbertsen. Generating schemes for long memory processes: regimes, aggregation and linearity. Journal of Econometrics, 128(2):253–282, October 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001563.

Doran:2006:GEI

[DS06] Howard E. Doran and Peter Schmidt. GMM estimators with improved finite sample properties using principal components of the weighting matrix, with an application to the dynamic panel data model. Journal of Econometrics, 133(1):387–409, July 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000953.

Daouia:2007:NEA

[DS07] Abdelaati Daouia and Léopold Simar. Nonparametric efficiency analysis: A multivariate conditional quantile approach. Journal of Econometrics, 140(2):375–400, October 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001400.

Delgado:2008:DFS

[DS08a] Miguel A. Delgado and Winfried Stute. Distribution-free specification tests of conditional models. Journal of Econometrics, 143(1): 37–55, March 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001625.

Djebbari:2008:HIP

[DS08b] Habiba Djebbari and Jeffrey Smith. Heterogeneous impacts in PROGRESA. Journal of Econometrics, 145(1–2):64–80, July 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608000493.

Duggal:2007:IPE

[DSK07] Vijaya G. Duggal, Cynthia Saltzman, and Lawrence R. Klein. Infrastructure and productivity: An extension to private infrastructure and it productivity. Journal of Econometrics, 140(2):485– 502, October 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760600145X.

Dadpay:2007:IMG

[DSS07a] Ali Dadpay, Ehsan S. Soofi, and Refik Soyer. Information measures for generalized gamma family. Journal of Econometrics, 138(2):568–585, June 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606000832.

Dueker:2007:CTA

[DSS07b] Michael J. Dueker, Martin Sola, and Fabio Spagnolo. Contemporaneous threshold autoregressive models: Estimation, testing and forecasting. Journal of Econometrics, 141(2):517–547, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606002053.

Dufour:2000:MPT

[DT00] Jean-Marie Dufour and Olivier Torrès. Markovian processes, two-sided autoregressions and finite-sample inference for stationary and nonstationary autoregressive processes. Journal of Econometrics, 99(2):255–289, December 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000269.

Davidson:2002:LMN

[DT02] James Davidson and Timo Teräsvirta. Long memory and nonlinear time series. Journal of Econometrics, 110(2):105–112, October 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760200088X.

Dufour:2007:FRP

[DT07] Jean-Marie Dufour and Mohamed Taamouti. Further results on projection-based inference in IV regressions with weak, collinear or missing instruments. Journal of Econometrics, 139(1):133–153, July 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407606001060.

Devereux:2009:OCC

[DT09] Paul J. Devereux and Gautam Tripathi. Optimally combining censored and uncensored datasets. Journal of Econometrics, 151 (1):17–32, July 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609000852.

Dubin:2007:VIA

[Dub07] Jeffrey A. Dubin. Valuing intangible assets with a nested logit market share model. Journal of Econometrics, 139(2):285–302, August 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606002156.

Dufour:2006:MCT

[Duf06] Jean-Marie Dufour. Monte Carlo tests with nuisance parameters: A general approach to finite-sample inference and nonstandard asymptotics. Journal of Econometrics, 133(2):443–477, August 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605001260.

Durlauf:2001:MGE

[Dur01] Steven N. Durlauf. Manifesto for a growth econometrics. Journal of Econometrics, 100(1):65–69, January 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000555.

Durham:2006:MCM

[Dur06] Garland B. Durham. Monte Carlo methods for estimating, smoothing, and filtering one- and two-factor stochastic volatility models. Journal of Econometrics, 133(1):273–305, July 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000916.

Delgado:2005:STL

[DV05] Miguel A. Delgado and Carlos Velasco. Sign tests for long-memory time series. Journal of Econometrics, 128(2):215–251, October 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001551.

Dufour:2009:EAT

[DV09] Jean-Marie Dufour and Pascale Valéry. Exact and asymptotic tests for possibly non-regular hypotheses on stochastic volatility models. Journal of Econometrics, 150(2):193–206, June 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002182.

Diebold:2001:FEM

[DW01] F. X. Diebold and Kenneth D. West. Forecasting and empirical methods in finance and macroeconomics. Journal of Econometrics, 105(1):1–3, November 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000677.

Ellison:2000:SFN

[EE00] Glenn Ellison and Sara Fisher Ellison. A simple framework for nonparametric specification testing. Journal of Econometrics, 96 (1):1–23, May 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407699000482.

Engle:2006:MIM

[EG06] Robert F. Engle and Giampiero M. Gallo. A multiple indicators model for volatility using intra-daily data. Journal of Econometrics, 131(1–2):3–27, March/April 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000047.

Egorov:2006:VFJ

[EHL06] Alexei V. Egorov, Yongmiao Hong, and Haitao Li. Validating forecasts of the joint probability density of bond yields: Can affine models beat random walk? Journal of Econometrics, 135(1–2):255–284, November/December 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001697.

Ehrenberg:2004:ESH

[Ehr04] Ronald G. Ehrenberg. Econometric studies of higher education. Journal of Econometrics, 121(1–2):19–37, July 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002586.

Eichler:2007:GCP

[Eic07] Michael Eichler. Granger causality and path diagrams for multivariate time series. Journal of Econometrics, 137(2):334–353, April 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606000480.

Elliott:2003:TUR

[EJ03] Graham Elliott and Michael Jansson. Testing for unit roots with stationary covariates. Journal of Econometrics, 115(1):75–89, July 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603000939.

Eklof:2003:OOA

[EL03] Matias Eklöf and Anders Lunander. Open outcry auctions with secret reserve prices: an empirical application to executive auctions of tenant owner’s apartments in Sweden. Journal of Econometrics, 114(2):243–260, June 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603000848.

Evans:2005:BPC

[EL05] William N. Evans and Diana S. Lien. The benefits of prenatal care: evidence from the PAT bus strike. Journal of Econometrics, 125(1–2):207–239, March 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000788.

Escanciano:2009:APT

[EL09] J. Carlos Escanciano and Ignacio N. Lobato. An automatic portmanteau test for serial correlation. Journal of Econometrics, 151(2):140–149, August 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609000773.

Egorov:2003:MLE

[ELX03] Alexei V. Egorov, Haitao Li, and Yuewu Xu. Maximum likelihood estimation of time-inhomogeneous diffusions. Journal of Econometrics, 114(1):107–139, May 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760200221X.

Elliott:2006:MII

[EM06a] Graham Elliott and Ulrich K. Müller. Minimizing the impact of the initial condition on testing for unit roots. Journal of Econometrics, 135(1–2):285–310, November/December 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001740.

Engle:2006:LRP

[EM06b] Robert F. Engle and Juri Marcucci. A long-run Pure Variance Common Features model for the common volatilities of the Dow Jones. Journal of Econometrics, 132(1):7–42, May 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000394.

Elliott:2007:CSD

[EM07] Graham Elliott and Ulrich K. Müller. Confidence sets for the date of a single break in linear time series regressions. Journal of Econometrics, 141(2):1196–1218, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000401.

Engle:2001:FEN

[Eng01] Robert Engle. Financial econometrics — a new discipline with new methods. Journal of Econometrics, 100(1):53–56, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000531. See comments [Zel01].

Elliott:2001:CIA

[ES01] Graham Elliott and James H. Stock. Confidence intervals for autoregressive coefficients near one. Journal of Econometrics, 103(1–2):155–181, July 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407601000422.

Escanciano:2008:JMS

[Esc08] J. Carlos Escanciano. Joint and marginal specification tests for conditional mean and variance models. Journal of Econometrics, 143(1):74–87, March 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407607001674.

Elliott:2004:OFC

[ET04] Graham Elliott and Allan Timmermann. Optimal forecast combinations under general loss functions and forecast error distributions. Journal of Econometrics, 122(1):47–79, September 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603002690.

Eklund:2007:TCE

[ET07]  Bruno Eklund and Timo Teräsvirta.  Testing constancy

of the error covariance matrix in vector models. Journal of Econometrics, 140(2):753–780, October 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001540.

Escanciano:2006:GST

[EV06] J. Carlos Escanciano and Carlos Velasco. Generalized spectral tests for the martingale difference hypothesis. Journal of Econometrics, 134(1):151–185, September 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001417.

Einmahl:2008:STN

[EV08] John H. J. Einmahl and Ingrid Van Keilegom. Specification tests in nonparametric regression. Journal of Econometrics, 143(1):88– 102, March 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001704.

Eckstein:2007:ELS

[EvdB07] Zvi Eckstein and Gerard J. van den Berg. Empirical labor search: A survey. Journal of Econometrics, 136(2):531–564, February 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605002241.

ElBabsiri:2001:CAG

[EZ01] Mohamed El Babsiri and Jean-Michel Zakoian. Contemporaneous asymmetry in GARCH processes. Journal of Econometrics, 101(2):257–294, April 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407600000841.

Francke:2007:MLU

[FdV07] Marc K. Francke and Aart F. de Vos. Marginal likelihood and unit roots. Journal of Econometrics, 137(2):708–728, April 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760600073X.

Fok:2007:MDS

[FF07] Dennis Fok and Philip Hans Franses. Modeling the diffusion of scientific publications. Journal of Econometrics, 139(2):376– 390, August 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760600220X.

Fan:2008:HDC

[FFL08] Jianqing Fan, Yingying Fan, and Jinchi Lv. High dimensional covariance matrix estimation using a factor model. Journal of Econometrics, 147(1):186–197, November 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001346.

Fok:2007:SNL

[FFP07] Dennis Fok, Philip Hans Franses, and Richard Paap. Seasonality and non-linear price effects in scanner-data-based marketresponse models. Journal of Econometrics, 138(1):231–251, May 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606000947.

Fernandes:2005:NST

[FG05] Marcelo Fernandes and Joachim Grammig. Nonparametric specification tests for conditional duration models. Journal of Econometrics, 127(1):35–68, July 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760400137X.

Fernandes:2006:FAC

[FG06] Marcelo Fernandes and Joachim Grammig. A family of autoregressive conditional duration models. Journal of Econometrics, 130(1):1–23, January 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604001587.

Fare:2008:EDD

[FGHM08] Rolf Färe, Shawna Grosskopf, Kathy J. Hayes, and Dimitris Margaritis. Estimating demand with distance functions: Parameterization in the primal and dual. Journal of Econometrics, 147(2):266–274, December 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001425.

Fare:2005:CPT

[FGNW05] Rolf Färe, Shawna Grosskopf, Dong-Woon Noh, and William Weber. Characteristics of a polluting technology: theory and practice. Journal of Econometrics, 126(2):469–492, June 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001204.

Fienberg:2003:EST

[FGS03] Stephen E. Fienberg, Clark Glymour, and Richard Scheines. Expert statistical testimony and epidemiological evidence: the toxic effects of lead exposure on children. Journal of Econometrics, 113(1):33–48, March 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602001653.

Frederiksen:2007:DTD

[FHH07] Anders Frederiksen, Bo E. Honoré, and Luojia Hu. Discrete time duration models with group-level heterogeneity. Journal of Econometrics, 141(2):1014–1043, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002569.

Forni:2004:GDF

[FHLR04] Mario Forni, Marc Hallin, Marco Lippi, and Lucrezia Reichlin. The generalized dynamic factor model consistency and rates. Journal of Econometrics, 119(2):231–255, April 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001969.

Fiteni:2004:ERM

[Fit04] Inmaculada Fiteni. τ-estimators of regression models with structural change of unknown location. Journal of Econometrics, 119(1):19–44, March 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603001532.

Fernandez:2000:BAM

[FKS00] Carmen Fernández, Gary Koop, and Mark Steel. A Bayesian analysis of multiple-output production frontiers. Journal of Econometrics, 98(1):47–79, September 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000743.

Fernandez:2005:AEM

[FKS05] Carmen Fernández, Gary Koop, and Mark F. J. Steel. Alternative efficiency measures for multiple-output production. Journal of Econometrics, 126(2):411–444, June 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001186.

Lee:2000:NSQ

[fL00] Lung fei Lee. A numerically stable quadrature procedure for the one-factor random-component discrete choice model. Journal of Econometrics, 95(1):117–129, March 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000329.

Frazis:2003:ELR

[FL03] Harley Frazis and Mark A. Loewenstein. Estimating linear regressions with mismeasured, possibly endogenous, binary explanatory variables. Journal of Econometrics, 117(1):151–178, November 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603001210.

Lee:2007:GEM

[fL07a] Lung fei Lee. GMM and 2SLS estimation of mixed regressive, spatial autoregressive models. Journal of Econometrics, 137(2): 489–514, April 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000662.

Lee:2007:IEE

[fL07b] Lung fei Lee. Identification and estimation of econometric models with group interactions, contextual factors and fixed effects. Journal of Econometrics, 140(2):333–374, October 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001394.

Lee:2007:MES

[fL07c] Lung fei Lee. The method of elimination and substitution in the GMM estimation of mixed regressive, spatial autoregressive models. Journal of Econometrics, 140(1):155–189, September 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606002272.

Florens:2003:STI

[Flo03] Jean-Pierre Florens. Some technical issues in defining causality. Journal of Econometrics, 112(1):127–128, January 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001550.

Fernandez:2001:BPB

[FLS01] Carmen Fernández, Eduardo Ley, and Mark F. J. Steel. Benchmark priors for Bayesian model averaging. Journal of Econometrics, 100(2):381–427, February 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000762.

Fernandes:2007:SME

[FLS07] Marcelo Fernandes, Oliver Linton, and Olivier Scaillet. Semiparametric methods in econometrics. Journal of Econometrics, 141(1): 1–4, November 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760700005X.

Fay:2009:ELM

[FMRT09] Gilles Faÿ, Eric Moulines, François Roueff, and Murad S. Taqqu. Estimators of long-memory: Fourier versus wavelets. Journal of Econometrics, 151(2):159–177, August 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000803.

Francq:2008:CSU

[FMZ08] Christian Francq, Svetlana Makarova, and Jean-Michel Zakoiän. A class of stochastic unit-root bilinear processes: Mixing properties and unit-root test. Journal of Econometrics, 142(1):312– 326, January 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001327.

Franke:2004:BNE

[FNS04] Jürgen Franke, Michael H. Neumann, and Jean-Pierre Stockis. Bootstrapping nonparametric estimators of the volatility function. Journal of Econometrics, 118(1–2):189–218, January 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001404.

Forchini:2005:OWA

[For05] Giovanni Forchini. Optimal weighted average power similar tests for the covariance structure in the linear regression model. Journal of Econometrics, 124(2):253–267, February 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000338.

Fortuna:2008:LRT

[For08] Natércia Fortuna. Local rank tests in a multivariate nonparametric relationship. Journal of Econometrics, 142(1):162–182, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760700125X.

Findley:2004:MTS

[FPW04] David F. Findley, Benedikt M. Pötscher, and Ching-Zong Wei. Modeling of time series arrays by multistep prediction or likelihood methods. Journal of Econometrics, 118(1–2):151–187, January 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603001398.

Frechette:2009:LMB

[Fré09] Guillaume R. Fréchette. Learning in a multilateral bargaining experiment. Journal of Econometrics, 153(2):183–195, December 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609001444.

Frolich:2007:NIE

[Frö07] Markus Frölich. Nonparametric IV estimation of local average treatment effects with covariates. Journal of Econometrics, 139 (1):35–75, July 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407606001023.

Fraser:2005:HAL

[FRW05] D. A. S. Fraser, M. Rekkas, and A. Wong. Highly accurate likelihood analysis for the seemingly unrelated regression problem. Journal of Econometrics, 127(1):17–33, July 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001356.

Florens:2005:PAN

[FS05] Jean-Pierre Florens and Léopold Simar. Parametric approximations of nonparametric frontiers. Journal of Econometrics, 124(1):91–116, January 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604000272.

Ferreira:2007:MCC

[FS07] José T. A. S. Ferreira and Mark F. J. Steel. Model comparison of coordinate-free multivariate skewed distributions with an application to stochastic frontiers. Journal of Econometrics, 137(2): 641–673, April 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000716.

Feng:2008:PTU

[FS08a] Guohua Feng and Apostolos Serletis. Productivity trends in U.S. manufacturing: Evidence from the NQ and AIM cost functions. Journal of Econometrics, 142(1):281–311, January 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001315.

Florios:2008:ECM

[FS08b] Kostas Florios and Spyros Skouras. Exact computation of max weighted score estimators. Journal of Econometrics, 146(1):86– 91, September 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000778.

Formby:2004:MMT

[FSZ04] John P. Formby, W. James Smith, and Buhong Zheng. Mobility measurement, transition matrices and statistical inference. Journal of Econometrics, 120(1):181–205, May 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002112.

Froeb:2003:BCC

[FTC03] Luke Froeb, Steven Tschantz, and Philip Crooke. Bertrand competition with capacity constraints: mergers among parking lots. Journal of Econometrics, 113(1):49–67, March 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001665.

Fernandez-Val:2009:FEE

[FV09] Iván Fernández-Val. Fixed effects estimation of structural parameters and marginal effects in panel probit models. Journal of Econometrics, 150(1):71–85, May 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000463.

Franses:2007:PCE

[FvD07] Philip Hans Franses and Herman K. van Dijk. Progress and challenges in econometrics. Journal of Econometrics, 138(1):1–2, May 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606000844.

Fernandez-Villaverde:2004:CDE

[FVRR04] Jesús Fernández-Villaverde and Juan Francisco Rubio-Ramírez. Comparing dynamic equilibrium models to data: a Bayesian approach. Journal of Econometrics, 123(1):153–187, November 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407603002860.

Faust:2008:EFT

[FW08a] Jon Faust and Jonathan H. Wright. Efficient forecast tests for conditional policy forecasts. Journal of Econometrics, 146(2):293– 303, October 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001073.

Fleissig:2008:NTW

[FW08b] Adrian R. Fleissig and Gerald A. Whitney. A nonparametric test of weak separability and consumer preferences. Journal of Econometrics, 147(2):275–281, December 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001437.

Francq:2001:SMM

[FZ01] C. Francq and J.-M. Zakoïan. Stationarity of multivariate Markovswitching ARMA models. Journal of Econometrics, 102(2):339– 364, June 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000574.

Galvao:2009:URQ

[Gal09] Antonio F. Galvao. Unit root quantile autoregression testing using covariates. Journal of Econometrics, 152(2):165–178, October 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609000165.

Gastwirth:2003:IAU

[Gas03] Joseph L. Gastwirth. Issues arising in using samples as evidence in trademark cases. Journal of Econometrics, 113(1):69– 82, March 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602001677.

Gao:2008:STD

[GC08] Jiti Gao and Isabel Casas. Specification testing in discretized diffusion models: Theory and practice. Journal of Econometrics, 147(1):131–140, November 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001309.

Geweke:2001:BEF

[Gew01] John Geweke. Bayesian econometrics and forecasting. Journal of Econometrics, 100(1):11–15, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0304407600000464. See comments [LW01].

Geweke:2003:EIU

[Gew03] John Geweke. Econometric issues in using the AHEAD panel. Journal of Econometrics, 112(1):115–120, January 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001537.

Garcia:2000:PHD

[GG00] René Garcia and Ramazan Gençay. Pricing and hedging derivative securities with neural networks and a homogeneity hint. Journal of Econometrics, 94(1–2):93–115, January 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000184.

Ghysels:2003:SCT

[GG03] Eric Ghysels and Alain Guay. Structural change tests for simulated method of moments. Journal of Econometrics, 115(1): 91–123, July 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603000940.

Giacomini:2004:AST

[GG04] Raffaella Giacomini and Clive W. J. Granger. Aggregation of space-time processes. Journal of Econometrics, 118(1–2):7–26, January 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603001325.

Gagliardini:2007:ENE

[GG07] Patrick Gagliardini and Christian Gouriéroux. An efficient nonparametric estimator for models with nonlinear dependence. Journal of Econometrics, 137(1):189–229, March 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000431.

Giacomini:2008:MDF

[GGHW08] Raffaella Giacomini, Andreas Gottschling, Christian Haefke, and Halbert White. Mixtures of t-distributions for finance and forecasting. Journal of Econometrics, 144(1):175–192, May 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000055.

Ghysels:2004:SVD

[GGJ04] Eric Ghysels, Christian Gouriéroux, and Joann Jasiak. Stochastic volatility duration models. Journal of Econometrics, 119(2):413– 433, April 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603002021.

Gao:2008:NST

[GGV08] Jiti Gao, Irène Gijbels, and Sébastien Van Bellegem. Nonparametric simultaneous testing for structural breaks. Journal of Econometrics, 143(1):123–142, March 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001686.

Granger:2006:IP

[GH06a] Clive W. J. Granger and Namwon Hyung. Introduction to m– m processes. Journal of Econometrics, 130(1):143–164, January 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605000515.

Guell:2006:EPL

[GH06b] Maia Güell and Luojia Hu. Estimating the probability of leaving unemployment using uncompleted spells from repeated crosssection data. Journal of Econometrics, 133(1):307–341, July 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605000928.

Galichon:2009:TNI

[GH09] Alfred Galichon and Marc Henry. A test of non-identifying restrictions and confidence regions for partially identified parameters. Journal of Econometrics, 152(2):186–196, October 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000141.

Girma:2000:QDA

[Gir00] Sourafel Girma. A quasi-differencing approach to dynamic modelling from a time series of independent cross-sections. Journal of Econometrics, 98(2):365–383, October 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000245.

Gourieroux:2006:MJP

[GJ06] Christian Gourieroux and Joann Jasiak. Multivariate Jacobi process with application to smooth transitions. Journal of Econometrics, 131(1–2):475–505, March/April 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000199.

Gourieroux:2008:DQM

[GJ08] C. Gourieroux and J. Jasiak. Dynamic quantile models. Journal of Econometrics, 147(1):198–205, November 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001358.

Gourieroux:2009:WAP

[GJS09] C. Gourieroux, J. Jasiak, and R. Sufana. The Wishart autoregressive process of multivariate stochastic volatility. Journal of Econometrics, 150(2):167–181, June 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002169.

Geweke:2000:EAE

[GK00] John Geweke and Michael Keane. An empirical analysis of earnings dynamics among men in the PSID: 1968–1989. Journal of Econometrics, 96(2):293–356, June 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000639.

Goncalves:2004:BAC

[GK04] Sílvia Gonçalves and Lutz Kilian. Bootstrapping autoregressions with conditional heteroskedasticity of unknown form. Journal of Econometrics, 123(1):89–120, November 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/S0304407603002847.

Geweke:2007:SMR

[GK07] John Geweke and Michael Keane. Smoothly mixing regressions. Journal of Econometrics, 138(1):252–290, May 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000959.

Giraitis:2003:RVR

[GKLT03] Liudas Giraitis, Piotr Kokoszka, Remigijus Leipus, and Gilles Teyssière. Rescaled variance and related tests for long memory in volatility and levels. Journal of Econometrics, 112(2):265– 294, February 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602001975. See corrigendum

[GKLT05].

Giraitis:2005:CRV

[GKLT05] Liudas Giraitis, Piotr Kokoszka, Remigijus Leipus, and Gilles Teyssière. Corrigendum to “Rescaled variance and related tests for long memory in volatility and levels”: [J. Econom. 112 (2003) 265–294]. Journal of Econometrics, 126(2):571–572, June 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001423. See [GKLT03].

Giordani:2007:UAN

[GKvD07] Paolo Giordani, Robert Kohn, and Dick van Dijk. A unified approach to nonlinearity, structural change, and outliers. Journal of Econometrics, 137(1):112–133, March 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000406.

Gao:2000:FCV

[GL00a] Chuanming Gao and Kajal Lahiri. Further consequences of viewing LIML as an iterated Aitken estimator. Journal of Econometrics, 98(2):187–202, October 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000834.

Gozalo:2000:LNL

[GL00b] Pedro Gozalo and Oliver Linton. Local nonlinear least squares: Using parametric information in nonparametric regression. Journal of Econometrics, 99(1):63–106, November 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000312.

Gozalo:2001:TAG

[GL01] Pedro L. Gozalo and Oliver B. Linton. Testing additivity in generalized nonparametric regression models with estimated parameters. Journal of Econometrics, 104(1):1–48, August 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000495.

Gao:2002:NDC

[GL02] Chuanming Gao and Kajal Lahiri. A note on the double kclass estimator in simultaneous equations. Journal of Econometrics, 108(1):101–111, May 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407601001075.

Garcia:2003:EAI

[GLR03] René Garcia, Richard Luger, and Eric Renault. Empirical assessment of an intertemporal option pricing model with latent variables. Journal of Econometrics, 116(1–2):49–83, September 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603001039.

Gregory:2002:ITE

[GLS02] Allan W. Gregory, Jean-François Lamarche, and Gregor W. Smith. Information-theoretic estimation of preference parameters: macroeconomic applications and simulation evidence. Journal of Econometrics, 107(1–2):213–233, March 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760100121X.

Gonzalo:2006:LSV

[GM06a] Jesús Gonzalo and Oscar Martínez. Large shocks vs. small shocks. (or does size matter? May be so.). Journal of Econometrics, 135(1–2):311–347, November/December 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001727.

Granger:2006:SAO

[GM06b] Clive W. J. Granger and Mark J. Machina. Structural attribution of observed volatility clustering. Journal of Econometrics, 135(1–2):15–29, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001673.

Gourieroux:2007:ESS

[GM07] C. Gourieroux and A. Monfort. Econometric specification of stochastic discount factor models. Journal of Econometrics, 136(2):509–530, February 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760500223X.

Gao:2008:EMF

[GMA08] Jiti Gao, Michael McAleer, and David E. Allen. Econometric modelling in finance and risk management: An overview.

Journal of Econometrics, 147(1):1–4, November 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760800119X.

Griffiths:2005:EVR

[GO05] William E. Griffiths and Christopher J. O’Donnell. Estimating variable returns to scale production frontiers with alternative stochastic assumptions. Journal of Econometrics, 126(2): 385–409, June 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604001174.

Greasley:2007:PIP

[GO07] David Greasley and Les Oxley. Patenting, intellectual property rights and sectoral outputs in Industrial Revolution Britain, 1780– 1851. Journal of Econometrics, 139(2):340–354, August 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002181.

Golan:2001:SEV

[Gol01] Amos Golan. A simultaneous estimation and variable selection rule. Journal of Econometrics, 101(1):165–193, March 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000786.

Golan:2002:IEE

[Gol02] Amos Golan. Information and entropy econometrics — Editor’s view. Journal of Econometrics, 107(1–2):1–15, March 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001105.

Golan:2007:IEE

[Gol07] Amos Golan. Information and entropy econometrics — volume overview and synthesis. Journal of Econometrics, 138(2):379– 387, June 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000741.

Gospodinov:2002:MUF

[Gos02] Nikolay Gospodinov. Median unbiased forecasts for highly persistent autoregressive processes. Journal of Econometrics, 111(1):85–101, November 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602001239.

Gospodinov:2008:ABT

[Gos08] Nikolay Gospodinov. Asymptotic and bootstrap tests for linearity in a TAR-GARCH(1,1) model with a unit root. Journal of Econometrics, 146(1):146–161, September 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000821.

Golan:2002:CME

[GP02a] Amos Golan and Jeffrey M. Perloff. Comparison of maximum entropy and higher-order entropy estimators. Journal of Econometrics, 107(1–2):195–211, March 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001208.

Gonzalo:2002:EMS

[GP02b] Jesús Gonzalo and Jean-Yves Pitarakis. Estimation and model selection based inference in single and multiple threshold models. Journal of Econometrics, 110(2):319–352, October 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602000982.

Granger:2001:MPF

[Gra01] Clive W. J. Granger. Macroeconometrics — past and future. Journal of Econometrics, 100(1):17–19, January 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0304407600000476. See [Zel01, Dei01].

Granger:2003:SAC

[Gra03] Clive W. J. Granger. Some aspects of causal relationships. Journal of Econometrics, 112(1):69–71, January 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001483.

Granger:2005:PFE

[Gra05] Clive W. J. Granger. The past and future of empirical finance: some personal comments. Journal of Econometrics, 129(1–2):35– 40, November 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604001642.

Granger:2006:OCP

[Gra06] Clive W. J. Granger. Opening comments: Predictive methodology and application in economics and finance: Presentation for the San Diego Conference, January, 2004. Journal of Econometrics, 135(1–2):11–13, November/December 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001624.

Granger:2007:FLB

[Gra07] Clive W. J. Granger. Forecasting-looking back and forward: Paper to celebrate the 50th anniversary of the Econometrics Institute at the Erasmus University, Rotterdam. Journal of Econometrics, 138 (1):3–13, May 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000856.

Greene:2005:RHP

[Gre05] William Greene. Reconsidering heterogeneity in panel data estimators of the stochastic frontier model. Journal of Econometrics, 126(2):269–303, June 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604001137.

Gregoir:2006:ETP

[Gre06] Stéphane Gregoir. Efficient tests for the presence of a pair of complex conjugate unit roots in real time series. Journal of Econometrics, 130(1):45–100, January 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001617.

Groen:2004:ECL

[Gro04] Jeffrey A. Groen. The effect of college location on migration of college-educated labor. Journal of Econometrics, 121(1–2):125– 142, July 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603002525.

Giannone:2006:VCF

[GRS06] Domenico Giannone, Lucrezia Reichlin, and Luca Sala. VARs, common factors and the empirical validation of equilibrium business cycle models. Journal of Econometrics, 132(1):257–279, May 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605000485.

Gaure:2007:TCM

[GRZ07] Simen Gaure, Knut Røed, and Tao Zhang. Time and causality: A Monte Carlo assessment of the timing-of-events approach. Journal of Econometrics, 141(2):1159–1195, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000280.

Griffin:2004:SBI

[GS04] J. E. Griffin and M. F. J. Steel. Semiparametric Bayesian inference for stochastic frontier models. Journal of Econometrics, 123(1):121–152, November 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407603002859.

Griffin:2006:ING

[GS06] J. E. Griffin and M. F. J. Steel. Inference with non-Gaussian Ornstein–Uhlenbeck processes for stochastic volatility. Journal of Econometrics, 134(2):605–644, October 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001570.

Guggenberger:2008:GEL

[GS08] Patrik Guggenberger and Richard J. Smith. Generalized empirical likelihood tests in time series models with potential identification failure. Journal of Econometrics, 142(1):134–161, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607001248.

Ghysels:2006:PVG

[GSCV06] Eric Ghysels, Pedro Santa-Clara, and Rossen Valkanov. Predicting volatility: getting the most out of return data sampled at different frequencies. Journal of Econometrics, 131(1–2):59–95, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605000060.

George:2008:BSS

[GSN08] Edward I. George, Dongchu Sun, and Shawn Ni. Bayesian stochastic search for VAR model restrictions. Journal of Econometrics, 142(1):553–580, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407607001753.

Ghysels:2003:FFE

[GT03] Eric Ghysels and George Tauchen. Frontiers of financial econometrics and financial engineering. Journal of Econometrics, 116(1–2): 1–7, September 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603001015.

Guidolin:2006:TSR

[GT06] Massimo Guidolin and Allan Timmermann. Term structure of risk under alternative econometric specifications. Journal of Econometrics, 131(1–2):285–308, March/April 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000138.

Guidolin:2009:FUS

[GT09] Massimo Guidolin and Allan Timmermann. Forecasts of US short-term interest rates: A flexible forecast combination approach. Journal of Econometrics, 150(2):297–311, June 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002261.

Granger:2006:CFC

[GTP06] Clive W. J. Granger, Timo Teräsvirta, and Andrew J. Patton. Common factors in conditional distributions for bivariate time series. Journal of Econometrics, 132(1):43–57, May 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000400.

Gagliardini:2005:RGT

[GTU05] Patrick Gagliardini, Fabio Trojani, and Giovanni Urga. Robust GMM tests for structural breaks. Journal of Econometrics, 129(1–2):139–182, November 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760400168X.

Gayle:2007:RCS

[GV07] George-Levi Gayle and Christelle Viauroux. Root-N consistent semiparametric estimators of a dynamic panel-sample-selection model. Journal of Econometrics, 141(1):179–212, November 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607000139.

Grammig:2002:MIV

[GW02] Joachim Grammig and Marc Wellner. Modeling the interdependence of volatility and inter-transaction duration processes. Journal of Econometrics, 106(2):369–400, February 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001051.

Goncalves:2004:MLB

[GW04] Sílvia Gonçalves and Halbert White. Maximum likelihood and the bootstrap for nonlinear dynamic models. Journal of Econometrics, 119(1):199–219, March 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002045.

Gonzalo:2005:SIT

[GW05] Jesús Gonzalo and Michael Wolf. Subsampling inference in threshold autoregressive models. Journal of Econometrics, 127(2):201– 224, August 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604001460.

Gan:2006:TME

[GZ06] Li Gan and Qinghua Zhang. The thick market effect on local unemployment rate fluctuations. Journal of Econometrics, 133(1): 127–152, July 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605000862.

Hafner:2008:TAM

[Haf08] Christian M. Hafner. Temporal aggregation of multivariate GARCH processes. Journal of Econometrics, 142(1):467–483, January 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001583.

Hamilton:2003:WOS

[Ham03]  James D. Hamilton. What is an oil shock?  Jour-

nal of Econometrics, 113(2):363–398, April 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602002075.

Hansen:2000:TSC

[Han00] Bruce E. Hansen. Testing for structural change in conditional models. Journal of Econometrics, 97(1):93–115, July 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000688.

Hansen:2003:SCC

[Han03] Peter Reinhard Hansen. Structural changes in the cointegrated vector autoregressive model. Journal of Econometrics, 114(2): 261–295, June 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760300085X.

Hansen:2006:IFP

[Han06] Bruce E. Hansen. Interval forecasts and parameter uncertainty. Journal of Econometrics, 135(1–2):377–398, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605001806.

Hansen:2007:APR

[Han07a] Christian B. Hansen. Asymptotic properties of a robust variance matrix estimator for panel data when t is large. Journal of Econometrics, 141(2):597–620, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002089.

Hansen:2007:GLS

[Han07b] Christian B. Hansen. Generalized least squares inference in panel and multilevel models with serial correlation and fixed effects. Journal of Econometrics, 140(2):670–694, October 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001515.

Hansen:2008:LSF

[Han08] Bruce E. Hansen. Least-squares forecast averaging. Journal of Econometrics, 146(2):342–350, October 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001097.

Hasan:2001:RTU

[Has01] Mohammad N. Hasan. Rank tests of unit root hypothesis with infinite variance errors. Journal of Econometrics, 104(1):49–65, August 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601000501.

Hausman:2001:M

[Hau01a] Jerry Hausman. Microeconometrics. Journal of Econometrics, 100(1):33–35, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407600000506.

See comment [WWM01].

Hausman:2001:R

[Hau01b] Jerry Hausman. Rejoinder. Journal of Econometrics, 100(1):95– 96, January 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000579.

Hausman:2003:TSM

[Hau03] Jerry A. Hausman. Triangular structural model specification and estimation with application to causality. Journal of Econometrics, 112(1):107–113, January 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001525.

Hayakawa:2009:EMN

[Hay09] Kazuhiko Hayakawa. On the effect of mean-nonstationarity in dynamic panel data models. Journal of Econometrics, 153(2):133– 135, December 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001286.

Henderson:2008:NET

[HCL08] Daniel J. Henderson, Raymond J. Carroll, and Qi Li. Nonparametric estimation and testing of fixed effects panel data models. Journal of Econometrics, 144(1):257–275, May 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760800016X.

Hamermesh:2008:ECC

[HD08] Daniel S. Hamermesh and Stephen G. Donald. The effect of college curriculum on earnings: An affinity identifier for nonignorable non-response bias. Journal of Econometrics, 144(2): 479–491, June 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000419.

Heckman:2001:EEE

[Hec01] James J. Heckman. Econometrics and empirical economics. Journal of Econometrics, 100(1):3–5, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000440. See [Dei01].

Heckman:2003:CCP

[Hec03] James Heckman. Conditioning, causality and policy analysis. Journal of Econometrics, 112(1):73–78, January 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001495.

Hendry:2001:ACE

[Hen01] David F. Hendry. Achievements and challenges in econometric methodology. Journal of Econometrics, 100(1):7–10, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000452.

Hendry:2006:RFE

[Hen06] David F. Hendry. Robustifying forecasts from equilibriumcorrection systems. Journal of Econometrics, 135(1–2):399– 426, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760500179X.

Hardle:2000:IBE

[HH00] W. Härdle and J. Horowitz. Internet-based econometric computing. Journal of Econometrics, 95(2):333–345, April 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769900041X.

Honore:2004:ECS

[HH04] Bo E. Honoré and Luojia Hu. Estimation of cross sectional and panel data censored regression models with endogeneity. Journal of Econometrics, 122(2):293–316, October 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760300277X.

Hardle:2009:DSP

[HH09] Wolfgang Härdle and Zdenek Hlávka. Dynamics of state price densities. Journal of Econometrics, 150(1):1–15, May 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000244.

Hahn:2007:LDI

[HHK07] Jinyong Hahn, Jerry Hausman, and Guido Kuersteiner. Long difference instrumental variables estimation for dynamic panel models with fixed effects. Journal of Econometrics, 140(2):574– 617, October 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606001485.

Hansen:2004:ESA

[HHM04] Karsten T. Hansen, James J. Heckman, and Kathleen J. Mullen. The effect of schooling and ability on achievement test scores. Journal of Econometrics, 121(1–2):39–98, July 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002598.

Haag:2009:TIS

[HHP09] Berthold R. Haag, Stefan Hoderlein, and Krishna Pendakur. Testing and imposing Slutsky symmetry in nonparametric demand systems. Journal of Econometrics, 153(1):33–50, November 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609001092.

Hsieh:2007:ADD

[HHS07] Meng-Chen Hsieh, Clifford M. Hurvich, and Philippe Soulier. Asymptotics for duration-driven long range dependent processes. Journal of Econometrics, 141(2):913–949, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002521.

Hall:2003:LSB

[HI03] Alastair R. Hall and Atsushi Inoue. The large sample behaviour of the generalized method of moments estimator in misspecified models. Journal of Econometrics, 114(2):361–394, June 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0304407603000897. See corrigendum [HI07].

Hall:2007:CLS

[HI07] Alastair R. Hall and Atsushi Inoue. Corrigendum to: “The large sample behaviour of the generalized method of moments estimator in misspecified models”: [Journal of Econometrics 114 (2003) 361–394]. Journal of Econometrics, 141(2):1418, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607000449. See [HI03].

Hidalgo:2002:COS

[Hid02] Javier Hidalgo. Consistent order selection with strongly dependent data and its application to efficient estimation. Journal of Econometrics, 110(2):213–239, October 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602000945.

Hidalgo:2003:ABM

[Hid03] Javier Hidalgo. An alternative bootstrap to moving blocks for time series regression models. Journal of Econometrics, 117(2):369– 399, December 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603001544.

Hidalgo:2005:BCT

[Hid05] J. Hidalgo. A bootstrap causality test for covariance stationary processes. Journal of Econometrics, 126(1):115–143, May 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604000429.

Hidalgo:2008:STR

[Hid08] J. Hidalgo. Specification testing for regression models with dependent data. Journal of Econometrics, 143(1):143–165, March 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607001649.

Hidalgo:2009:GFL

[Hid09] Javier Hidalgo. Goodness of fit for lattice processes. Journal of Econometrics, 151(2):113–128, August 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000761.

Hall:2007:IGM

[HIJS07] Alastair R. Hall, Atsushi Inoue, Kalidas Jana, and Changmock Shin. Information in generalized method of moments estimation and entropy-based moment selection. Journal of Econometrics, 138(2):488–512, June 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000790.

Hillebrand:2005:NPC

[Hil05] Eric Hillebrand. Neglecting parameter changes in GARCH models. Journal of Econometrics, 129(1–2):121–138, November 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001678.

Hill:2006:SIN

[Hil06] Robert J. Hill. Superlative index numbers: not all of them are super. Journal of Econometrics, 130(1):25–43, January 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001605.

Hotz:2005:PEF

[HIM05] V. Joseph Hotz, Guido W. Imbens, and Julie H. Mortimer. Predicting the efficacy of future training programs using past experiences at other locations. Journal of Econometrics, 125(1–2):241– 270, March 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604000806.

Hasegawa:2003:ELC

[HK03] Hikaru Hasegawa and Hideo Kozumi. Estimation of Lorenz curves: a Bayesian nonparametric approach. Journal of Econometrics, 115(2):277–291, August 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603000988.

Hidalgo:2006:BST

[HK06] J. Hidalgo and J.-P. Kreiss. Bootstrap specification tests for linear covariance stationary processes. Journal of Econometrics, 133(2):807–839, August 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760500134X.

Hausman:2008:DDM

[HK08a] Jerry Hausman and Guido Kuersteiner. Difference in difference meets generalized least squares: Higher order properties of hypotheses tests. Journal of Econometrics, 144(2):371–391, June 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608000341.

Haywood:2008:DFG

[HK08b] John Haywood and Estate Khmaladze. On distribution-free goodness-of-fit testing of exponentiality. Journal of Econometrics, 143(1):5–18, March 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407607001613.

Ham:2005:REL

[HKL05] John C. Ham, John H. Kagel, and Steven F. Lehrer. Randomization, endogeneity and laboratory experiments: the role of cash balances in private value auctions. Journal of Econometrics, 125(1–2):175–205, March 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760400079X.

Honore:2002:QRU

[HKP02] Bo Honoré, Shakeeb Khan, and James L. Powell. Quantile regression under random censoring. Journal of Econometrics, 109(1): 67–105, July 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601001427.

Hoogerheide:2007:NCP

[HKvD07a] Lennart Hoogerheide, Frank Kleibergen, and Herman K. van Dijk. Natural conjugate priors for the instrumental variables regression model applied to the Angrist–Krueger data. Journal of Econometrics, 138(1):63–103, May 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000893.

Hoogerheide:2007:SPD

[HKvD07b] Lennart F. Hoogerheide, Johan F. Kaashoek, and Herman K. van Dijk. On the shape of posterior densities and credible sets in instrumental variable regression models with reduced rank: An application of flexible sampling methods using neural networks. Journal of Econometrics, 139(1):154–180, July 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407606001072.

Horvath:2006:TSD

[HKZ06] Lajos Horváth, Piotr Kokoszka, and Ricardas Zitikis. Testing for stochastic dominance using the weighted McFadden-type statistic. Journal of Econometrics, 133(1):191–205, July 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000886.

Horowitz:2004:SEP

[HL04] Joel L. Horowitz and Sokbae Lee. Semiparametric estimation of a panel data proportional hazards model with fixed effects. Journal of Econometrics, 119(1):155–198, March 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002033.

Ham:2005:SIE

[HL05] John C. Ham and Robert J. LaLonde. Special issue on experimental and non-experimental evaluation of economic policy and models. Journal of Econometrics, 125(1–2):1–13, March 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000727.

Hansen:2006:CRV

[HL06] Peter Reinhard Hansen and Asger Lunde. Consistent ranking of volatility models. Journal of Econometrics, 131(1–2):97–121, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605000072.

Hausman:2007:EPL

[HL07] Jerry A. Hausman and Gregory K. Leonard. Estimation of patent licensing value using a flexible demand specification. Journal of Econometrics, 139(2):242–258, August 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002132.

Horowitz:2009:TPQ

[HL09] Joel L. Horowitz and Sokbae Lee. Testing a parametric quantile-regression model with an endogenous explanatory variable against a nonparametric alternative. Journal of Econometrics, 152(2):141–152, October 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000189.

Horowitz:2006:BBP

[HLNS06] Joel L. Horowitz, I. N. Lobato, John C. Nankervis, and N. E. Savin. Bootstrapping the Box–Pierce Q test: a robust test of uncorrelatedness. Journal of Econometrics, 133(2):841–862, August 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605001351.

Hsiao:2007:CMS

[HLR07] Cheng Hsiao, Qi Li, and Jeffrey S. Racine. A consistent model specification test with mixed discrete and continuous data. Journal of Econometrics, 140(2):802–826, October 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001564.

Hirschberg:2008:IME

[HLS08] J. G. Hirschberg, J. N. Lye, and D. J. Slottje. Inferential methods for elasticity estimates. Journal of Econometrics, 147(2):299– 315, December 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001450.

Harvey:2006:MTC

[HLT06] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Modified tests for a change in persistence. Journal of Econometrics, 134(2):441–469, October 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0304407605001521. See corrigendum [HLT12].

Harvey:2007:SRP

[HLT07] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. A simple, robust and powerful test of the trend hypothesis.

Journal of Econometrics, 141(2):1302–1330, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0304407607000450. See erratum [HLT08].

Harvey:2008:ESR

[HLT08] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Erratum to “A simple, robust and powerful test of the trend hypothesis” [Journal of Econometrics 141(2) (2007) 1302– 1330]. Journal of Econometrics, 143(2):396–397, April 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607002448. See [HLT07].

Harvey:2012:CMT

[HLT12] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Corrigendum to “Modified tests for a change in persistence” [J. Econom. 134 (2006) 441–469]. Journal of Econometrics, 168 (2):407, June 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002570. See [HLT06].

Hong:2009:GCR

[HLW09] Yongmiao Hong, Yanhui Liu, and Shouyang Wang. Granger causality in risk and detection of extreme risk spillover between financial markets. Journal of Econometrics, 150(2):271–287, June 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608002248.

Hong:2007:CRW

[HLZ07] Yongmiao Hong, Haitao Li, and Feng Zhao. Can the random walk model be beaten in out-of-sample density forecasts? Evidence from intraday foreign exchange rates. Journal of Econometrics, 141(2):736–776, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002351.

Heckelei:2003:BBM

[HM03] Thomas Heckelei and Ron C. Mittelhammer. Bayesian bootstrap multivariate regression. Journal of Econometrics, 112(2):241– 264, February 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602001963.

Horowitz:2006:IES

[HM06] Joel L. Horowitz and Charles F. Manski. Identification and estimation of statistical functionals using incomplete data. Journal of Econometrics, 132(2):445–459, June 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760500062X.

Hoderlein:2008:IPV

[HM08] Stefan Hoderlein and Sonya Mihaleva. Increasing the price variation in a repeated cross section. Journal of Econometrics, 147(2):316–325, December 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407608001462.

Harris:2002:SCE

[HML02] David Harris, Brendan McCabe, and Stephen Leybourne. Stochastic cointegration: estimation and inference. Journal of Econometrics, 111(2):363–384, December 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001112.

Haldrup:2005:MEO

[HMS05] Niels Haldrup, Antonio Montanés, and Andreu Sanso. Measurement errors and outliers in seasonal unit root testing. Journal of Econometrics, 127(1):103–128, July 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760400140X.

Hirschberg:2003:AII

[HMSA03] Joseph G. Hirschberg, Esfandiar Maasoumi, Daniel Slottje, and Augustine C. Arize. Antitrust issues in international comparisons of market structure. Journal of Econometrics, 113(1):129– 158, March 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602001707.

Hassler:2006:RLP

[HMV06] U. Hassler, F. Marmol, and C. Velasco. Residual logperiodogram inference for long-run relationships. Journal of Econometrics, 130(1):165–207, January 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000667.

Huang:2000:NMC

[HN00] Ju-Chin Huang and Douglas W. Nychka. A nonparametric multiple choice method within the random utility framework. Journal of Econometrics, 97(2):207–225, August 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769900072X.

Herwartz:2005:BIS

[HN05] Helmut Herwartz and Michael H. Neumann. Bootstrap inference in systems of single equation error correction models. Journal of Econometrics, 128(1):165–193, September 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001538.

Haldrup:2006:RSL

[HN06] Niels Haldrup and Morten Ørregaard Nielsen. A regime switching long memory model for electricity prices. Journal of Econometrics, 135(1–2):349–376, November/December 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001715.

Heckman:2007:DDC

[HN07] James J. Heckman and Salvador Navarro. Dynamic discrete choice and dynamic treatment effects. Journal of Econometrics, 136(2):341–396, February 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605002186.

Hong:2001:TVS

[Hon01] Yongmiao Hong. A test for volatility spillover with application to exchange rates. Journal of Econometrics, 103(1–2):183–224, July 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601000434.

Hoover:2003:SCL

[Hoo03] Kevin D. Hoover. Some causal lessons from macroeconomics. Journal of Econometrics, 112(1):121–125, January 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001549.

Horowitz:2001:RCE

[Hor01a] J. L. Horowitz. Response to comments of esfandiar maasoumi. Journal of Econometrics, 100(1):97–98, January 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000580.

Horowitz:2001:BHT

[Hor01b] Joel L. Horowitz. The bootstrap and hypothesis tests in econometrics. Journal of Econometrics, 100(1):37–40, January 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000518.

Horowitz:2002:BCV

[Hor02] Joel L. Horowitz. Bootstrap critical values for tests based on the smoothed maximum score estimator. Journal of Econometrics, 111(2):141–167, December 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001021.

Horrace:2005:RSI

[Hor05] William C. Horrace. On ranking and selection from independent truncated normal distributions. Journal of Econometrics, 126(2): 335–354, June 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604001150.

Han:2005:EPD

[HOS05] Chirok Han, Luis Orea, and Peter Schmidt. Estimation of a panel data model with parametric temporal variation in individual effects. Journal of Econometrics, 126(2):241–267, June 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001125.

Houser:2003:BAD

[Hou03] Daniel Houser. Bayesian analysis of a dynamic stochastic model of labor supply and saving. Journal of Econometrics, 113(2):289– 335, April 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602002051.

Hsiao:2001:SET

[HP01] Cheng Hsiao and Isabelle Perrigne. Studies in estimation and testing. Journal of Econometrics, 103(1–2):1–4, July 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000380.

Hu:2004:NDC

[HP04] Ling Hu and Peter C. B. Phillips. Nonstationary discrete choice. Journal of Econometrics, 120(1):103–138, May 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002082.

Harding:2006:SC

[HP06] Don Harding and Adrian Pagan. Synchronization of cycles. Journal of Econometrics, 132(1):59–79, May 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000412.

Han:2008:TSP

[HP08] Heejoon Han and Joon Y. Park. Time series properties of ARCH processes with persistent covariates. Journal of Econometrics, 146(2):275–292, October 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407608001152.

Hsiao:2002:MLE

[HPT02] Cheng Hsiao, M. Hashem Pesaran, and A. Kamil Tahmiscioglu. Maximum likelihood estimation of fixed effects dynamic panel data models covering short time periods. Journal of Econometrics, 109(1):107–150, July 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001439.

Herriges:2008:EDS

[HPT08] Joseph A. Herriges, Daniel J. Phaneuf, and Justin L. Tobias. Estimating demand systems when outcomes are correlated counts. Journal of Econometrics, 147(2):282–298, December 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001449.

Hecq:2006:CCF

[HPU06] Alain Hecq, Franz C. Palm, and Jean-Pierre Urbain. Common cyclical features analysis in VAR models with cointegration. Journal of Econometrics, 132(1):117–141, May 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000436.

Hualde:2007:RCE

[HR07] J. Hualde and P. M. Robinson. Root-n-consistent estimation of weak fractional cointegration. Journal of Econometrics, 140(2):450–484, October 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606001448.

Horowitz:2000:ERC

[HS00] Joel L. Horowitz and N. E. Savin. Empirically relevant critical values for hypothesis tests: A bootstrap approach. Journal of Econometrics, 95(2):375–389, April 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000421.

Hansen:2002:TTR

[HS02a] Bruce E. Hansen and Byeongseon Seo. Testing for two-regime threshold cointegration in vector error-correction models. Journal of Econometrics, 110(2):293–318, October 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602000970.

Huang:2002:SCC

[HS02b] Tai-Hsin Huang and Chung-Hua Shen. Seasonal cointegration and cross-equation restrictions on a forward-looking buffer stock model of money demand. Journal of Econometrics, 111(1):11– 46, November 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602001148.

Hong:2003:EMA

[HS03] Han Hong and Matthew Shum. Econometric models of asymmetric ascending auctions. Journal of Econometrics, 112(2):327– 358, February 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602001999.

Hong:2006:FSM

[HS06] H. Hong and O. Scaillet. A fast subsampling method for nonlinear dynamic models. Journal of Econometrics, 133(2):557– 578, August 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001223.

Hagmann:2007:LMB

[HS07] M. Hagmann and O. Scaillet. Local multiplicative bias correction for asymmetric kernel density estimators. Journal of Econometrics, 141(1):213–249, November 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000140.

Hsiao:2001:OFC

[Hsi01] C. Hsiao. Open forum on the current state and future challenges of econometrics. Journal of Econometrics, 100(1):1, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000439.

Hsiao:2009:AEA

[Hsi09] Cheng Hsiao. Announcement of the establishment of the Amemiya lecture series. Journal of Econometrics, 149(1):1, April 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000682.

Hsiao:2008:EEW

[HSWW08] Cheng Hsiao, Yan Shen, Boqing Wang, and Greg Weeks. Evaluating the effectiveness of Washington state repeated job search services on the employment rate of prime-age female welfare recipients. Journal of Econometrics, 145(1–2):98–108, July 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000511.

Hong:2003:SEN

[HT03] Han Hong and Elie Tamer. A simple estimator for nonlinear error in variable models. Journal of Econometrics, 117(1):1–19, November 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603001167.

Hosoya:2009:ITS

[HT09] Yuzo Hosoya and Takahiro Terasaka. Inference on transformed stationary time series. Journal of Econometrics, 151(2):129– 139, August 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609000785.

Hordahl:2006:JEM

[HTV06] Peter Hördahl, Oreste Tristani, and David Vestin. A joint econometric model of macroeconomic and term-structure dynamics. Journal of Econometrics, 131(1–2):405–444, March/April 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605000175.

Harvey:2007:TCE

[HTV07] Andrew C. Harvey, Thomas M. Trimbur, and Herman K. Van Dijk. Trends and cycles in economic time series: A Bayesian approach. Journal of Econometrics, 140(2):618–649, October 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606001497.

Hu:2006:BPL

[Hu06] Yingyao Hu. Bounding parameters in a linear regression model with a mismeasured regressor using additional information. Journal of Econometrics, 133(1):51–70, July 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000837.

Hu:2008:IEN

[Hu08] Yingyao Hu. Identification and estimation of nonlinear models with misclassification error using instrumental variables: A general solution. Journal of Econometrics, 144(1):27–61, May 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607002436.

Hsiao:2006:MTS

[HW06] Cheng Hsiao and Siyan Wang. Modified two-stage leastsquares estimators for the estimation of a structural vector autoregressive integrated process. Journal of Econometrics, 135(1–2):427–463, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001703.

Heiss:2008:LAN

[HW08a] Florian Heiss and Viktor Winschel. Likelihood approximation by numerical integration on sparse grids. Journal of Econometrics, 144(1):62–80, May 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607002552.

Hong:2008:RPE

[HW08b] Seung-Hyun Hong and Frank A. Wolak. Relative prices and electronic substitution: Changes in household-level demand for postal delivery services from 1986 to 2004. Journal of Econometrics, 145(1–2):226–242, July 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000596.

Hall:2005:UAT

[HY05] Peter Hall and Adonis Yatchew. Unified approach to testing functional hypotheses in semiparametric contexts. Journal of Econometrics, 127(2):225–252, August 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001472.

Harris:2007:ZIO

[HZ07a] Mark N. Harris and Xueyan Zhao. A zero-inflated ordered probit model, with an application to modelling tobacco consumption. Journal of Econometrics, 141(2):1073–1099, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000048.

Hidalgo:2007:GFTa

[HZ07b] Javier Hidalgo and Paolo Zaffaroni. A goodness-of-fit test for ARCH(∞) models. Journal of Econometrics, 141(2):835–875, December 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606002387.

Hidalgo:2007:GFTb

[HZ07c] Javier Hidalgo and Paolo Zaffaroni. A goodness-of-fit test for ARCH(∞) models. Journal of Econometrics, 141(2):973–1013, December 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606002557.

Inoue:2006:SFM

[IK06] Atsushi Inoue and Lutz Kilian. On the selection of forecasting models. Journal of Econometrics, 130(2):273–306, February 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605000771.

Imbens:2008:SIE

[IL08a] Guido Imbens and Thomas Lemieux. Special issue editors’ introduction: The regression discontinuity design — theory and applications. Journal of Econometrics, 142(2):611–614, February 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607001200.

Imbens:2008:RDD

[IL08b] Guido W. Imbens and Thomas Lemieux. Regression discontinuity designs: A guide to practice. Journal of Econometrics, 142(2):615–635, February 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407607001091.

Issler:2009:PDA

[IL09] João Victor Issler and Luiz Renato Lima. A panel data approach to economic forecasting: The bias-corrected average forecast. Journal of Econometrics, 152(2):153–164, October 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000177.

Im:2000:RGT

[Im00] Kyung So Im. Robustifying glejser test of heteroskedasticity. Journal of Econometrics, 97(1):179–188, July 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000615.

Inkmann:2000:MHP

[Ink00] Joachim Inkmann. Misspecified heteroskedasticity in the panel probit model: A small sample comparison of GMM and SML estimators. Journal of Econometrics, 97(2):227–259, August 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000706.

Inoue:2008:EEI

[Ino08] Atsushi Inoue. Efficient estimation and inference in linear pseudopanel data models. Journal of Econometrics, 142(1):449–466, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607001571.

Im:2003:TUR

[IPS03] Kyung So Im, M. Hashem Pesaran, and Yongcheol Shin. Testing for unit roots in heterogeneous panels. Journal of Econometrics, 115(1):53–74, July 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603000927.

Imbens:2002:CIG

[IS02] Guido W. Imbens and Richard Spady. Confidence intervals in generalized method of moments models. Journal of Econometrics, 107(1–2):87–98, March 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001142.

Inoue:2006:BGE

[IS06] Atsushi Inoue and Mototsugu Shintani. Bootstrapping GMM estimators for time series. Journal of Econometrics, 133(2):531– 555, August 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001235.

Im:2008:MEE

[IS08] Kyung So Im and Peter Schmidt. More efficient estimation under non-normality when higher moments do not depend on the regressors, using residual augmented least squares. Journal of Econometrics, 144(1):219–233, May 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000079.

Issler:2006:MLU

[IV06] João Victor Issler and Farshid Vahid. The missing link: using the NBER recession indicator to construct coincident and leading indices of economic activity. Journal of Econometrics, 132(1): 281–303, May 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605000497.

Imai:2005:BAM

[IvD05] Kosuke Imai and David A. van Dyk. A Bayesian analysis of the multinomial probit model using marginal data augmentation. Journal of Econometrics, 124(2):311–334, February 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000351.

Jansson:2005:POT

[Jan05] Michael Jansson. Point optimal tests of the null hypothesis of cointegration. Journal of Econometrics, 124(1):187–201, January 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604000302.

Johnson:2003:DCT

[JDGZ03] Elizabeth Johnson, Francesca Dominici, Michael Griswold, and Scott L. Zeger. Disease cases and their medical costs attributable to smoking: an analysis of the national medical expenditure survey. Journal of Econometrics, 112(1):135–151, January 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001574.

Jin:2009:DCM

[Jin09] Sainan Jin. Discrete choice modeling with nonstationary panels applied to exchange rate regime choice. Journal of Econometrics, 150(2):312–321, June 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407608002273.

Jacquier:2000:BAC

[JJ00] Eric Jacquier and Robert Jarrow. Bayesian analysis of contingent claim model error. Journal of Econometrics, 94(1–2):145– 180, January 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407699000202.

Jacquier:2007:MML

[JJP07] Eric Jacquier, Michael Johannes, and Nicholas Polson. MCMC maximum likelihood for latent state models. Journal of Econometrics, 137(2):615–640, April 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606000704.

Jagannathan:2003:EMF

[JKS03] Ravi Jagannathan, Andrew Kaplin, and Steve Sun. An evaluation of multi-factor CIR models using LIBOR, swap rates, and cap and swaption prices. Journal of Econometrics, 116(1–2):113–146, September 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603001052.

Jondeau:2008:EBE

[JL08] Eric Jondeau and Hervé Le Bihan. Examining bias in estimators of linear rational expectations models under misspecification. Journal of Econometrics, 143(2):375–395, April 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607002308.

Jacobson:2005:ERC

[JLS05] Louis Jacobson, Robert LaLonde, and Daniel G. Sullivan. Estimating the returns to community college schooling for displaced workers. Journal of Econometrics, 125(1–2):271–304, March 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604000818.

Lee:2006:SPE

[jLV06] Myoung jae Lee and Francis Vella. A semi-parametric estimator for censored selection models with endogeneity. Journal of Econometrics, 130(2):235–252, February 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000539.

Jansen:2008:FPA

[JLWY08] Dennis W. Jansen, Qi Li, Zijun Wang, and Jian Yang. Fiscal policy and asset markets: A semiparametric analysis. Journal of Econometrics, 147(1):141–150, November 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001310.

Judge:2007:EIC

[JM07] George G. Judge and Ron C. Mittelhammer. Estimation and inference in the case of competing sets of estimating equations. Journal of Econometrics, 138(2):513–531, June 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000807.

Jiang:2009:BAR

[JMR09] Renna Jiang, Puneet Manchanda, and Peter E. Rossi. Bayesian analysis of random coefficient logit models using aggregate data. Journal of Econometrics, 149(2):136–148, April 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002297.

Jiang:2008:TJW

[JO08] George J. Jiang and Roel C. A. Oomen. Testing for jumps when asset prices are observed with noise — a “swap variance” approach. Journal of Econometrics, 144(2):352–370, June 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760800033X.

Johansen:2002:SSC

[Joh02] Søren Johansen. A small sample correction for tests of hypotheses on the cointegrating vectors. Journal of Econometrics, 111(2):195–221, December 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602001045.

Johansen:2006:SAH

[Joh06] Søren Johansen. Statistical analysis of hypotheses on the cointegrating relations in the i(2) model. Journal of Econometrics, 132(1):81–115, May 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605000424.

Jones:2003:DSV

[Jon03] Christopher S. Jones. The dynamics of stochastic volatility: evidence from underlying and options markets. Journal of Econometrics, 116(1–2):181–224, September 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001076.

Jenish:2009:CLT

[JP09a] Nazgul Jenish and Ingmar R. Prucha. Central limit theorems and uniform laws of large numbers for arrays of random fields. Journal of Econometrics, 150(1):86–98, May 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000475.

Jun:2009:STC

[JP09b] Sung Jae Jun and Joris Pinkse. Semiparametric tests of conditional moment restrictions under weak or partial identification. Journal of Econometrics, 152(1):3–18, September 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000578.

Jacquier:2004:BAS

[JPR04] Eric Jacquier, Nicholas G. Polson, and Peter E. Rossi. Bayesian analysis of stochastic volatility models with fat-tails and correlated errors. Journal of Econometrics, 122(1):185–212, September 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603002732.

Jalan:2003:DPW

[JR03] Jyotsna Jalan and Martin Ravallion. Does piped water reduce diarrhea for children in rural India? Journal of Econometrics, 112(1):153–173, January 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001586.

Jorgenson:2008:CLS

[JS08] Dale W. Jorgenson and Daniel T. Slesnick. Consumption and labor supply. Journal of Econometrics, 147(2):326–335, December 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608001474.

Jouneau-Sion:2006:MTL

[JST06] Frédéric Jouneau-Sion and Olivier Torrès. MMC techniques for limited dependent variables models: Implementation by the branch-and-bound algorithm. Journal of Econometrics, 133(2):479–512, August 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605001259.

Jun:2008:WIR

[Jun08] Sung Jae Jun. Weak identification robust tests in an instrumental quantile model. Journal of Econometrics, 144(1):118– 138, May 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607002576.

Jun:2009:LSQ

[Jun09] Sung Jae Jun. Local structural quantile effects in a model with a nonseparable control variable. Journal of Econometrics, 151(1):

82–97, July 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609000736.

Jurajda:2002:EEU

[Jur02] Stepán Jurajda. Estimating the effect of unemployment insurance compensation on the labor market histories of displaced workers. Journal of Econometrics, 108(2):227–252, June 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760100135X.

Juhl:2005:NTC

[JX05a] Ted Juhl and Zhijie Xiao. A nonparametric test for changing trends. Journal of Econometrics, 127(2):179–199, August 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001459.

Juhl:2005:TCU

[JX05b] Ted Juhl and Zhijie Xiao. Testing for cointegration using partially linear models. Journal of Econometrics, 124(2):363–394, February 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604000405.

Juhl:2009:TCM

[JX09] Ted Juhl and Zhijie Xiao. Tests for changing mean with monotonic power. Journal of Econometrics, 148(1):14–24, January 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608001176.

Kim:2002:CDC

[KA02] Jae-Young Kim and Rosa Badillo Amador. Corrigendum to “Detection of change in persistence of a linear time series” [J. Econom. 95 (2000) 97–116]. Journal of Econometrics, 109(2):389–392, August 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602000878. See [Kim00].

Karantininis:2002:IBE

[Kar02] Kostas Karantininis. Information-based estimators for the nonstationary transition probability matrix: an application to the Danish pork industry. Journal of Econometrics, 107(1–2):275– 290, March 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601001245.

Kim:2007:PAM

[KAR07] Jaehwan Kim, Greg M. Allenby, and Peter E. Rossi. Product attributes and models of multiple discreteness. Journal of Econometrics, 138(1):208–230, May 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000935.

Kawakatsu:2006:MEG

[Kaw06] Hiroyuki Kawakatsu. Matrix exponential GARCH. Journal of Econometrics, 134(1):95–128, September 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001399.

Kemp:2001:IWT

[Kem01] Gordon C. R. Kemp. Invariance and the Wald test. Journal of Econometrics, 104(2):209–217, September 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000483.

Kuan:2009:GEI

[KH09a] Chung-Ming Kuan and Yongmiao Hong. Guest editors’ introduction. Journal of Econometrics, 150(2):117–118, June 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002121.

Kurozumi:2009:APE

[KH09b] Eiji Kurozumi and Kazuhiko Hayakawa. Asymptotic properties of the efficient estimators for cointegrating regression models with serially dependent errors. Journal of Econometrics, 149(2):118– 135, April 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760800211X.

Khan:2001:TSR

[Kha01] Shakeeb Khan. Two-stage rank estimation of quantile index models. Journal of Econometrics, 100(2):319–355, February 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000403.

Kim:2000:DCP

[Kim00] Jae-Young Kim. Detection of change in persistence of a linear time series. Journal of Econometrics, 95(1):97–116, March 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000317. See corrigendum [KA02].

Kim:2002:LIL

[Kim02] Jae-Young Kim. Limited information likelihood and Bayesian analysis. Journal of Econometrics, 107(1–2):175–193, March 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601001191.

Kim:2004:MSM

[Kim04] Chang-Jin Kim. Markov-switching models with endogenous explanatory variables. Journal of Econometrics, 122(1):127–136, September 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603002719.

Kim:2008:MSB

[Kim08] Chang-Jin Kim. Markov-switching and the Beveridge–Nelson decomposition: Has US output persistence changed since 1984? Journal of Econometrics, 146(2):227–240, October 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001048.

Kim:2009:MSM

[Kim09] Chang-Jin Kim. Markov-switching models with endogenous explanatory variables II: A two-step MLE procedure. Journal of Econometrics, 148(1):46–55, January 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760800153X.

Kapoor:2007:PDM

[KKP07] Mudit Kapoor, Harry H. Kelejian, and Ingmar R. Prucha. Panel data models with spatially correlated error components. Journal of Econometrics, 140(1):97–130, September 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002259.

Kalnina:2008:EQV

[KL08] Ilze Kalnina and Oliver Linton. Estimating quadratic variation consistently in the presence of endogenous and diurnal measurement error. Journal of Econometrics, 147(1):47–59, November 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608001231.

Kleibergen:2004:IBI

[Kle04] Frank Kleibergen. Invariant Bayesian inference in regression models that is robust against the Jeffreys–Lindley’s paradox. Journal of Econometrics, 123(2):227–258, December 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002902.

Kleibergen:2007:GWI

[Kle07] Frank Kleibergen. Generalizing weak instrument robust IV statistics towards multiple parameters, unrestricted covariance matrices and identification statistics. Journal of Econometrics, 139(1): 181–216, July 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407606001084.

Kleibergen:2009:TRP

[Kle09] Frank Kleibergen. Tests of risk premia in linear factor models. Journal of Econometrics, 149(2):149–173, April 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000311.

Koopman:2008:MSL

[KLM08] Siem Jan Koopman, André Lucas, and André Monteiro. The multi-state latent factor intensity model for credit rating transitions. Journal of Econometrics, 142(1):399–424, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607001467.

Kim:2002:URT

[KLN02] Tae-Hwan Kim, Stephen Leybourne, and Paul Newbold. Unit root tests with a break in innovation variance. Journal of Econometrics, 109(2):365–387, August 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001185.

Kazakevicius:2004:SRC

[KLV04] Vytautas Kazakevicius, Remigijus Leipus, and Marie-Claude Viano. Stability of random coefficient ARCH models and aggregation schemes. Journal of Econometrics, 120(1):139–158, May 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603002094.

Klaassen:2009:ETA

[KM09] Franc J. G. M. Klaassen and Jan R. Magnus. The efficiency of top agents: An analysis through service strategy in tennis. Journal of Econometrics, 148(1):72–85, January 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001644.

Korinek:2007:EMC

[KMR07] Anton Korinek, Johan A. Mistiaen, and Martin Ravallion. An econometric method of correcting for unit nonresponse bias in surveys. Journal of Econometrics, 136(1):213–235, January 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606000327.

Koenker:2000:GEF

[Koe00] Roger Koenker. Galton, Edgeworth, frisch, and prospects for quantile regression in econometrics. Journal of Econometrics, 95(2):347–374, April 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407699000433.

Komunjer:2005:QML

[Kom05] Ivana Komunjer. Quasi-maximum likelihood estimation for conditional quantiles. Journal of Econometrics, 128(1):137–164, September 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604001526.

Kongsted:2005:TNR

[Kon05] Hans Christian Kongsted. Testing the nominal-to-real transformation. Journal of Econometrics, 124(2):205–225, February 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604000314.

Koop:2001:BIM

[Koo01] Gary Koop. Bayesian inference in models based on equilibrium search theory. Journal of Econometrics, 102(2):311–338, June 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601000562.

Kelejian:2001:ADM

[KP01a] Harry H. Kelejian and Ingmar R. Prucha. On the asymptotic distribution of the Moran I test statistic with applications. Journal of Econometrics, 104(2):219–257, September 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000641.

Khan:2001:TSE

[KP01b] Shakeeb Khan and James L. Powell. Two-step estimation of semiparametric censored regression models. Journal of Econometrics, 103(1–2):73–110, July 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000409.

Kleibergen:2002:PPB

[KP02] Frank Kleibergen and Richard Paap. Priors, posteriors and Bayes factors for a Bayesian analysis of cointegration. Journal of Econometrics, 111(2):223–249, December 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001057.

Kelejian:2004:ESS

[KP04a] Harry H. Kelejian and Ingmar R. Prucha. Estimation of simultaneous systems of spatially interrelated cross sectional equations. Journal of Econometrics, 118(1–2):27–50, January 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001337.

Koop:2004:BVS

[KP04b] Gary Koop and Dale J. Poirier. Bayesian variants of some classical semiparametric regression techniques. Journal of Econometrics, 123(2):259–282, December 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002914.

Kleibergen:2006:GRR

[KP06] Frank Kleibergen and Richard Paap. Generalized reduced rank tests using the singular value decomposition. Journal of Econometrics, 133(1):97–126, July 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000850.

Kelejian:2007:HES

[KP07] Harry H. Kelejian and Ingmar R. Prucha. HAC estimation in a spatial framework. Journal of Econometrics, 140(1):131–154, September 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606002260.

Kejriwal:2008:LDE

[KP08] Mohitosh Kejriwal and Pierre Perron. The limit distribution of the estimates in cointegrated regression models with multiple structural changes. Journal of Econometrics, 146(1):59–73, September 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000754.

Kim:2009:ARP

[KP09a] Dukpa Kim and Pierre Perron. Assessing the relative power of structural break tests using a framework based on the approximate Bahadur slope. Journal of Econometrics, 149(1):26–51, April 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608002078.

Kim:2009:URT

[KP09b] Dukpa Kim and Pierre Perron. Unit root tests allowing for a break in the trend function at an unknown time under both the null and alternative hypotheses. Journal of Econometrics, 148(1): 1–13, January 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000961.

Kapetanios:2007:MMC

[KPS07a] G. Kapetanios, A. Pagan, and A. Scott. Making a match: Combining theory and evidence in policy-oriented macroeconomic modeling. Journal of Econometrics, 136(2):565–594, February 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605002253.

Kuosmanen:2007:NPT

[KPS07b] Timo Kuosmanen, Thierry Post, and Stefan Scholtes. Nonparametric tests of productive efficiency with errors-in-variables. Journal of Econometrics, 136(1):131–162, January 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002046.

Kim:2008:EMR

[KPS08] Chang-Jin Kim, Jeremy Piger, and Richard Startz. Estimation of Markov regime-switching regression models with endogenous switching. Journal of Econometrics, 143(2):263–273, April 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607002126.

Kumbhakar:2007:NSF

[KPST07] Subal C. Kumbhakar, Byeong U. Park, Léopold Simar, and Efthymios G. Tsionas. Nonparametric stochastic frontiers: A local maximum likelihood approach. Journal of Econometrics, 137(1): 1–27, March 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000376.

Krauth:2006:SBE

[Kra06] Brian V. Krauth. Simulation-based estimation of peer effects. Journal of Econometrics, 133(1):243–271, July 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000904.

Krishnakumar:2001:SCJ

[Kri01] Jaya Krishnakumar. A short comment on the JE open forum essays. Journal of Econometrics, 100(1):77–78, January 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000609.

Kristensen:2008:EPD

[Kri08] Dennis Kristensen. Estimation of partial differential equations with applications in finance. Journal of Econometrics, 144(2): 392–408, June 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000353.

Kruiniger:2008:MLE

[Kru08] Hugo Kruiniger. Maximum likelihood estimation and inference methods for the covariance stationary panel AR(1)/unit root model. Journal of Econometrics, 144(2):447–464, June 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0304407608000390. See corrigendum [Kru14].

Kruiniger:2014:CML

[Kru14] Hugo Kruiniger. Corrigendum to “Maximum likelihood estimation and inference methods for the covariance stationary panel AR(1) /unit root model” [J. Econom. 144 (2008) 447– 464]. Journal of Econometrics, 178(2):824, February 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407613002340. See [Kru08].

Kitamura:2002:CBE

[KS02] Yuichi Kitamura and Michael Stutzer. Connections between entropic and linear projections in asset pricing estimation. Journal of Econometrics, 107(1–2):159–174, March 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760100118X.

Keller:2007:OSI

[KS07] Wolfgang Keller and Carol H. Shiue. The origin of spatial interaction. Journal of Econometrics, 140(1):304–332, September 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606002338.

Kasahara:2008:PLE

[KS08a] Hiroyuki Kasahara and Katsumi Shimotsu. Pseudo-likelihood estimation and bootstrap inference for structural discrete Markov decision models. Journal of Econometrics, 146(1):92–106, September 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760800078X.

Kim:2008:VTW

[KS08b] Myungsup Kim and Peter Schmidt. Valid tests of whether technical inefficiency depends on firm characteristics. Journal of Econometrics, 144(2):409–427, June 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000365.

Koopman:2009:TAB

[KSC09] Siem Jan Koopman, Neil Shephard, and Drew Creal. Testing the assumptions behind importance sampling. Journal of Econometrics, 149(1):2–11, April 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001693.

Kapetanios:2003:TUR

[KSS03] George Kapetanios, Yongcheol Shin, and Andy Snell. Testing for a unit root in the nonlinear STAR framework. Journal of Econometrics, 112(2):359–379, February 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602002026.

Kumbhakar:2005:MTA

[KT05] Subal C. Kumbhakar and Efthymios G. Tsionas. Measuring technical and allocative inefficiency in the translog cost system: a Bayesian approach. Journal of Econometrics, 126(2):355–384, June 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001162.

Koop:2006:SBI

[KT06a] Gary Koop and Justin L. Tobias. Semiparametric Bayesian inference in smooth coefficient models. Journal of Econometrics, 134(1):283–315, September 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605001491.

Kumbhakar:2006:ESF

[KT06b] Subal C. Kumbhakar and Efthymios G. Tsionas. Estimation of stochastic frontier production functions with input-oriented technical efficiency. Journal of Econometrics, 133(1):71–96, July 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605000849.

Khan:2007:PRE

[KT07] Shakeeb Khan and Elie Tamer. Partial rank estimation of duration models with general forms of censoring. Journal of Econometrics, 136(1):251–280, January 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000340.

Khan:2009:IEC

[KT09] Shakeeb Khan and Elie Tamer. Inference on endogenously censored regression models using conditional moment inequalities. Journal of Econometrics, 152(2):104–119, October 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000219.

Kuersteiner:2001:OIV

[Kue01] Guido M. Kuersteiner. Optimal instrumental variables estimation for ARMA models. Journal of Econometrics, 104(2):359–405, September 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000884.

Kurozumi:2002:TSB

[Kur02] Eiji Kurozumi. Testing for stationarity with a break. Journal of Econometrics, 108(1):63–99, May 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001063.

Koop:2000:TIU

[KV00] Gary Koop and Herman K. Van Dijk. Testing for integration using evolving trend and seasonals models: A Bayesian approach. Journal of Econometrics, 97(2):261–291, August 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000718.

Kumbhakar:2006:ETA

[KW06a] Subal C. Kumbhakar and Hung-Jen Wang. Estimation of technical and allocative inefficiency: A primal system approach. Journal of Econometrics, 134(2):419–440, October 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760500151X.

Kumbhakar:2006:PEC

[KW06b] Subal C. Kumbhakar and Hung-Jen Wang. Pitfalls in the estimation of a cost function that ignores allocative inefficiency: A Monte Carlo analysis. Journal of Econometrics, 134(2):317– 340, October 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001466.

Koenker:2009:PLB

[KY09] Roger Koenker and Jungmo Yoon. Parametric links for binary choice models: A fisherian-Bayesian colloquy. Journal of Econometrics, 152(2):120–130, October 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000207.

Kuan:2009:AVR

[KYH09] Chung-Ming Kuan, Jin-Huei Yeh, and Yu-Chin Hsu. Assessing value at risk with CARE, the conditional autoregressive expectile models. Journal of Econometrics, 150(2):261–270, June 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608002236.

Kleibergen:2003:BCA

[KZ03]  Frank Kleibergen and Eric Zivot.  Bayesian and classi-

cal approaches to instrumental variable regression. Journal of Econometrics, 114(1):29–72, May 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602002191.

LaFrance:2008:SUF

[LaF08] Jeffrey T. LaFrance. The structure of US food demand. Journal of Econometrics, 147(2):336–349, December 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001486.

Lalive:2008:HDE

[Lal08] Rafael Lalive. How do extended benefits affect unemployment duration? A regression discontinuity approach. Journal of Econometrics, 142(2):785–806, February 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001170.

Lamont:2001:ETP

[Lam01] Owen A. Lamont. Economic tracking portfolios. Journal of Econometrics, 105(1):161–184, November 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000744.

Lancaster:2000:IPP

[Lan00] Tony Lancaster. The incidental parameter problem since 1948. Journal of Econometrics, 95(2):391–413, April 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000445.

Lavergne:2001:ETA

[Lav01] Pascal Lavergne. An equality test across nonparametric regressions. Journal of Econometrics, 103(1–2):307–344, July 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760100046X.

Lazarova:2005:TSC

[Laz05] Stepána Lazarová. Testing for structural change in regression with long memory processes. Journal of Econometrics, 129(1–2):329– 372, November 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604001733.

LaFrance:2002:ITM

[LBPA02] J. T. LaFrance, T. K. M. Beatty, R. D. Pope, and G. K. Agnew. Information theoretic measures of the income distribution in food demand. Journal of Econometrics, 107(1–2):235– 257, March 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601001221.

Lee:2008:RDI

[LC08] David S. Lee and David Card. Regression discontinuity inference with specification error. Journal of Econometrics, 142(2):655– 674, February 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760700111X.

Lee:2007:EQR

[Lee07] Sokbae Lee. Endogeneity in quantile regression models: A control function approach. Journal of Econometrics, 141(2):1131–1158, December 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607000279.

Lee:2008:REN

[Lee08] David S. Lee. Randomized experiments from non-random selection in U.S. House elections. Journal of Econometrics, 142(2):675– 697, February 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001121.

Lehmann:2009:RBI

[Leh09] Bruce N. Lehmann. The role of beliefs in inference for rational expectations models. Journal of Econometrics, 150(2):322– 331, June 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608002285.

Lewbel:2000:SQR

[Lew00] Arthur Lewbel. Semiparametric qualitative response model estimation with unknown heteroscedasticity or instrumental variables. Journal of Econometrics, 97(1):145–177, July 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000154.

Lewbel:2003:CCC

[Lew03] Arthur Lewbel. Calculating compensation in cases of wrongful death. Journal of Econometrics, 113(1):115–128, March 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602001690.

Lewbel:2007:EST

[Lew07] Arthur Lewbel. Endogenous selection or treatment model estimation. Journal of Econometrics, 141(2):777–806, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606002363.

Lahiri:2002:BAN

[LG02] Kajal Lahiri and Jian Gao. Bayesian analysis of nested logit model by Markov chain Monte Carlo. Journal of Econometrics, 111(1):103–133, November 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001252.

Li:2004:REG

[LH04] Tong Li and Cheng Hsiao. Robust estimation of generalized linear models with measurement errors. Journal of Econometrics, 118(1–2):51–65, January 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001349.

Liu:2009:MFT

[LHMC09] Echu Liu, Cheng Hsiao, Tomoya Matsumoto, and Shinyi Chou. Maternal full-time employment and overweight children: Parametric, semi-parametric, and non-parametric assessment. Journal of Econometrics, 152(1):61–69, September 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000542.

Li:2003:CST

[LHZ03] Qi Li, Cheng Hsiao, and Joel Zinn. Consistent specification tests for semiparametric/nonparametric models based on series estimation methods. Journal of Econometrics, 112(2):295–325, February 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602001987.

Li:2002:RCE

[Li02] Tong Li. Robust and consistent estimation of nonlinear errors-in-variables models. Journal of Econometrics, 110(1):1–26, September 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602001203.

Li:2005:EFP

[Li05] Tong Li. Econometrics of first-price auctions with entry and binding reservation prices. Journal of Econometrics, 126(1):173–200, May 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001344.

Li:2008:ETE

[Li08] Hong Li. Estimation and testing of Euler equation models with time-varying reduced-form coefficients. Journal of Econometrics, 142(1):425–448, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407607001479.

Li:2009:SBS

[Li09] Tong Li. Simulation based selection of competing structural econometric models. Journal of Econometrics, 148(2):114–123, February 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001681.

Liesenfeld:2001:GBM

[Lie01] Roman Liesenfeld. A generalized bivariate mixture model for stock price volatility and trading volume. Journal of Econometrics, 104(1):141–178, August 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407601000628.

Linton:2002:EAS

[Lin02] Oliver Linton. Edgeworth approximations for semiparametric instrumental variable estimators and test statistics. Journal of Econometrics, 106(2):325–368, February 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760100104X.

Ling:2007:SWL

[Lin07] Shiqing Ling. Self-weighted and local quasi-maximum likelihood estimators for ARMA-GARCH/IGARCH models. Journal of Econometrics, 140(2):849–873, October 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001588.

Liu:2000:MLM

[Liu00] Ming Liu. Modeling long memory in stock market volatility. Journal of Econometrics, 99(1):139–171, November 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000336.

Lee:2009:CBM

[LL09] Tae-Hwy Lee and Xiangdong Long. Copula-based multivariate GARCH model with uncorrelated dependent errors. Journal of Econometrics, 150(2):207–218, June 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002194.

Levin:2002:URT

[LLC02] Andrew Levin, Chien-Fu Lin, and Chia-Shang James Chu. Unit root tests in panel data: asymptotic and finite-sample properties. Journal of Econometrics, 108(1):1–24, May 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000987.

Ling:2002:SEM

[LM02] Shiqing Ling and Michael McAleer. Stationarity and the existence of moments of a family of GARCH processes. Journal of Econometrics, 106(1):109–117, January 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000902.

Lemieux:2008:IES

[LM08a] Thomas Lemieux and Kevin Milligan. Incentive effects of social assistance: A regression discontinuity approach. Journal of Econometrics, 142(2):807–828, February 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001182.

Linton:2008:NTW

[LM08b] Oliver B. Linton and Enno Mammen. Nonparametric transformation to white noise. Journal of Econometrics, 142(1):241– 264, January 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001297.

Linton:2001:YCE

[LMNT01] Oliver Linton, Enno Mammen, Jans Perch Nielsen, and Carsten Tanggaard. Yield curve estimation by kernel smoothing methods. Journal of Econometrics, 105(1):185–223, November 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000756.

Lima:2009:NRM

[LMPX09] Luiz Renato Lima, Marcelo Moreira, Jack Porter, and Zhijie Xiao. Nonparametric and robust methods in econometrics. Journal of Econometrics, 152(2):79–80, October 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000992.

Li:2009:NTE

[LMR09] Qi Li, Esfandiar Maasoumi, and Jeffrey S. Racine. A nonparametric test for equality of distributions with mixed categorical and continuous data. Journal of Econometrics, 148(2):186–200, February 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608002054.

Long:2004:HCD

[Lon04a] Bridget Terry Long. How have college decisions changed over time? An application of the conditional logistic choice model. Journal of Econometrics, 121(1–2):271–296, July 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002549.

Long:2004:CAE

[Lon04b] Mark C. Long. College applications and the effect of affirmative action. Journal of Econometrics, 121(1–2):319–342, July 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603002513.

Li:2003:EMB

[LP03] Kai Li and Dale J. Poirier. An econometric model of birth inputs and outputs for native Americans. Journal of Econometrics, 113(2):337–361, April 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602002063.

Lubrano:2004:DIR

[LP04] Michel Lubrano and Camelia Protopopescu. Density inference for ranking European research systems in the field of economics. Journal of Econometrics, 123(2):345–369, December 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002926.

LeSage:2007:MES

[LP07] James P. LeSage and R. Kelley Pace. A matrix exponential spatial specification. Journal of Econometrics, 140(1):190–214, September 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606002284.

Lavergne:2008:BCD

[LP08a] Pascal Lavergne and Valentin Patilea. Breaking the curse of dimensionality in nonparametric testing. Journal of Econometrics, 143(1):103–122, March 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407607001601.

Leeb:2008:SEO

[LP08b] Hannes Leeb and Benedikt M. Pötscher. Sparse estimators and the oracle property, or the return of Hodges’ estimator. Journal of Econometrics, 142(1):201–211, January 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001273.

Lewbel:2008:ECH

[LP08c] Arthur Lewbel and Krishna Pendakur. Estimation of collective household models with Engel curves. Journal of Econometrics, 147(2):350–358, December 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001498.

Lieberman:2008:CAS

[LP08d] Offer Lieberman and Peter C. B. Phillips. A complete asymptotic series for the autocovariance function of a long memory process. Journal of Econometrics, 147(1):99–103, November 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001279.

Leipus:2005:RRS

[LPS05] Remigijus Leipus, Vygantas Paulauskas, and Donatas Surgailis. Renewal regime switching and stable limit laws. Journal of Econometrics, 129(1–2):299–327, November 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001721.

Li:2000:CIP

[LPV00] Tong Li, Isabelle Perrigne, and Quang Vuong. Conditionally independent private information in OCS wildcat auctions. Journal of Econometrics, 98(1):129–161, September 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000810.

Lutkepohl:2000:TCR

[LS00] Helmut Lütkepohl and Pentti Saikkonen. Testing for the cointegrating rank of a VAR process with a time trend. Journal of Econometrics, 95(1):177–198, March 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000354.

Lewbel:2007:SOD

[LS07] Arthur Lewbel and Susanne M. Schennach. A simple ordered data estimator for inverse density weighted expectations. Journal of Econometrics, 136(1):189–211, January 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760500206X.

Lahiri:2008:EFD

[LS08a] Kajal Lahiri and Xuguang Sheng. Evolution of forecast disagreement in a Bayesian learning model. Journal of Econometrics, 144(2):325–340, June 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407608000213.

Lee:2008:SEB

[LS08b] Sokbae Lee and Myung Hwan Seo. Semiparametric estimation of a binary response model with a change-point due to a covariate threshold. Journal of Econometrics, 144(2):492–499, June 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608000420.

Lawford:2009:FSE

[LS09a] Steve Lawford and Michalis P. Stamatogiannis. The finite-sample effects of VAR dimensions on OLS bias, OLS variance, and minimum MSE estimators. Journal of Econometrics, 148(2):124– 130, February 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001802.

Linton:2009:CEG

[LS09b] Oliver Linton and Alessio Sancetta. Consistent estimation of a general nonparametric regression function in time series. Journal of Econometrics, 152(1):70–78, September 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000554.

Lutkepohl:2003:CTC

[LST03] Helmut Lütkepohl, Pentti Saikkonen, and Carsten Trenkler. Comparison of tests for the cointegrating rank of a VAR process with a structural shift. Journal of Econometrics, 113(2):201–229, April 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602002002.

Lahiri:2008:MSS

[LSW08] Kajal Lahiri, Jae Song, and Bernard Wixon. A model of social security disability insurance using matched SIPP/Administrative data. Journal of Econometrics, 145(1–2):4–20, July 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000468.

Lundbergh:2002:EGM

[LT02] Stefan Lundbergh and Timo Teräsvirta. Evaluating GARCH models. Journal of Econometrics, 110(2):417–435, October 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602000969.

Li:2006:CBT

[LT06a] Fuchun Li and Greg Tkacz. A consistent bootstrap test for conditional density functions with time-series data. Journal of Econometrics, 133(2):863–886, August 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001363.

Lundbergh:2006:TSM

[LT06b] Stefan Lundbergh and Timo Teräsvirta. A time series model for an exchange rate in a target zone with applications. Journal of Econometrics, 131(1–2):579–609, March/April 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000229.

Luger:2003:ENP

[Lug03] Richard Luger. Exact non-parametric tests for a random walk with unknown drift under conditional heteroscedasticity. Journal of Econometrics, 115(2):259–276, August 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603000976.

Luger:2006:EPT

[Lug06] Richard Luger. Exact permutation tests for non-nested nonlinear regression models. Journal of Econometrics, 133(2):513– 529, August 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001247.

Lutkepohl:2001:CEC

[Lüt01] Helmut Lütkepohl. Comment on essays on current state and future challenges of econometrics. Journal of Econometrics, 100(1): 81–82, January 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000622.

Lutkepohl:2007:GSS

[Lüt07] Helmut Lütkepohl. General-to-specific or specific-to-general modelling? An opinion on current econometric terminology. Journal of Econometrics, 136(1):319–324, January 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002307.

Lenk:2001:BER

[LW01] Peter Lenk and Michel Wedel. Bayesian econometrics:: A reaction to Geweke. Journal of Econometrics, 100(1):79–80, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000610. See [Gew01, Zel01].

Linton:2007:QAE

[LW07] O. Linton and Yoon-Jae Whang. The quantilogram: With an application to evaluating directional predictability. Journal of Econometrics, 141(1):250–282, November 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000152.

Lee:2006:BBQ

[LY06] Tae-Hwy Lee and Yang Yang. Bagging binary and quantile predictors for time series. Journal of Econometrics, 135(1–2): 465–497, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605001685.

Maasoumi:2001:RFO

[Maa01] Esfandiar Maasoumi. On the relevance of first-order asymptotic theory to economics. Journal of Econometrics, 100(1):83–86, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000634.

Machado:2004:CEB

[Mac04] Matilde P. Machado. A consistent estimator for the binomial distribution in the presence of “incidental parameters”: an application to patent data. Journal of Econometrics, 119(1):73– 98, March 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603001568.

Malani:2008:PEM

[Mal08] Anup Malani. Patient enrollment in medical trials: Selection bias in a randomized experiment. Journal of Econometrics, 144(2): 341–351, June 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000328.

Manski:2000:IPD

[Man00] Charles F. Manski. Identification problems and decisions under ambiguity: Empirical analysis of treatment response and normative analysis of treatment choice. Journal of Econometrics, 95(2): 415–442, April 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407699000457.

Manski:2007:MRT

[Man07] Charles F. Manski. Minimax-regret treatment choice with missing outcome data. Journal of Econometrics, 139(1):105–115, July 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407606001047.

Marsh:2004:EHE

[Mar04] Lawrence C. Marsh. The econometrics of higher education: editor’s view. Journal of Econometrics, 121(1–2):1–18, July 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603002574.

Marmer:2008:NNS

[Mar08] Vadim Marmer. Nonlinearity, nonstationarity, and spurious forecasts. Journal of Econometrics, 142(1):1–27, January 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760700108X.

Matsudaira:2008:MSS

[Mat08] Jordan D. Matsudaira. Mandatory summer school and student achievement. Journal of Econometrics, 142(2):829–850, February 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607001194.

McCrorie:2006:GCS

[MC06] J. Roderick McCrorie and Marcus J. Chambers. Granger causality and the sampling of economic processes. Journal of Econometrics, 132(2):311–336, June 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000576.

Millimet:2008:EPS

[MC08] Daniel L. Millimet and Trevor Collier. Efficiency in public schools: Does competition matter? Journal of Econometrics, 145(1–2): 134–157, July 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760800047X.

McAleer:2007:EIP

[McA07] Michael McAleer. The econometrics of intellectual property: An overview. Journal of Econometrics, 139(2):237–241, August 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606002120.

McAdams:2008:PIT

[McA08] David McAdams. Partial identification and testable restrictions in multi-unit auctions. Journal of Econometrics, 146(1):74–85, September 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000766.

McCracken:2000:RSI

[McC00] Michael W. McCracken. Robust out-of-sample inference. Journal of Econometrics, 99(2):195–223, December 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000221.

McCausland:2007:TRS

[McC07a] William J. McCausland. Time reversibility of stationary regular finite-state Markov chains. Journal of Econometrics, 136(1):303– 318, January 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000534.

McCracken:2007:AST

[McC07b] Michael W. McCracken. Asymptotics for out of sample tests of Granger causality. Journal of Econometrics, 140(2):719–752, October 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606001539.

McCausland:2008:BAC

[McC08a] William J. McCausland. On Bayesian analysis and computation for functions with monotonicity and curvature restrictions. Journal of Econometrics, 142(1):484–507, January 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001728.

McCrary:2008:MRV

[McC08b] Justin McCrary. Manipulation of the running variable in the regression discontinuity design: A density test. Journal of Econometrics, 142(2):698–714, February 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001133.

McKenzie:2004:ATH

[McK04] David J. McKenzie. Asymptotic theory for heterogeneous dynamic pseudo-panels. Journal of Econometrics, 120(2):235–262, June 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603002136.

McAleer:2007:EAA

[MCM07] Michael McAleer, Felix Chan, and Dora Marinova. An econometric analysis of asymmetric volatility: Theory and application to patents. Journal of Econometrics, 139(2):259–284, August 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606002144.

Martins-Filho:2007:NFE

[MFY07] Carlos Martins-Filho and Feng Yao. Nonparametric frontier estimation via local linear regression. Journal of Econometrics, 141(1):283–319, November 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407607000164.

Martins-Filho:2008:SNC

[MFY08] Carlos Martins-Filho and Feng Yao. A smooth nonparametric conditional quantile frontier estimator. Journal of Econometrics, 143(2):317–333, April 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407607002151.

Matilla-Garcia:2008:NPI

[MGM08] Mariano Matilla-García and Manuel Ruiz Marín. A nonparametric independence test using permutation entropy. Journal of Econometrics, 144(1):139–155, May 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760800002X.

Miller:2002:RJD

[MhL02] Douglas J. Miller and Wei han Liu. On the recovery of joint distributions from limited information. Journal of Econometrics, 107(1–2):259–274, March 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001233.

Mittelhammer:2005:CEI

[MJ05] Ron C. Mittelhammer and George G. Judge. Combining estimators to improve structural model estimation and inference under quadratic loss. Journal of Econometrics, 128(1):1–29, September 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001484.

Ma:2006:QRM

[MK06] Lingjie Ma and Roger Koenker. Quantile regression methods for recursive structural equation models. Journal of Econometrics, 134(2):471–506, October 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605001533.

Magnac:2007:IIM

[MM07] Thierry Magnac and Eric Maurin. Identification and information in monotone binary models. Journal of Econometrics, 139(1): 76–104, July 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407606001035.

McAleer:2008:MRS

[MM08] Michael McAleer and Marcelo C. Medeiros. A multiple regime smooth transition heterogeneous autoregressive model for long memory and asymmetries. Journal of Econometrics, 147(1):104– 119, November 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001280.

Mora:2008:STO

[MME08] Juan Mora and Ana I. Moro-Egido. On specification testing of ordered discrete choice models. Journal of Econometrics, 143(1): 191–205, March 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001637.

McAleer:2008:NND

[MMS08a] Michael McAleer, Marcelo C. Medeiros, and Daniel Slottje. A neural network demand system with heteroskedastic errors. Journal of Econometrics, 147(2):359–371, December 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001504.

Medeiros:2008:AAE

[MMS^{+}08b] Marcelo C. Medeiros, Michael McAleer, Daniel Slottje, Vicente Ramos, and Javier Rey-Maquieira. An alternative approach to estimating demand: Neural network regression with conditional volatility for high frequency air passenger arrivals. Journal of Econometrics, 147(2):372–383, December 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001516.

Marriott:2000:SEU

[MN00a] John Marriott and Paul Newbold. The strength of evidence for unit autoregressive roots and structural breaks: A Bayesian perspective. Journal of Econometrics, 98(1):1–25, September 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000779.

Michaelides:2000:ERE

[MN00b] Alexander Michaelides and Serena Ng. Estimating the rational expectations model of speculative storage: A Monte Carlo comparison of three simulation estimators. Journal of Econometrics, 96(2):231–266, June 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000585.

Millimet:2003:EWE

[MNRS03] Daniel L. Millimet, Michael Nieswiadomy, Hang Ryu, and Daniel Slottje. Estimating worklife expectancy: an econometric approach. Journal of Econometrics, 113(1):83–113, March 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602001689.

Montanes:2005:SBP

[MOC05] Antonio Montañés, Irene Olloqui, and Elena Calvo. Selection of the break in the Perron-type tests. Journal of Econometrics, 129(1–2):41–64, November 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001654.

Moench:2008:FYC

[Moe08] Emanuel Moench. Forecasting the yield curve in a data-rich environment: A no-arbitrage factor-augmented VAR approach. Journal of Econometrics, 146(1):26–43, September 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000730.

Molinari:2008:PIP

[Mol08] Francesca Molinari. Partial identification of probability distributions with misclassified data. Journal of Econometrics, 144(1): 81–117, May 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607002564.

Moon:2004:MSE

[Moo04] Hyungsik Roger Moon. Maximum score estimation of a nonstationary binary choice model. Journal of Econometrics, 122(2):385–403, October 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603002811.

Moretti:2004:ESR

[Mor04] Enrico Moretti. Estimating the social return to higher education: evidence from longitudinal and repeated cross-sectional data. Journal of Econometrics, 121(1–2):175–212, July 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002653.

Moreira:2009:TCS

[Mor09] Marcelo J. Moreira. Tests with correct size when instruments can be arbitrarily weak. Journal of Econometrics, 152(2):131– 140, October 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609000190.

Moschini:2001:PRE

[Mos01] GianCarlo Moschini. Production risk and the estimation of exante cost functions. Journal of Econometrics, 100(2):357–380, February 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000415.

Moon:2004:TUR

[MP04] Hyungsik Roger Moon and Benoît Perron. Testing for a unit root in panels with dynamic factors. Journal of Econometrics, 122(1):81–126, September 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603002707.

Morley:2008:TCD

[MP08] James Morley and Jeremy Piger. Trend/cycle decomposition of regime-switching processes. Journal of Econometrics, 146(2):220– 226, October 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001036.

Moon:2007:ITP

[MPP07] Hyungsik Roger Moon, Benoît Perron, and Peter C. B. Phillips. Incidental trends and the power of panel unit root tests. Journal of Econometrics, 141(2):416–459, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002016.

McCulloch:2000:BAM

[MPR00] Robert E. McCulloch, Nicholas G. Polson, and Peter E. Rossi. A Bayesian analysis of the multinomial probit model with fully identified parameters. Journal of Econometrics, 99(1):173–193, November 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000348.

Mittnik:2002:SSP

[MPR02] Stefan Mittnik, Marc S. Paolella, and Svetlozar T. Rachev. Stationarity of stable power-GARCH processes. Journal of Econometrics, 106(1):97–107, January 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000896.

Moreira:2009:BVS

[MPS09] Marcelo J. Moreira, Jack R. Porter, and Gustavo A. Suarez. Bootstrap validity for the score test when instruments may be weak. Journal of Econometrics, 149(1):52–64, April 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002030.

McCulloch:2000:RN

[MR00] Robert E. McCulloch and Peter E. Rossi. Reply to nobile. Journal of Econometrics, 99(2):347–348, December 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000361.

Marinucci:2001:SFC

[MR01] D. Marinucci and P. M. Robinson. Semiparametric fractional cointegration analysis. Journal of Econometrics, 105(1):225–247, November 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000768.

Maasoumi:2002:EPS

[MR02] Esfandiar Maasoumi and Jeff Racine. Entropy and predictability of stock market returns. Journal of Econometrics, 107(1–2):291– 312, March 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601001257.

Mealli:2003:AAE

[MR03] Fabrizia Mealli and Donald B. Rubin. Assumptions allowing the estimation of direct causal effects. Journal of Econometrics, 112(1):79–87, January 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602001501.

Meddahi:2004:TAV

[MR04] Nour Meddahi and Eric Renault. Temporal aggregation of volatility models. Journal of Econometrics, 119(2):355–379, April 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603002008.

Maasoumi:2007:GCP

[MRS07] Esfandiar Maasoumi, Jeff Racine, and Thanasis Stengos. Growth and convergence: A profile of distribution dynamics and mobility. Journal of Econometrics, 136(2):483–508, February 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002228.

Machado:2000:GTR

[MS00] José A. F. Machado and J. M. C. Santos Silva. Glejser’s test revisited. Journal of Econometrics, 97(1):189–202, July 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000166.

Mouchart:2004:BEN

[MS04] Michel Mouchart and Eliana Scheihing. Bayesian evaluation of non-admissible conditioning. Journal of Econometrics, 123(2):283–306, December 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603002938.

Mosconi:2006:NCB

[MS06] Rocco Mosconi and Raffaello Seri. Non-causality in bivariate binary time series. Journal of Econometrics, 132(2):379–407, June 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605000606.

Mencia:2009:MLS

[MS09a] Javier Mencía and Enrique Sentana. Multivariate location-scale mixtures of normals and mean-variance-skewness portfolio allocation. Journal of Econometrics, 153(2):105–121, December 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609001262.

Moon:2009:EOI

[MS09b] Hyungsik Roger Moon and Frank Schorfheide. Estimation with overidentifying inequality moment conditions. Journal of Econometrics, 153(2):136–154, December 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609001407.

Marcellino:2006:CDI

[MSW06] Massimiliano Marcellino, James H. Stock, and Mark W. Watson. A comparison of direct and iterated multistep AR methods for forecasting macroeconomic time series. Journal of Econometrics, 135(1–2):499–526, November/December 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760500165X.

Munkin:2003:BAS

[MT03] Murat K. Munkin and Pravin K. Trivedi. Bayesian analysis of a self-selection model with multiple outcomes using simulationbased estimation: an application to the demand for healthcare. Journal of Econometrics, 114(2):197–220, June 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602002233.

Millimet:2008:EHD

[MT08a] Daniel L. Millimet and Rusty Tchernis. Estimating highdimensional demand systems in the presence of many binding non-negativity constraints. Journal of Econometrics, 147(2):384– 395, December 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001528.

Munkin:2008:BAO

[MT08b] Murat K. Munkin and Pravin K. Trivedi. Bayesian analysis of the ordered probit model with endogenous selection. Journal of Econometrics, 143(2):334–348, April 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607002163.

Mukherjee:2007:GRE

[Muk07] Kanchan Mukherjee. Generalized r-estimators under conditional heteroscedasticity. Journal of Econometrics, 141(2):383–415, December 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606002004.

Muller:2005:SPT

[Mül05] Ulrich K. Müller. Size and power of tests of stationarity in highly autocorrelated time series. Journal of Econometrics, 128(2):195– 213, October 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760400154X.

Mullin:2006:IEC

[Mul06] Charles H. Mullin. Identification and estimation with contaminated data: When do covariate data sharpen inference? Journal of Econometrics, 130(2):253–272, February 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000540.

Muller:2007:TRL

[Mül07] Ulrich K. Müller. A theory of robust long-run variance estimation. Journal of Econometrics, 141(2):1331–1352, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607000462.

Marmol:2002:TSV

[MV02] Francesc Marmol and Carlos Velasco. Trend stationarity versus long-range dependence in time series analysis. Journal of Econometrics, 108(1):25–42, May 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000999.

Martens:2007:MVR

[MvD07] Martin Martens and Dick van Dijk. Measuring volatility with the realized range. Journal of Econometrics, 138(1):181–207, May 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606000923.

Murtazashvili:2008:FEI

[MW08] Irina Murtazashvili and Jeffrey M. Wooldridge. Fixed effects instrumental variables estimation in correlated random coefficient panel data models. Journal of Econometrics, 142(1):539–552, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407607001741.

Marsh:2004:BSG

[MZ04] Lawrence C. Marsh and Arnold Zellner. Bayesian solutions to graduate admissions and related selection problems. Journal of Econometrics, 121(1–2):405–426, July 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002550.

Nakatsuma:2000:BAA

[Nak00] Teruo Nakatsuma. Bayesian analysis of ARMA-GARCH models: A Markov chain sampling approach. Journal of Econometrics, 95(1):57–69, March 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407699000299.

Nelson:2008:BND

[Nel08] Charles R. Nelson. The Beveridge–Nelson decomposition in retrospect and prospect. Journal of Econometrics, 146(2):202–206, October 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608001024.

Nevo:2002:SSI

[Nev02] Aviv Nevo. Sample selection and information-theoretic alternatives to GMM. Journal of Econometrics, 107(1–2):149–157, March 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601001178.

Nicoletti:2006:NDP

[Nic06] Cheti Nicoletti. Nonresponse in dynamic panel data models. Journal of Econometrics, 132(2):461–489, June 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000631.

Nielsen:2000:SEA

[Nie00] Soren Feodor Nielsen. On simulated EM algorithms. Journal of Econometrics, 96(2):267–292, June 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000603.

Nobile:2000:CBM

[Nob00] Agostino Nobile. Comment: Bayesian multinomial probit models with a normalization constraint. Journal of Econometrics, 99(2):335–345, December 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760000035X.

Ni:2003:NPF

[NS03] Shawn Ni and Dongchu Sun. Noninformative priors and frequentist risks of Bayesian estimators of vector-autoregressive models. Journal of Econometrics, 115(1):159–197, July 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760300099X.

Nelson:2007:ZIL

[NS07a] Charles R. Nelson and Richard Startz. The zero-information-limit condition and spurious inference in weakly identified models. Journal of Econometrics, 138(1):47–62, May 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000881.

Nielsen:2007:DCR

[NS07b] Morten Ørregaard Nielsen and Katsumi Shimotsu. Determining the cointegrating rank in nonstationary fractional systems by the exact local Whittle approach. Journal of Econometrics, 141(2):574–596, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002077.

ODonnell:2005:BAI

[OC05] Christopher J. O’Donnell and Timothy J. Coelli. A Bayesian approach to imposing curvature on distance functions. Journal of Econometrics, 126(2):493–523, June 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001216.

Omori:2007:SVL

[OCSN07] Yasuhiro Omori, Siddhartha Chib, Neil Shephard, and Jouchi Nakajima. Stochastic volatility with leverage: Fast and efficient likelihood inference. Journal of Econometrics, 140(2):425– 449, October 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606001436.

Orbe:2005:NET

[OFRP05] Susan Orbe, Eva Ferreira, and Juan Rodriguez-Poo. Nonparametric estimation of time varying parameters under shape restrictions. Journal of Econometrics, 126(1):53–77, May 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000399.

Oberhelman:2000:APC

[OK00] Dennis Oberhelman and K. Rao Kadiyala. Asymptotic probability concentrations and finite sample properties of modified LIML estimators for equations with more than two endogenous variables. Journal of Econometrics, 98(1):163–185, September 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000822.

Okui:2009:OCM

[Oku09] Ryo Okui. The optimal choice of moments in dynamic panel data models. Journal of Econometrics, 151(1):1–16, July 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609001109.

Osiewalski:2004:BCB

[OP04] Jacek Osiewalski and Mateusz Pipień. Bayesian comparison of bivariate ARCH-type models for the main exchange rates in Poland. Journal of Econometrics, 123(2):371–391, December 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760300294X.

Omtzigt:2005:IF

[OP05] Pieter Omtzigt and Paolo Paruolo. Impact factors. Journal of Econometrics, 128(1):31–68, September 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001496.

Okhrin:2006:DPP

[OS06] Yarema Okhrin and Wolfgang Schmid. Distributional properties of portfolio weights. Journal of Econometrics, 134(1):235–256, September 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001442.

Ortelli:2005:REM

[OT05] Claudio Ortelli and Fabio Trojani. Robust efficient method of moments. Journal of Econometrics, 128(1):69–97, September 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001502.

Otsu:2008:CEL

[Ots08] Taisuke Otsu. Conditional empirical likelihood estimation and inference for quantile regression models. Journal of Econometrics, 142(1):508–538, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760700173X.

Oulton:2008:CIC

[Oul08] Nicholas Oulton. Chain indices of the cost-of-living and the path-dependence problem: An empirical solution. Journal of Econometrics, 144(1):306–324, May 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000201.

Oh:2008:RBB

[OZC08] Kum Hwa Oh, Eric Zivot, and Drew Creal. The relationship between the Beveridge–Nelson decomposition and other permanent-transitory decompositions that are popular in economics. Journal of Econometrics, 146(2):207–219, October 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760800105X.

Palandri:2009:SCC

[Pal09] Alessandro Palandri. Sequential conditional correlations: Inference and evaluation. Journal of Econometrics, 153(2):122–132, December 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001274.

Park:2002:NNH

[Par02] Joon Y. Park. Nonstationary nonlinear heteroskedasticity. Journal of Econometrics, 110(2):383–415, October 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001008.

Park:2003:SIV

[Par03] Sangin Park. Semiparametric instrumental variables estimation. Journal of Econometrics, 112(2):381–399, February 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602002038.

Park:2006:BTW

[Par06a] Joon Y. Park. A bootstrap theory for weakly integrated processes. Journal of Econometrics, 133(2):639–672, August 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001296.

Paruolo:2006:CTC

[Par06b] Paolo Paruolo. Common trends and cycles in i(2) VAR systems. Journal of Econometrics, 132(1):143–168, May 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000448.

Park:2009:MEA

[PB09] Sung Y. Park and Anil K. Bera. Maximum entropy autoregressive conditional heteroskedasticity model. Journal of Econometrics, 150(2):219–230, June 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407608002200.

Pelletier:2006:RSD

[Pel06] Denis Pelletier. Regime switching for dynamic correlations. Journal of Econometrics, 131(1–2):445–473, March/April 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000187.

Pere:2000:AEW

[Per00] Pekka Pere. Adjusted estimates and Wald statistics for the AR(1) model with constant. Journal of Econometrics, 98(2):335– 363, October 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000233.

Pesavento:2004:AEP

[Pes04] Elena Pesavento. Analytical evaluation of the power of tests for the absence of cointegration. Journal of Econometrics, 122(2):349–384, October 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603002793.

Pesaran:2007:PWA

[Pes07] M. Hashem Pesaran. A pair-wise approach to testing for output and growth convergence. Journal of Econometrics, 138(1): 312–355, May 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000972.

Pizzinga:2008:RKF

[PFC08] Adrian Pizzinga, Cristiano Fernandes, and Sergio Contreras. Restricted Kalman filtering revisited. Journal of Econometrics, 144(2):428–429, June 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407608000377.

Phillips:2000:AAO

[Phi00] Garry D. A. Phillips. An alternative approach to obtaining nagartype moment approximations in simultaneous equation models. Journal of Econometrics, 97(2):345–364, August 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000755.

Phillips:2001:TTS

[Phi01] Peter C. B. Phillips. Trending time series and macroeconomic activity: Some present and future challenges. Journal of Econometrics, 100(1):21–27, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000488.

Phillips:2002:NUR

[Phi02] Peter C. B. Phillips. New unit root asymptotics in the presence of deterministic trends. Journal of Econometrics, 111(2):323–353, December 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407602001094.

Phillips:2007:URL

[Phi07] Peter C. B. Phillips. Unit root log periodogram regression. Journal of Econometrics, 138(1):104–124, May 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760600090X.

Phillips:2009:LML

[Phi09] Peter C. B. Phillips. Long memory and long run variation. Journal of Econometrics, 151(2):150–158, August 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000797.

Phillips:2007:NDC

[PJH07] Peter C. B. Phillips, Sainan Jin, and Ling Hu. Nonstationary discrete choice: A corrigendum and addendum. Journal of Econometrics, 141(2):1115–1130, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000267.

Ploberger:2004:CCT

[Plo04] Werner Ploberger. A complete class of tests when the likelihood is locally asymptotically quadratic. Journal of Econometrics, 118(1–2):67–94, January 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603001350.

Phillips:2007:LTM

[PM07] Peter C. B. Phillips and Tassos Magdalinos. Limit theory for moderate deviations from a unit root. Journal of Econometrics, 136(1):115–130, January 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002034.

Paul:2005:PDP

[PN05] Catherine J. Morrison Paul and Richard Nehring. Product diversification, production systems, and economic performance in U.S. agricultural production. Journal of Econometrics, 126(2): 525–548, June 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604001228.

Pollock:2000:TET

[Pol00] D. S. G. Pollock. Trend estimation and de-trending via rational square-wave filters. Journal of Econometrics, 99(2):317–334, December 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000282.

Posch:2009:SEJ

[Pos09] Olaf Posch. Structural estimation of jump-diffusion processes in macroeconomics. Journal of Econometrics, 153(2):196–210, December 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001456.

Poterba:2003:SOH

[Pot03] James M. Poterba. Some observations on health status and economic status. Journal of Econometrics, 112(1):65–67, January 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602001471.

Pena:2004:FND

[PP04a] Daniel Peña and Pilar Poncela. Forecasting with nonstationary dynamic factor models. Journal of Econometrics, 119(2):291– 321, April 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603001982.

Potscher:2004:CET

[PP04b] Benedikt M. Pötscher and Ingmar R. Prucha. Contributions to econometrics, time-series analysis, and systems identification: a Festschrift in honor of manfred deistler. Journal of Econometrics, 118(1–2):1–5, January 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001313.

Phillips:2004:NIV

[PPC04] Peter C. B. Phillips, Joon Y. Park, and Yoosoon Chang. Nonlinear instrumental variable estimation of an autoregression. Journal of Econometrics, 118(1–2):219–246, January 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001416.

Parker:2006:URT

[PPP06] Cameron Parker, Efstathios Paparoditis, and Dimitris N. Politis. Unit root testing via the stationary bootstrap. Journal of Econometrics, 133(2):601–638, August 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001284.

Perron:2006:ERS

[PQ06] Pierre Perron and Zhongjun Qu. Estimating restricted structural change models. Journal of Econometrics, 134(2):373–399, October 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760500148X.

Perez-Quiros:2001:BCA

[PQT01] Gabriel Perez-Quiros and Allan Timmermann. Business cycle asymmetries in stock returns: Evidence from higher order moments and conditional densities. Journal of Econometrics, 103(1–2):259–306, July 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407601000458.

Perron:2003:GDE

[PR03] Pierre Perron and Gabriel Rodríguez. GLS detrending, efficient unit root tests and structural change. Journal of Econometrics, 115(1):1–27, July 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603000903.

Perraudin:2000:DRA

[PS00] William R. M. Perraudin and Bent E. Sørensen. The demand for risky assets: Sample selection and household portfolios. Journal of Econometrics, 97(1):117–144, July 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769900069X.

Phillips:2007:BDP

[PS07a] Peter C. B. Phillips and Donggyu Sul. Bias in dynamic panel estimation with fixed effects, incidental trends and cross section dependence. Journal of Econometrics, 137(1):162–188, March 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760600042X.

Poskitt:2007:ADT

[PS07b] D. S. Poskitt and C. L. Skeels. Approximating the distribution of the two-stage least squares estimator when the concentration parameter is small. Journal of Econometrics, 139(1):217–236, July 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407606001096.

Panagiotelis:2008:BIS

[PS08] Anastasios Panagiotelis and Michael Smith. Bayesian identification, selection and estimation of semiparametric functions in highdimensional additive models. Journal of Econometrics, 143(2):

291–316, April 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760700214X.

Prokhorov:2009:GRR

[PS09a] Artem Prokhorov and Peter Schmidt. GMM redundancy results for general missing data problems. Journal of Econometrics, 151 (1):47–55, July 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609000700.

Prokhorov:2009:LBE

[PS09b] Artem Prokhorov and Peter Schmidt. Likelihood-based estimation in a panel setting: Robustness, redundancy and validity of copulas. Journal of Econometrics, 153(1):93–104, November 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609001432.

Pesaran:2000:SAV

[PSS00] M. Hashem Pesaran, Yongcheol Shin, and Richard J. Smith. Structural analysis of vector error correction models with exogenous i (1) variables. Journal of Econometrics, 97(2):293–343, August 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000731.

Park:2003:CSE

[PSS03a] Byeong U. Park, Robin C. Sickles, and Léopold Simar. Corrigendum to “Semiparametric-efficient estimation of AR(1) panel data models”: [J. Econom. 117 (2003) 279–309]. Journal of Econometrics, 117(2):311, December 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001945.

Park:2003:SEE

[PSS03b] Byeong U. Park, Robin C. Sickles, and Léopold Simar. Semiparametric-efficient estimation of AR(1) panel data models. Journal of Econometrics, 117(2):279–309, December 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0304407603001490. See corrigendum [PSS03a].

Park:2007:SEE

[PSS07a] Byeong U. Park, Robin C. Sickles, and Léopold Simar. Semiparametric efficient estimation of dynamic panel data models. Journal of Econometrics, 136(1):281–301, January 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000352.

Pinkse:2007:CLT

[PSS07b] Joris Pinkse, Lihong Shen, and Margaret Slade. A central limit theorem for endogenous locations and complex spatial interactions. Journal of Econometrics, 140(1):215–225, September 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606002296.

Park:2008:LLE

[PSZ08] Byeong U. Park, Léopold Simar, and Valentin Zelenyuk. Local likelihood estimation of truncated regression and its partial derivatives: Theory and application. Journal of Econometrics, 146(1):185–198, September 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760800095X.

Pesaran:2005:SSP

[PT05] M. Hashem Pesaran and Allan Timmermann. Small sample properties of forecasts from autoregressive models under structural breaks. Journal of Econometrics, 129(1–2):183–217, November 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001691.

Patton:2007:POF

[PT07a] Andrew J. Patton and Allan Timmermann. Properties of optimal forecasts under asymmetric loss and nonlinearity. Journal of Econometrics, 140(2):884–918, October 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606001606.

Pesaran:2007:SEW

[PT07b] M. Hashem Pesaran and Allan Timmermann. Selection of estimation window in the presence of breaks. Journal of Econometrics, 137(1):134–161, March 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000418.

Papke:2008:PDM

[PW08] Leslie E. Papke and Jeffrey M. Wooldridge. Panel data methods for fractional response variables with an application to test pass rates. Journal of Econometrics, 145(1–2):121–133, July 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760800050X.

Pan:2008:ETP

[PWT08] Jiazhu Pan, Hui Wang, and Howell Tong. Estimation and tests for power-transformed and threshold GARCH models. Journal of Econometrics, 142(1):352–378, January 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001443.

Pesaran:2008:TSH

[PY08a] M. Hashem Pesaran and Takashi Yamagata. Testing slope homogeneity in large panels. Journal of Econometrics, 142(1):50– 93, January 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607001224.

Polonik:2008:TMV

[PY08b] Wolfgang Polonik and Qiwei Yao. Testing for multivariate volatility functions using minimum volume sets and inverse regression. Journal of Econometrics, 147(1):151–162, November 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001322.

Perron:2009:EDT

[PY09a] Pierre Perron and Tomoyoshi Yabu. Estimating deterministic trends with an integrated or stationary noise component. Journal of Econometrics, 151(1):56–69, July 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000712.

Phillips:2009:TSR

[PY09b] Peter C. B. Phillips and Jun Yu. A two-stage realized volatility approach to estimation of diffusion processes with discrete data. Journal of Econometrics, 150(2):139–150, June 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608002145.

Perron:2005:SBD

[PZ05] Pierre Perron and Xiaokang Zhu. Structural breaks with deterministic and stochastic trends. Journal of Econometrics, 129(1–2):65–119, November 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604001666.

Qu:2008:TSC

[Qu08] Zhongjun Qu. Testing for structural change in regression quantiles. Journal of Econometrics, 146(1):170–184, September 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608000948.

Racine:2000:CCV

[Rac00] Jeff Racine. Consistent cross-validatory model-selection for dependent data: hv-block cross-validation. Journal of Econometrics, 99(1):39–61, November 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407600000300.

Ramalho:2002:RMC

[Ram02] Esmeralda A. Ramalho. Regression models for choice-based samples with misclassification in the response variable. Journal of Econometrics, 106(1):171–201, January 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760100094X.

Reback:2004:ICC

[Reb04] Randall Reback. The impact of college course offerings on the supply of academically talented public school teachers.

Journal of Econometrics, 121(1–2):377–404, July 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002616.

Reif:2007:ABR

[Rei07] Jirí Reif. Asymptotic behaviour of regression pre-test estimators with minimal Bayes risk. Journal of Econometrics, 140(2):413– 424, October 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606001424.

Robinson:2003:HOK

[RH03] Peter M. Robinson and Marc Henry. Higher-order kernel semiparametric m-estimation of long memory. Journal of Econometrics, 114(1):1–27, May 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407602002087.

Rabe-Hesketh:2005:MLE

[RHSP05] Sophia Rabe-Hesketh, Anders Skrondal, and Andrew Pickles. Maximum likelihood estimation of limited and discrete dependent variable models with nested random effects. Journal of Econometrics, 128(2):301–323, October 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001599.

Robinson:2005:CFS

[RI05] P. M. Robinson and F. Iacone. Cointegration in fractional systems with deterministic trends. Journal of Econometrics, 129(1–2):263–298, November 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760400171X.

Rockinger:2002:EDA

[RJ02] Michael Rockinger and Eric Jondeau. Entropy densities with an application to autoregressive conditional skewness and kurtosis. Journal of Econometrics, 106(1):119–142, January 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000926.

Racine:2004:NER

[RL04] Jeff Racine and Qi Li. Nonparametric estimation of regression functions with both categorical and continuous data. Journal of Econometrics, 119(1):99–130, March 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760300157X.

Racine:2007:VRM

[RM07] Jeffrey S. Racine and Esfandiar Maasoumi. A versatile and robust metric entropy test of time-reversibility, and other hypotheses. Journal of Econometrics, 138(2):547–567, June 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000820.

Cheng:2008:GAS

[rMCGJL08] Ai ru (Meg) Cheng, A. Ronald Gallant, Chuanshu Ji, and Beom S. Lee. A Gaussian approximation scheme for computation of option prices in stochastic volatility models. Journal of Econometrics, 146(1):44–58, September 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000742.

Robinson:2001:MSV

[Rob01] P. M. Robinson. The memory of stochastic volatility models. Journal of Econometrics, 101(2):195–218, April 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000798.

Robins:2003:GMC

[Rob03] James M. Robins. General methodological considerations. Journal of Econometrics, 112(1):89–106, January 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001513.

Robinson:2005:DBR

[Rob05] P. M. Robinson. The distance between rival nonstationary fractional processes. Journal of Econometrics, 128(2):283–300, October 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001575.

Robinson:2008:CTT

[Rob08a] P. M. Robinson. Correlation testing in time series, spatial and cross-sectional data. Journal of Econometrics, 147(1):5–16, November 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608001206.

Robinson:2008:DTC

[Rob08b] P. M. Robinson. Diagnostic testing for cointegration. Journal of Econometrics, 143(1):206–225, March 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001698.

Rosen:2008:CSP

[Ros08] Adam M. Rosen. Confidence sets for partially identified parameters that satisfy a finite number of moment inequalities. Journal of Econometrics, 146(1):107–117, September 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000791.

Rothenberg:2002:SED

[Rot02] Thomas J. Rothenberg. Some elementary distribution theory for an autoregression fitted to a random walk. Journal of Econometrics, 111(2):355–361, December 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001100.

Rothstein:2004:CPP

[Rot04] Jesse M. Rothstein. College performance predictions and the SAT. Journal of Econometrics, 121(1–2):297–317, July 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002537.

Rothe:2009:SEB

[Rot09] Christoph Rothe. Semiparametric estimation of binary response models with endogenous regressors. Journal of Econometrics, 153(1):51–64, November 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609001110.

Ramsay:2002:FDA

[RR02] James O. Ramsay and James B. Ramsey. Functional data analysis of the dynamics of the monthly index of nondurable goods production. Journal of Econometrics, 107(1–2):327–344, March 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601001270.

Ryu:2000:EDU

[RS00] Hang K. Ryu and Daniel J. Slottje. Estimating the density of unemployment duration based on contaminated samples or small samples. Journal of Econometrics, 95(1):131–156, March 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000330.

Ramalho:2002:GEL

[RS02] Joaquim J. S. Ramalho and Richard J. Smith. Generalized empirical likelihood non-nested tests. Journal of Econometrics, 107(1–2):99–125, March 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407601001154.

Rilstone:1996:SOB

[RSU96] Paul Rilstone, V. K. Srivastava, and Aman Ullah. The secondorder bias and mean squared error of nonlinear estimators. Journal of Econometrics, 75(2):369–395, December 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0304407696894577. See corrigendum [RU05].

Ronchetti:2001:RIG

[RT01] Elvezio Ronchetti and Fabio Trojani. Robust inference with GMM estimators. Journal of Econometrics, 101(1):37–69, March 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000737.

Rodrigues:2004:AEU

[RT04] Paulo M. M. Rodrigues and A. M. Robert Taylor. Alternative estimators and unit root tests for seasonal autoregressive processes. Journal of Econometrics, 120(1):35–73, May 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002069.

Radchenko:2006:LIB

[RT06] Stanislav Radchenko and Hiroki Tsurumi. Limited information Bayesian analysis of a simultaneous equation with an autocorrelated error term and its application to the U.S. gasoline market. Journal of Econometrics, 133(1):31–49, July 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000825.

Rodrigues:2007:ETS

[RT07] Paulo M. M. Rodrigues and A. M. Robert Taylor. Efficient tests of the seasonal unit root hypothesis. Journal of Econometrics, 141(2):548–573, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002065.

Rilstone:2005:CSO

[RU05] Paul Rilstone and Aman Ullah. Corrigendum to “The secondorder bias and mean squared error of nonlinear estimators”: [Journal of Econometrics 75(2) (1996) 369–395]. Journal of Econometrics, 124(1):203–204, January 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000521. See [RSU96].

Reif:2002:OPT

[RV02] Jiri Reif and Karel Vlcek. Optimal pre-test estimators in regression. Journal of Econometrics, 110(1):91–102, September 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602001240.

Robinson:2002:DCR

[RY02] Peter M. Robinson and Yoshihiro Yajima. Determination of cointegrating rank in fractional systems. Journal of Econometrics, 106(2):217–241, February 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407601000963.

Richard:2007:EHD

[RZ07] Jean-François Richard and Wei Zhang. Efficient high-dimensional importance sampling. Journal of Econometrics, 141(2):1385– 1411, December 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607000486.

Sandor:2004:ASM

[SA04] Zsolt Sándor and Péter András. Alternative sampling methods for estimating multivariate normal probabilities. Journal of Econometrics, 120(2):207–234, June 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002124.

Saikkonen:2005:SRN

[Sai05] Pentti Saikkonen. Stability results for nonlinear error correction models. Journal of Econometrics, 127(1):69–81, July 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001381.

Sakata:2007:IVE

[Sak07] Shinichi Sakata. Instrumental variable estimation based on conditional median restriction. Journal of Econometrics, 141(2):350– 382, December 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606001990.

Sancetta:2007:OFC

[San07] Alessio Sancetta. Online forecast combinations of distributions: Worst case bounds. Journal of Econometrics, 141(2):621–651, December 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606002090.

Sareen:2003:RBI

[Sar03] Samita Sareen. Reference Bayesian inference in nonregular models. Journal of Econometrics, 113(2):265–288, April 2003. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760200204X.

Smith:2002:FSA

[SB02] Richard J. Smith and H. Peter Boswijk. Finite sample and asymptotic methods in econometrics. Journal of Econometrics, 111(2):135–140, December 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760200101X.

Stoker:2005:PDA

[SBES05] Thomas M. Stoker, Ernst R. Berndt, A. Denny Ellerman, and Susanne M. Schennach. Panel data analysis of U.S. coal productivity. Journal of Econometrics, 127(2):131–164, August 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001411.

Sain:2007:SMM

[SC07] Stephan R. Sain and Noel Cressie. A spatial model for multivariate lattice data. Journal of Econometrics, 140(1):226–259, September 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606002302.

Sentana:2008:IEL

[SCF08] Enrique Sentana, Giorgio Calzolari, and Gabriele Fiorentini. Indirect estimation of large conditionally heteroskedastic factor models, with an application to the Dow 30 stocks. Journal of Econometrics, 146(1):10–25, September 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000626.

Schonfeld:2004:LSG

[Sch04] Peter Schönfeld. Least squares in general vector spaces revisited. Journal of Econometrics, 118(1–2):95–109, January 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001362.

Schorfheide:2005:VFU

[Sch05] Frank Schorfheide. VAR forecasting under misspecification. Journal of Econometrics, 128(1):99–136, September 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001514.

Swanson:2006:PMA

[SEGG06] Norman R. Swanson, Graham Elliott, Eric Ghysels, and Jesus Gonzalo. Predictive methodology and application in economics and finance: Volume in honor of the accomplishments of Clive W. J. Granger. Journal of Econometrics, 135(1–2):1–9, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001612.

Sentana:2004:FRP

[Sen04] Enrique Sentana. Factor representing portfolios in large asset markets. Journal of Econometrics, 119(2):257–289, April 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001970.

Seo:2006:BTN

[Seo06] Myunghwan Seo. Bootstrap testing for the null of no cointegration in a threshold vector error correction model. Journal of Econometrics, 134(1):129–150, September 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001405.

Seo:2007:ADC

[Seo07] Byeongseon Seo. Asymptotic distribution of the cointegrating vector estimator in error correction models with conditional heteroskedasticity. Journal of Econometrics, 137(1):68–111, March 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760600039X.

Sentana:2001:IET

[SF01] Enrique Sentana and Gabriele Fiorentini. Identification, estimation and testing of conditionally heteroskedastic factor models. Journal of Econometrics, 102(2):143–164, June 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000513.

Silver:2007:WEP

[SH07] Mick Silver and Saeed Heravi. Why elementary price index number formulas differ: Evidence on price dispersion. Journal of Econometrics, 140(2):874–883, October 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760600159X.

Shintani:2001:SCR

[Shi01] Mototsugu Shintani. A simple cointegrating rank test without vector autoregression. Journal of Econometrics, 105(2):337–362, December 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000847.

Shimotsu:2007:GSE

[Shi07] Katsumi Shimotsu. Gaussian semiparametric estimation of multivariate fractionally integrated processes. Journal of Econometrics, 137(2):277–310, April 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606000443.

Strachan:2004:BAE

[SI04] Rodney W. Strachan and Brett Inder. Bayesian analysis of the error correction model. Journal of Econometrics, 123(2):307–325, December 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603002951.

Sickles:2005:PEI

[Sic05] Robin C. Sickles. Panel estimators and the identification of firmspecific efficiency levels in parametric, semiparametric and nonparametric settings. Journal of Econometrics, 126(2):305–334, June 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604001149.

Sims:2000:ULP

[Sim00] Christopher A. Sims. Using a likelihood perspective to sharpen econometric discourse: Three examples. Journal of Econometrics, 95(2):443–462, April 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407699000469.

Singleton:2001:EAA

[Sin01] Kenneth J. Singleton. Estimation of affine asset pricing models using the empirical characteristic function. Journal of Econometrics, 102(1):111–141, May 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000920.

Smith:2000:NSU

[SK00] Michael Smith and Robert Kohn. Nonparametric seemingly unrelated regression. Journal of Econometrics, 98(2):257–281, October 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760000018X.

Shin:2006:IVA

[SK06a] Dong Wan Shin and Seungho Kang. An instrumental variable approach for panel unit root tests under cross-sectional dependence. Journal of Econometrics, 134(1):215–234, September 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001430.

Sriananthakumar:2006:NAP

[SK06b] Sivagowry Sriananthakumar and Maxwell L. King. A new approximate point optimal test of a composite null hypothesis. Journal of Econometrics, 130(1):101–122, January 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001629.

Skouras:2007:DDB

[Sko07] Spyros Skouras. Decisionmetrics: A decision-based approach to econometric modelling. Journal of Econometrics, 137(2):414– 440, April 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000510.

Shin:2003:IVA

[SL03] Dong Wan Shin and Oesook Lee. An instrumental variable approach for tests of unit roots and seasonal unit roots in asymmetric time series models. Journal of Econometrics, 115(1):29–52, July 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603000915.

Shintani:2004:NNN

[SL04] Mototsugu Shintani and Oliver Linton. Nonparametric neural network estimation of Lyapunov exponents and a direct test for chaos. Journal of Econometrics, 120(1):1–33, May 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603002057.

Seo:2007:SLS

[SL07a] Myung Hwan Seo and Oliver Linton. A smoothed least squares estimator for threshold regression models. Journal of Econometrics, 141(2):704–735, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760600234X.

Shin:2007:ANS

[SL07b] Dong Wan Shin and Oesook Lee. Asymmetry and nonstationarity for a seasonal time series model. Journal of Econometrics, 136(1):89–114, January 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407605002010.

Slottje:2008:EDS

[Slo08] Daniel Slottje. Estimating demand systems and measuring consumer preferences. Journal of Econometrics, 147(2):207–209, December 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760800136X.

Spanos:2002:PNM

[SM02] Aris Spanos and Anya McGuirk. The problem of nearmulticollinearity revisited: erratic vs systematic volatility. Journal of Econometrics, 108(2):365–393, June 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001440.

Slottje:2007:EAC

[SMB07] Daniel J. Slottje, Daniel L. Millimet, and Michael J. Buchanan. Econometric analysis of copyrights. Journal of Econometrics, 139(2):303–317, August 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407606002168.

Smith:2007:EIT

[Smi07] Richard J. Smith. Efficient information theoretic inference for conditional moment restrictions. Journal of Econometrics, 138(2): 430–460, June 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000777.

Smith:2006:MSM

[SNT06] Aaron Smith, Prasad A. Naik, and Chih-Ling Tsai. Markovswitching model selection using Kullback–Leibler divergence. Journal of Econometrics, 134(2):553–577, October 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001557.

Shin:2004:FMS

[SO04] Dong Wan Shin and Man-Suk Oh. Fully modified semiparametric GLS estimation for regressions with nonstationary seasonal regressors. Journal of Econometrics, 122(2):247–280, October 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603002756.

Shen:2001:MEB

[SP01] Edward Z. Shen and Jeffrey M. Perloff. Maximum entropy and Bayesian approaches to the ratio problem. Journal of Econometrics, 104(2):289–313, September 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000823.

Sun:2003:NLP

[SP03] Yixiao Sun and Peter C. B. Phillips. Nonlinear log-periodogram regression for perturbed fractional processes. Journal of Econometrics, 115(2):355–389, August 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603001155.

Shimotsu:2006:LWE

[SP06] Katsumi Shimotsu and Peter C. B. Phillips. Local Whittle estimation of fractional integration and some of its variants. Journal of Econometrics, 130(2):209–233, February 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000527.

Soofi:2002:IIU

[SR02] E. S. Soofi and J. J. Retzer. Information indices: unification and applications. Journal of Econometrics, 107(1–2):17–40, March 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601001117.

Shin:2000:GTS

[SS00] Dong Wan Shin and Beong Soo So. Gaussian tests for seasonal unit roots based on Cauchy estimation and recursive mean adjustments. Journal of Econometrics, 99(1):107–137, November 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000324.

So:2001:IST

[SS01] Beong Soo So and Dong Wan Shin. An invariant sign test for random walks based on recursive median adjustment. Journal of Econometrics, 102(2):197–229, June 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000537.

Stinebrickner:2004:TUC

[SS04] Ralph Stinebrickner and Todd R. Stinebrickner. Time-use and college outcomes. Journal of Econometrics, 121(1–2):243–269, July 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407603002628.

Sun:2006:SEA

[SS06] Yiguo Sun and Thanasis Stengos. Semiparametric efficient adaptive estimation of asymmetric GARCH models. Journal of Econometrics, 133(1):373–386, July 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605000941.

Schmid:2009:EMI

[SS09] Matthias Schmid and Hans Schneeweiss. The effect of microaggregation by individual ranking on the estimation of moments. Journal of Econometrics, 153(2):174–182, December 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609001420.

Shaikh:2009:STP

[SSVY09] Azeem M. Shaikh, Marianne Simonsen, Edward J. Vytlacil, and Nese Yildiz. A specification test for the propensity score using its distribution conditional on participation. Journal of Econometrics, 151(1):33–46, July 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000499.

Severini:2001:SAC

[ST01a] Thomas A. Severini and Gautam Tripathi. A simplified approach to computing efficiency bounds in semiparametric models. Journal of Econometrics, 102(1):23–66, May 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000907.

Smith:2001:RRR

[ST01b] Richard J. Smith and A. M. Robert Taylor. Recursive and rolling regression-based tests of the seasonal unit root hypothesis. Journal of Econometrics, 105(2):309–336, December 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000835.

Schluter:2002:TLC

[ST02] Christian Schluter and Mark Trede. Tails of Lorenz curves. Journal of Econometrics, 109(1):151–166, July 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001452.

Smith:2005:R

[ST05a] Jeffrey Smith and Petra Todd. Rejoinder. Journal of Econometrics, 125(1–2):365–375, March 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407604000843.

Smith:2005:DMO

[ST05b] Jeffrey A. Smith and Petra E. Todd. Does matching overcome LaLonde’s critique of nonexperimental estimators? Journal of Econometrics, 125(1–2):305–353, March 2005. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760400082X.

Stock:2001:ME

[Sto01] James H. Stock. Macro-econometrics. Journal of Econometrics, 100(1):29–32, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760000049X.

Stoye:2009:MRT

[Sto09] Jörg Stoye. Minimax regret treatment choice with finite samples. Journal of Econometrics, 151(1):70–81, July 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000724.

Stutzer:2003:PCE

[Stu03] Michael Stutzer. Portfolio choice with endogenous utility: a large deviations approach. Journal of Econometrics, 116(1–2):365–386, September 2003. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760300112X.

Sullivan:2001:DDM

[STW01] Ryan Sullivan, Allan Timmermann, and Halbert White. Dangers of data mining: The case of calendar effects in stock returns. Journal of Econometrics, 105(1):249–286, November 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760100077X.

Su:2008:LPE

[SU08] Liangjun Su and Aman Ullah. Local polynomial estimation of nonparametric simultaneous equations models. Journal of Econometrics, 144(1):193–218, May 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000067.

Silverberg:2007:SDI

[SV07a] Gerald Silverberg and Bart Verspagen. The size distribution of innovations revisited: An application of extreme value statistics to citation and value measures of patent significance. Journal of Econometrics, 139(2):318–339, August 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760600217X.

Stummer:2007:OSD

[SV07b] Wolfgang Stummer and Igor Vajda. Optimal statistical decisions about some alternative financial models. Journal of Econometrics, 137(2):441–471, April 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000522.

Schluter:2009:EEN

[SvG09] Christian Schluter and Kees Jan van Garderen. Edgeworth expansions and normalizing transforms for inequality measures. Journal of Econometrics, 150(1):16–29, May 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000293.

Sakata:2001:ENR

[SW01a] Shinichi Sakata and Halbert White. S-estimation of nonlinear regression models with dependent and heterogeneous observations. Journal of Econometrics, 103(1–2):5–72, July 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000392.

Silva:2001:TPM

[SW01b] João M. C. Santos Silva and Frank Windmeijer. Twopart multiple spell models for health care demand. Journal of Econometrics, 104(1):67–89, August 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000598.

Simar:2007:EIT

[SW07a] Léopold Simar and Paul W. Wilson. Estimation and inference in two-stage, semi-parametric models of production processes. Journal of Econometrics, 136(1):31–64, January 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001594.

Su:2007:CCF

[SW07b] Liangjun Su and Halbert White. A consistent characteristic function-based test for conditional independence. Journal of Econometrics, 141(2):807–834, December 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606002375.

Sickles:2008:SIE

[SW08a] Robin C. Sickles and Jennifer Williams. Special issue editors’ introduction: The use of econometrics in informing public policy makers. Journal of Econometrics, 145(1–2):1–3, July 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000456.

Sickles:2008:TCD

[SW08b] Robin C. Sickles and Jenny Williams. Turning from crime: A dynamic perspective. Journal of Econometrics, 145(1–2):158– 173, July 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000535.

Sims:2008:MIL

[SWZ08] Christopher A. Sims, Daniel F. Waggoner, and Tao Zha. Methods for inference in large multiple-equation Markov-switching models. Journal of Econometrics, 146(2):255–274, October 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001140.

Sarafidis:2009:TCS

[SYR09] Vasilis Sarafidis, Takashi Yamagata, and Donald Robertson. A test of cross section dependence for a linear dynamic panel model with regressors. Journal of Econometrics, 148(2):149–161, February 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608001826.

Taber:2000:SIH

[Tab00] Christopher R. Taber. Semiparametric identification and heterogeneity in discrete choice dynamic programming models. Journal of Econometrics, 96(2):201–229, June 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000573.

Tauchen:2001:NFE

[Tau01] George Tauchen. Notes on financial econometrics. Journal of Econometrics, 100(1):57–64, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000543.

Taylor:2005:VRT

[Tay05] A. M. Robert Taylor. Variance ratio tests of the seasonal unit root hypothesis. Journal of Econometrics, 124(1):33–54, January 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407604000259.

Tsay:2000:SRF

[TC00] Wen-Jen Tsay and Ching-Fan Chung. The spurious regression of fractionally integrated processes. Journal of Econometrics, 96(1): 155–182, May 2000. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407699000561.

Tang:2009:PEB

[TC09] Cheng Yong Tang and Song Xi Chen. Parameter estimation and bias correction for diffusion processes. Journal of Econometrics, 149(1):65–81, April 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760800208X.

Thomsen:2000:SCD

[Tho00] Thomas Thomsen. Short cuts to dynamic factor demand modelling. Journal of Econometrics, 97(1):1–23, July 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000627.

Timmermann:2000:MMS

[Tim00] Allan Timmermann. Moments of Markov switching models. Journal of Econometrics, 96(1):75–111, May 2000. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000512.

Todorov:2009:ECT

[Tod09] Viktor Todorov. Estimation of continuous-time stochastic volatility models with jumps using high-frequency data. Journal of Econometrics, 148(2):131–148, February 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001814.

Tse:2000:TCC

[Tse00] Y. K. Tse. A test for constant correlations in a multivariate GARCH model. Journal of Econometrics, 98(1):107–127, September 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000809.

Turkington:2000:GVO

[Tur00] Darrell Turkington. Generalised vec operators and the seemingly unrelated regression equations model with vector correlated disturbances. Journal of Econometrics, 99(2):225–253, December 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000257.

Ullah:2002:UED

[Ull02] Aman Ullah. Uses of entropy and divergence measures for evaluating econometric approximations and inference. Journal of Econometrics, 107(1–2):313–326, March 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001269.

vanAkkeren:2002:GMB

[vAJM02] Marco van Akkeren, George Judge, and Ron Mittelhammer. Generalized moment based estimation and inference. Journal of Econometrics, 107(1–2):127–148, March 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001166.

vandenBerg:2007:UOP

[vdB07] Gerard J. van den Berg. On the uniqueness of optimal prices set by monopolistic sellers. Journal of Econometrics, 141(2):482– 491, December 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760600203X.

vandenBerg:2001:CMM

[vdBvdK01] Gerard J. van den Berg and Bas van der Klaauw. Combining micro and macro unemployment duration data. Journal of Econometrics, 102(2):271–309, June 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000550.

vanDijk:2002:NLM

[vDFP02] Dick van Dijk, Philip Hans Franses, and Richard Paap. A nonlinear long memory model, with an application to US unemployment. Journal of Econometrics, 110(2):135–165, October 2002. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602000908.

vanderKlaauw:2008:BLB

[vdK08] Wilbert van der Klaauw. Breaking the link between poverty and low student achievement: An evaluation of Title I. Journal of Econometrics, 142(2):731–756, February 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001157.

vanderKlaauw:2008:SSR

[vdKW08] Wilbert van der Klaauw and Kenneth I. Wolpin. Social security and the retirement and savings behavior of low-income households. Journal of Econometrics, 145(1–2):21–42, July 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000547.

vanDijk:2008:EIR

[vDP08] Bram van Dijk and Richard Paap. Explaining individual response using aggregated data. Journal of Econometrics, 146(1): 1–9, September 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000614.

Vogelsang:2005:TCD

[VF05] Timothy J. Vogelsang and Philip Hans Franses. Testing for common deterministic trend slopes. Journal of Econometrics, 126(1): 1–24, May 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604000375.

vanGarderen:2001:OPL

[vG01] Kees Jan van Garderen. Optimal prediction in loglinear models. Journal of Econometrics, 104(1):119–140, August 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601000616.

vanGiersbergen:2002:HIB

[vGK02] Noud P. A. van Giersbergen and Jan F. Kiviet. How to implement the bootstrap in static or stable dynamic regression models: test statistic versus confidence region approach. Journal of Econometrics, 108(1):133–156, May 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001324.

vanGarderen:2000:CSA

[vGLP00] Kees Jan van Garderen, Kevin Lee, and M. Hashem Pesaran. Cross-sectional aggregation of non-linear models. Journal of Econometrics, 95(2):285–331, April 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000408.

Vahid:2002:ICC

[VI02] Farshid Vahid and João Victor Issler. The importance of common cyclical features in VAR analysis: a Monte-Carlo study. Journal of Econometrics, 109(2):341–363, August 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001173.

Vilasuso:2001:CTC

[Vil01] Jon Vilasuso. Causality tests and conditional heteroskedasticity:: Monte Carlo evidence. Journal of Econometrics, 101(1): 25–35, March 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407600000725.

Villani:2006:BPE

[Vil06] Mattias Villani. Bayesian point estimation of the cointegration space. Journal of Econometrics, 134(2):645–664, October 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605001582.

Vinod:2001:CFR

[Vin01] H. D. Vinod. Care and feeding of reproducible econometrics. Journal of Econometrics, 100(1):87–88, January 2001. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000646.

Villani:2009:RDE

[VKG09] Mattias Villani, Robert Kohn, and Paolo Giordani. Regression density estimation using smooth adaptive Gaussian mixtures.

Journal of Econometrics, 153(2):155–173, December 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609001419.

vanSchuppen:2004:STS

[vS04] Jan H. van Schuppen. System theory for system identification. Journal of Econometrics, 118(1–2):313–339, January 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001453.

vanSoest:2002:SLS

[vSDG02] Arthur van Soest, Marcel Das, and Xiaodong Gong. A structural labour supply model with flexible preferences. Journal of Econometrics, 107(1–2):345–374, March 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001282.

Verbeek:2005:EDM

[VV05] Marno Verbeek and Francis Vella. Estimating dynamic models from repeated cross-sections. Journal of Econometrics, 127(1): 83–102, July 2005. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407604001393.

Wansbeek:2001:GEP

[Wan01] Tom Wansbeek. GMM estimation in panel data models with measurement error. Journal of Econometrics, 104(2):259–268, September 2001. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601000793.

West:2001:ETW

[Wes01] Kenneth D. West. Encompassing tests when no model is encompassing. Journal of Econometrics, 105(1):287–308, November 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407601000781.

Whang:2000:CBT

[Wha00] Yoon-Jae Whang. Consistent bootstrap tests of parametric regression functions. Journal of Econometrics, 98(1):27–46, September 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000780.

White:2006:TSE

[Whi06] Halbert White. Time-series estimation of the effects of natural experiments. Journal of Econometrics, 135(1–2):527–566, November/December 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605001636.

Willems:2004:DLS

[Wil04] J. C. Willems. Deterministic least squares filtering. Journal of Econometrics, 118(1–2):341–373, January 2004. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001465.

Windmeijer:2005:FSC

[Win05] Frank Windmeijer. A finite sample correction for the variance of linear efficient two-step GMM estimators. Journal of Econometrics, 126(1):25–51, May 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000387.

Wachter:2003:PBH

[WKV03] Kenneth W. Wachter, John E. Knodel, and Mark VanLandingham. Parental bereavement: heterogeneous impacts of AIDS in Thailand. Journal of Econometrics, 112(1):193–206, January 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407602001604.

Wooldridge:2007:IPW

[Woo07] Jeffrey M. Wooldridge. Inverse probability weighted estimation for general missing data problems. Journal of Econometrics, 141(2):1281–1301, December 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607000437.

Wu:2007:GEM

[WP07] Ximing Wu and Jeffrey M. Perloff. GMM estimation of a maximum entropy distribution with interval data. Journal of Econometrics, 138(2):532–546, June 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407606000819.

Wright:2008:BMA

[Wri08] Jonathan H. Wright. Bayesian model averaging and exchange rate forecasts. Journal of Econometrics, 146(2):329–341, October 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608001115.

Wang:2009:DET

[WS09] Wei Siang Wang and Peter Schmidt. On the distribution of estimated technical efficiency in stochastic frontier models. Journal of Econometrics, 148(1):36–45, January 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001164.

Wu:2003:CME

[Wu03] Ximing Wu. Calculation of maximum entropy densities with application to income distribution. Journal of Econometrics, 115(2):347–354, August 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603001143.

Wagstaff:2003:DCH

[WvDW03] Adam Wagstaff, Eddy van Doorslaer, and Naoko Watanabe. On decomposing the causes of health sector inequalities with an application to malnutrition inequalities in Vietnam. Journal of Econometrics, 112(1):207–223, January 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602001616.

Wagenvoort:2002:BRI

[WW02] Rien Wagenvoort and Robert Waldmann. On b-robust instrumental variable estimation of the linear model with panel data.

Journal of Econometrics, 106(2):297–324, February 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001026.

Wheelock:2008:NPU

[WW08] David C. Wheelock and Paul W. Wilson. Non-parametric, unconditional quantile estimation for efficiency analysis with an application to Federal reserve check processing operations. Journal of Econometrics, 145(1–2):209–225, July 2008. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608000572.

Wachter:2009:PRA

[WW09] Jessica A. Wachter and Missaka Warusawitharana. Predictable returns and asset allocation: Should a skeptical investor time the market? Journal of Econometrics, 148(2):162–178, February 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608002029.

Wansbeek:2001:CMJ

[WWM01] Tom Wansbeek, Michel Wedel, and Erik Meijer. Comment on “Microeconometrics” by J. A. Hausman. Journal of Econometrics, 100(1):89–91, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000658. See [Hau01a].

Waggoner:2003:LPN

[WZ03a] Daniel F. Waggoner and Tao Zha. Likelihood preserving normalization in multiple equation models. Journal of Econometrics, 114(2):329–347, June 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407603000873.

Wan:2003:OCV

[WZ03b] Alan T. K. Wan and Guohua Zou. Optimal critical values of pre-tests when estimating the regression error variance: analytical findings under a general loss structure. Journal of Econometrics, 114(1):165–196, May 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407602002245.

Xiao:2009:FCC

[Xia09a] Zhijie Xiao. Functional-coefficient cointegration models. Journal of Econometrics, 152(2):81–92, October 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609000232.

Xiao:2009:QCR

[Xia09b] Zhijie Xiao. Quantile cointegrating regression. Journal of Econometrics, 150(2):248–260, June 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407608002224.

Xie:2004:VES

[XL04] Liang-Liang Xie and Lennart Ljung. Variance expressions for spectra estimated using auto-regressions. Journal of Econometrics, 118(1–2):247–256, January 2004. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603001428.

Xiao:2002:CTC

[XP02a] Zhijie Xiao and Peter C. B. Phillips. A CUSUM test for cointegration using regression residuals. Journal of Econometrics, 108 (1):43–61, May 2002. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407601001038.

Xiao:2002:HOA

[XP02b] Zhijie Xiao and Peter C. B. Phillips. Higher order approximations for Wald statistics in time series regressions with integrated processes. Journal of Econometrics, 108(1):157–198, May 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001348.

Xu:2008:AEA

[XP08] Ke-Li Xu and Peter C. B. Phillips. Adaptive estimation of autoregressive models with time-varying variances. Journal of Econometrics, 142(1):265–280, January 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407607001303.

Xu:2009:ELB

[Xu09] Ke-Li Xu. Empirical likelihood-based inference for nonparametric recurrent diffusions. Journal of Econometrics, 153(1):65–82, November 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001122.

Yamagata:2008:JSC

[Yam08] Takashi Yamagata. A joint serial correlation test for linear panel data models. Journal of Econometrics, 146(1):135–145, September 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760800081X.

Yang:2006:SGM

[Yan06] Lijian Yang. A semiparametric GARCH model for foreign exchange volatility. Journal of Econometrics, 130(2):365–384, February 2006. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407605000801.

Yu:2008:QML

[YdJfL08] Jihai Yu, Robert de Jong, and Lung fei Lee. Quasi-maximum likelihood estimators for spatial dynamic panel data with fixed effects when both n and t are large. Journal of Econometrics, 146(1):118– 134, September 2008. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407608000808.

Yao:2000:IOW

[YH00] Feng Yao and Yuzo Hosoya. Inference on one-way effect and evidence in Japanese macroeconomic data. Journal of Econometrics, 98(2):225–255, October 2000. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000846.

Yatchew:2006:NSP

[YH06] Adonis Yatchew and Wolfgang Härdle. Nonparametric state price density estimation using constrained least squares and the bootstrap. Journal of Econometrics, 133(2):579–599, August 2006. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407605001272.

Yu:2007:BAT

[YS07] Keming Yu and Julian Stander. Bayesian analysis of a Tobit quantile regression model. Journal of Econometrics, 137(1):260– 276, March 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407606000546.

Yu:2005:LSV

[Yu05] Jun Yu. On leverage in a stochastic volatility model. Journal of Econometrics, 127(2):165–178, August 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604001435.

Yu:2007:CFL

[Yu07] Jialin Yu. Closed-form likelihood approximation and estimation of jump-diffusions with an application to the realignment risk of the Chinese Yuan. Journal of Econometrics, 141(2):1245–1280, December 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407607000425.

Zaffaroni:2004:CAL

[Zaf04] Paolo Zaffaroni. Contemporaneous aggregation of linear dynamic models in large economies. Journal of Econometrics, 120(1):75– 102, May 2004. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407603002070.

Zaffaroni:2007:AMM

[Zaf07] Paolo Zaffaroni. Aggregation and memory of models of changing volatility. Journal of Econometrics, 136(1):237–249, January 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606000339.

Zaffaroni:2009:WEE

[Zaf09] Paolo Zaffaroni. Whittle estimation of EGARCH and other exponential volatility models. Journal of Econometrics, 151(2):190– 200, August 2009. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609000827.

Zhang:2009:BAB

[ZBK09] Xibin Zhang, Robert D. Brooks, and Maxwell L. King. A Bayesian approach to bandwidth selection for multivariate kernel regression with an application to state-price density estimation. Journal of Econometrics, 153(1):21–32, November 2009. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609001080.

Zheng:2001:SIT

[ZC01] Buhong Zheng and Brian J. Cushing. Statistical inference for testing inequality indices with dependent samples. Journal of Econometrics, 101(2):315–335, April 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407600000877.

Zaffaroni:2003:GIC

[Zd03]  Paolo Zaffaroni and Banca d’Italia.  Gaussian inference

on certain long-range dependent volatility models. Journal of Econometrics, 115(2):199–258, August 2003. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407603000964.

Zellner:2001:CPE

[Zel01] Arnold Zellner. Comments on papers by Engle, Geweke and Granger. Journal of Econometrics, 100(1):93–94, January 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000671. See [Eng01, Gew01, Gra01].

Zellner:2002:IPB

[Zel02] Arnold Zellner. Information processing and Bayesian analysis. Journal of Econometrics, 107(1–2):41–50, March 2002. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407601001129.

Zellner:2007:EGS

[Zel07a] Arnold Zellner. Erratum to “Generalizing the standard product rule of probability theory and Bayes’s Theorem”: [J. Econometrics 138 (1) (2007) 14–23]. Journal of Econometrics, 141(2):1419, December 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760700067X. See [Zel07b].

Zellner:2007:GSP

[Zel07b] Arnold Zellner. Generalizing the standard product rule of probability theory and Bayes’s theorem. Journal of Econometrics, 138 (1):14–23, May 2007. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407606000868. See erratum

[Zel07a].

Zellner:2007:POE

[Zel07c] Arnold Zellner. Philosophy and objectives of econometrics. Journal of Econometrics, 136(2):331–339, February 2007. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605002174.

Zellner:2007:SAH

[Zel07d] Arnold Zellner. Some aspects of the history of Bayesian information processing. Journal of Econometrics, 138(2):388–404, June 2007. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407606000753.

Zheng:2001:SIP

[Zhe01] Buhong Zheng. Statistical inference for poverty measures with relative poverty lines. Journal of Econometrics, 101(2):337–356, April 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407600000889.

Zhang:2001:NAC

[ZRT01] Michael Yuanjie Zhang, Jeffrey R. Russell, and Ruey S. Tsay. A nonlinear autoregressive conditional duration model with applications to financial transaction data. Journal of Econometrics, 104(1):179–207, August 2001. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440760100063X.

Zhao:2009:NID

[ZW09] Zhibiao Zhao and Wei Biao Wu. Nonparametric inference of discretely sampled stable Lévy processes. Journal of Econometrics, 153(1):83–92, November 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407609001250.

Zhu:2009:PEA

[ZZW09] Dongming Zhu and Victoria Zinde-Walsh. Properties and estimation of asymmetric exponential power distribution. Journal of Econometrics, 148(1):86–99, January 2009. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407608001668.